A Bibliography of Publications in Science and Public Policy: 2010–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 0.32

Title word cross-reference

18th [Ano13]. 1960 [Gut16].

2 [ZP17]. 2.0 [MFC12]. 2012 [Tan17]. 2020 [Eng18, Fit13]. 20th [AF14]. 21st [FM16, LHGW14]. 27 [Lor11].

3 [Wil16].

5 [Lov16].

6th [Ano13].

7 [Sti17]. 71 [Dra14].

9 [Alm16, Gut16]. 978 [Alm16, Gut16, Lov16, Sti17, Wil16]. 978-0-262-02850-9 [Gut16]. 978-0-262-52691-3 [Wil16]. 978-1-4094-5491-5 [Lov16]. 978-1-78195-396-9 [Alm16]. 978-1-78254-787-7 [Sti17].

Åke [BFE11]. Aalto [KBA17]. abandoned [Gro10]. Abatement [Sha17]. Abby [Cla13b]. abroad [HR19]. Absorptive [HL19, SK18]. Abu [MAS16]. academia [SRC17, SB11]. Academic [CD16, Fit17, HFWL19, Kwi18, LH11, Sha18, AS16, ADR15, AF14, AMLL18, BLS16, Bee13, BCSF17a, BCSF17b, BOP18, CS15, CH17, Cla10, FBB14, FL13, GRLR17, HGS11, HL11, HFW11, JVH14, JC15, JNP16, Kwo11, LS14, Mur13, MRR17, Nie16a, OPMGCM14, RVDWA18, RH19, RdMdAFY16, RdR17, SM10, VJ12, WKGK11, Wu13, vHWW19]. academics [GK10, KA16, LS15, WHLhC19]. Academy [FdSdS^{+}10, Hei18, LZ10]. Academy-industry [FdSdS^{+}10]. acceptance [KCW14]. Accepting [Esp13]. Access [CRS17, FWS19]. accommodations [SM10]. accountability [LS12b]. accrediting [LG16]. accumulation [TBvB15]. Achieving [DP13, ET19]. acquisition [Hem17]. across [CHH18, HSM16, KCW14, Kor14, MBD^{+}12]. Act [OL13, Tei14, BG16]. Action [Gut16, AFW10, Gut18, Mot12]. actionable [Bis19]. activate [SC19a, SC19b]. activist [Par10]. Activities [Cat12, COC14, Kwo11, Tam13, TRGA16, dR18]. activity [KWJ17]. actor [Isk17]. actor-network [Isk17]. actors [BCSF17a, BCSF17b, WO17]. actually [Car18]. Adam [Fra12, Hoy13, Tra14]. Adaptation [Ken16, Bis19, Gho18, HH12, Web11]. Adaptive [AFW10]. additionality [KDL19, ORB19]. address [Art12]. Addressing [HO19, Kee13, RG17]. Administrative [KdR17]. adoption [BG16]. ads [WLS12]. Advanced [Wes12]. advancement [ADR15]. Advancing [Xu15]. advantages [Lau11]. advent [Wes12]. adverse [Bor16]. advice [KCH19, MBHD12, POD19, Tho16]. advise [NdMSFC12]. Advising [Art12, Him18]. Advisory [CH17, Cro13, Kri15]. advocacy [Kin11]. aeronautics [dAFUFd19]. affair [Flo14]. affairs [FS10]. Affect [GAM^{+}19, Lv14]. affecting [LKP16]. affects [BAP19]. Africa [Boo11, CDD^{+}19, HWC^{+}18a, HWC^{+}18b, KG17b, TBvB15, TKM18, UJ14, Zav18, Sea13, Woo16]. African [Kah19]. after [IR17]. against [Men10]. Age [Lev15, Wic13a, ADM16, Ans14, Ren13, Tra14]. ageing [BCB18]. agencies [Bon14, LBMM13]. Agency [Gut18, Isk17]. agenda [BG19, MFM18, RVDWA18, VR18, dR18, dOB17]. agendas [PEV^{+}14]. agent [Har14]. Agglomeration [Les13, Sön14a]. Agreements [Hua17]. agri [HM10]. agri-food [HM10]. Agricultural [FSK11, Sea13, cC12, HGS11, LSKvA17, RHS12, TZ18]. agriculture [WHMS13]. agroecology [SG16]. ailing [TZ18]. AIMday [BLS16]. Aiming [vdBHC19, SBK^{+}19]. Air [Cal12, BSBD14]. aircraft [YHSH17]. Al [Art14]. Al-Dabbagh [Art14]. Alain [Egb14]. Albert [Fit18a]. Aldo [BK17]. Alexandra [Su14]. Aligning [Ric16, dR18]. alignment [CV12]. Alina [Kle12]. alleged [HBM19]. alliance [vdK13, Rei15]. alliances [dMFPP14]. allocation [CD16]. alternative [Wha19]. alternatives [Esp13]. altmetrics [RH19, RGvLR18]. ambiguous [Åm13]. ambition [Kro19a]. ambivalence [Mar11]. America [Gut16, Liv12, Arz10, Dia11, Dut10, DNA^{+}18, Kee13]. American [Su14, Doo18, DA10, Mat17, Wal15, Sha17]. Americans [Ans14, Lev15]. amidst [LKR19]. among [CFO16, GK10, HF17, Hua17, SFZ16]. Amr [Art14]. Anadon [Del16]. Analysing [AB19, KM13, KD10, YWH^{+}15, HNRM11]. Analysis [BW18, CPSL17, HH12, SK18, ABC^{+}18, Bee13, Blü18, COC14, CCR16, DSTS10, Dan17, DBD^{+}10, ET19, dMFPP14, HHJ10, HA11, HB16, JL19, Kam13, KdLABdP18, KJ18, LS15, LHGW14, Lor11, MBM19, Phi13, UG19, Wag15, WW17, WdJH14, WHKB12, WO17, YHR12, YH14, HLL14, FM16]. analysis* [PdSA13]. analytical [DRVWA18, RHJS15]. Analyzing [MSZ18, TFB12, LVG11]. Anarchism [Jac12]. Anderson [Alm16]. Andrea [Lus16]. Andrew [Ken16, Lus16, Mes14]. animal [SH14a]. Anique [Eva15]. Anna [Son14b]. Anneke [Lab14]. Ansolabehere [Lev15]. ante [Bul15]. Antecedents [ZLGG19].

Anthropocene [Rah15, Gal14, Rah15]. anticipating [AFW10]. Anticipatory [Con17]. Antitrust [Ors14]. Antonelli [Lig14]. Application [YWH^{+}15]. applications [LKP16]. applied [GK10]. appraisal [SSB19]. approach [AB19, ADS10, BLS16, BLKR17, BJS12, FS11, GP13, HFNW15, HOSV13, Kro19a, MAS13, NVAM14, NGB14, Tho16, Das15]. Approaches [Gib13, FS10]. appropriate [AdCBBB15]. Arabic [CAAH18]. Arabic-speaking [CAAH18]. Area [Kas10, Rya15, BHPW13, TS15]. Arenas [FT12]. Argentina [AV10, SBY18, TFB12]. argument [Nef19]. Argumentative [VG14]. Arm [Doo18]. ARRA [SH14b, SH14c]. art [GP13, WBWW15, Tei12]. articulation [tKBKS18]. artificial [SCE19]. Arts [Kle12]. Ashgate [Lov16]. Ashiabor [Su16]. Asia [ASSX16, DSTS10, LSZ16a, RHS12, SA16, ZLLU16]. Assessing [CZ19, Con17, HA11, Har14, Hig16]. Assessment [Bre12, HN15, RM14, BRB18, BSM14, Bou19, BL15, FvDvR16, Gru19, HFW^{+}12, Has11, HOSV13, Kas17, LS18, LS12b, LVG11, PD18, PH18a, PH18b, RRF^{+}19, SNT^{+}14, Tan17, WBWW15, Web11, WWL^{+}10, ZLRJ19, Bla14, Tan17]. Assessments [FTNdB14, WF13, vDFL14, vDH14]. asset [Art16]. asset-based [Art16]. Assimakopoulos [Sti17]. Assistant [HFW^{+}12]. assisted [BCB18]. associated [WSBTM15]. assumptions [Xu15]. Assurance [Lab14]. Astrid [Bey14, Grø14]. asymmetries [SD15, dBdLHB14]. Attaining [WS16]. attempts [Kas17, UJ14]. attitudes [SCB^{+}16, WCK^{+}14]. Attitudinal [SCB^{+}16]. Attracting [KG17a]. audiences [SCB^{+}16]. Audretch [Bra13]. Audretsch [Fit18a, Hoy13]. Australia [Bee13, Coo14, HFS^{+}13, LH13, TWGJ11, Van10]. Austria [SPH10]. Austrian [DT11]. authorities [tKBKS18]. authority [DS18a]. authorship [Kah18, MPP^{+}17]. awarding [VCdS16]. awards [BL10].

B [Bra13, Fit18a, Hoy13, Mar13]. Back [LDS19]. bad [BGY^{+}16]. balances [HBEvdV10]. Balancing [DLSS19, RVDWA18, SNT^{+}14, MTZ17]. Baldrige [LS12a]. Balkans [ORB19]. Baltic [Tõn19, TK17, Vīt15]. banking [MBHD12]. barrier [MHB14, Par10]. Barriers [FWS19, GAM^{+}19, BL16, BdCB14, FvDvR16, GST13, RH15, HN15]. Barry [Álv13, Kee13]. base [TS15]. Based [CPSL17, SK18, Art16, ADS10, BK19, BN15, Dan17, Das15, DMMPC18, DS18b, dMFPP14, FSHM19, Gal12, KB17, KVH13, Kro19a, LP16, MFC12, NGB14, RCRR19, Tan17, Tas13, TK11, YZ11, ZLRJ19, ZP17, FvDvR16]. bases [ZLLU16]. basic [GK10, Hem12, JL19, Ko15, RRMS12]. basis [MVFG15]. Basque [NVAM14]. Bayh [WS16]. be [Åm13, BdV14, MVFG15]. bedside [vdWVvdB12]. Beech [Pag11]. been [BFPCB10]. bees [PM19]. beginning [Flo14]. behavior [CGL13, HF17]. behaviors [HL11]. behaviour [Gök10, KKL17]. bench [vdWVvdB12]. benchmarking [DMRJ13]. Benedikt [Cad14]. Benefits [GAM^{+}19, Åm13, Arz10, DA10, DDT10, Dut10, FdSdS^{+}10]. Bengt [Tei14, BFE11]. Bengt-Åke [BFE11]. Berglund [Tei14]. Berlin [Ano13]. Bernhard [Man15]. best [SMR16]. Better [Fit13, Fit13]. Between [BG19, BUK16, KCH19, Tõn19, AWD16, AV10, BJB11, CPP^{+}13, DA10, DDT10, Dut10, Kah19, Klo13, KDL19, KB11, MB18, MK13, MRR17, OPMGCM14, OR10, RHJS15, SN13, SNL15, Tam13, TK11, UJ14, Yan18]. Beyer [Ren13]. Beyond [BJR17, BLKR17, CS14, LSTC17, Lov16, Tan13, Tas13, TYM^{+}14, AR11, Hub14, JC18, Sch15]. bibliometric [Aag15, PH18a, PH18b]. Bibliometrics [Sch15, CS14, DP13]. bibliometrics-informed [DP13]. Big [Joh17, JH12, MJH^{+}16, Hal15, Phi17, YPH17]. Bijker [Eva15]. bind [Tan18]. Bio [UJ14]. Bio-entrepreneurship [UJ14]. Biobank [HP12]. biobanking [MMGM12, Tup11]. Biodiversity [Ken17]. bioeconomy [Kea13, WHMS13]. bioenergy [Saa19]. Bioethics [Fra12, SH14a, Bro10, MBHD12]. biofuel [dMFPP14]. Biofuels [BSM14, TD12]. Biological [Mar13, FWS19, RJ14, WHMS13]. biology [MHB14, vDH14]. biomass [LP16]. biomedical [Aar17, AMLL18, LT11, WLA19, vdWVvdB12]. biomedicine [RdR17, SS10]. biopharmaceutical [Grø11a, Son17, Tuu11]. biosafety [KM13]. bioscience [LS15]. bioscientific [Fra12]. biosecurity [DSTS10, EG14]. biotechnologies [MBHD12]. Biotechnology [Fra12, SG16, AGS16, BFZFMT18, cC12, FSK11, Hei13, Hem12, HR12, HM10, Kin11, KKL17, vDTC19]. bird [EG14]. Birgitta [Tei14]. Bjerke [Mot12, Tei14]. Björn [Mot12, Tei14]. Black [Jac12, CLDS18]. Blaise [Sch15]. Blandine [Val14a]. block [RdMdAFY16]. blocks [Gök10]. blood [MBHD12]. bloom [Bou19]. blurring [SN13]. Bodies [Fra12]. Bolivia [PRZ^{+}10]. Booher [Afz13]. Book [Afz13, Ali15, Alm16, Álv13, And15, Art12, Art14, Bey14, BK17, Bra13, Cad14, Cal12, Cat12, Cla13a, Cla13b, Cla15, Das15, Del16, Egb13, Egb14, Esp13, Eva15, Fer16, Fit13, Fit18a, Fra12, Gib13, Gop14, Grø14, Gut16, Hal17, Haw12, Haw15, Hic12, Him16, Hol15, Hoy13, Hro14, Jac12, Kee13, Kee16, Ken16, Ken17, Kle12, Kor14, Kro19b, Lab14, Lem16, Les13, Lev15, Lig14, Liv12, Lov16, Lus16, Man15, Mar13, Mes14, Mil13, Mil15, Mot12, Mot16, Mus12, Ors14, Pri19, Rah15, Ren13, Rob14, Sch15, SL14, Sea13, Sha17, She15, Shi14, Sön14a, Son18, Sri14, Sti17, Su14, Su16, Tan13, Tei12, Tei14, Tra14, Val14a, Val14b, Wic13a, Woo16]. boom [SBD18]. Borders [TK17, Kor14]. Börje [Sön14a]. born [Wal15, vHWW19]. Borrás [Mus16]. Bosch [Mot16]. bottom [HOSV13]. bottom-up [HOSV13]. boundaries [RRH17]. Boundary [BK18, JRG19, POD19, SM10, SN13, Wha19, WWL^{+}10, dR18]. box [CLDS18]. Brain [AW10, Ciu10, MF17, Esc14]. Brazil [AG17b, BCSFC15, COC14, CRS17, FdSdS^{+}10, FSS18, PR10, TFB12, dMCL18]. Brazilian [CR14, CPP^{+}13, dMFPP14, dAFUFd19, KdLABdP18, SFBB10, TD12]. Breaking [Par10]. breeding [GG14]. bridge [Rah15, UJ14, WKGK11, Sen17]. Bridging [Kas17]. Briggle [Fra12]. Britain [Gut16]. Broadening [RG17]. broader [Kam13]. Brokering [SK18, TSK^{+}13]. brokers [SM13]. Brush [Das15]. Budget [KPG10]. build [Liv12]. Building [BK18, Gök10, Hel18, BFPCB10, Dag12, Gro11b, rL11, LF13, SFCR18a, SFCR18b, Egb13]. Bulgaria [Gal12]. Bunn [Del16]. Burden [Esp13]. Bureaucracy [FCLG16]. bureaucratization [KJ18]. Busch [BL13]. Business [AG17b, BL10, DLSS19, Grø11a, LTU10, LS18, Man15, Mot12, Sco14, DSZ14, MFC12, RPK16, SK11, Tas13, TZ18, Tuu11, UJ14]. businesses [Das17b].

C [Fra12]. Calestous [Sea13]. Cambodia [YWH^{+}15]. Cambridge [Gut16, Wil16]. campus [Cha16]. Can [BdV14, MMHV19, MVFG15, Rah15, TZ18, TK17, dC16, CAK12, LSTC17, Liv12, YHR12]. Canada [HFS^{+}13, KH12a, LH13, ML13, SL11, Tam17, WJOS14]. Canadian [Kee16, AL11, DCT18a, DCT18b, EJB11, IH11, Lem18, TK11, Wol14]. Cancer [SSB19]. Candeub [Tra14]. Candida [Das15]. Capabilities [CPSL17, JC18, Pre14, SK11, Les13]. capability [rL11, PJYL15]. capacities [Bor11]. Capacity [Hel18, SK18, Dag12, HL19, LT11, LF13, SFCR18a, SFCR18b]. capacity-building [LF13]. Capital [NJ18, ART10, DCT18a, DCT18b, Esp13, KA16, KWJ17, MF16, SBC16, YZ11]. Capitalism [vdK13, Kim14, Rei15]. Capitalist [Hlá17]. capitalists [LMDU16]. capture [Ric16, WCK^{+}14]. Carbon [Gib13, Lov16, ASSX16, Hub14, Liv12, Ric16, WCK^{+}14]. Care [Fit13]. career [ADR15, CAAH18, IG11, LS14, Waa17]. careers [LS15]. Caribbean [DNA^{+}18]. Carl [Cad14]. Carr [Cla13a]. carried [PR10]. Case [Bey14, DSP17, Lem16, RdMdAFY16, TK17, AAL19, AW10, ADLK12, BLS16, BFZFMT18, BJR17, BW18, Bro14, BAP19, CVCBA18, Cho19, CH17, cC12, DS11, FT12, FL13, dAFUFd19, GB19, GB15, HWC^{+}18a, HWC^{+}18b, HR12, HS16, JRG19, JNP16, JS18, Jul13, Kah19, Kam11, KBA17, KB11, KG17b, Kwo11, LDL^{+}10, LK17, MAS16, MMGM12, Men10, MBL16, NVAM14, PM18, PM19, PCS10, RB19, Rod10, RCRR19, SFBB10, SBC16, SMR16, SKY18a, SKY18b, SLvP10, SX19, SX20, SB11, SS17, Tam17, TBvB15, VJ12, VG14, Vil11, Vīt15, Web11, WBB11, WE18, Wu13, ZWL16, vdWVvdB12, vHDdJ17]. case-study [RB19]. cases [Gal12, HGS11, RHS12]. Cassidy [Sch15]. catalytic [MMHV19]. catch [CRS17, Cro13, LSTC17]. catch-up [CRS17, LSTC17]. catching [LZ10]. cautionary [Ekl11, Kah18, Löv11]. cautious [KCS12]. CCS [WCK^{+}14]. CCS-specific [WCK^{+}14]. celebrities [BRB18]. cell [BFPCB10, Flo14, Lev12, LLH13, Mar11, dJMA12, SS10, SFCR18a, SFCR18b]. cells [Dra14, Tho13]. Censorship [Mar15]. Center [Fit17, Kas19, YSRF19, YH14]. Central [LSL15, Rad11, SK10]. centrality [Eng18]. Centre [PH18a, PH18b]. Centres [Bor16, Hel18, TWGJ11, KG17a, SK11, Röd15]. Centrifugality [LGGK12]. Century [FM16, Swa17, AF14, Cap10, LHGW14, WHMS13]. chain [AFW10]. chains [ADLK12]. challenge [BJR17, KCM^{+}12, RdR17]. Challenges [ASvLW17, HMH^{+}15, IG11, KR18, LSL15, NdMSFC12, BJS12, BE18, CAK12, CM18, CH10, CBB12, Gie17, HJS17, Hem12, HO19, LF13, Mar16, Mer12, RM14, Ric16, SDP11, Tam17, WE18, WHMS13, DBD^{+}10, MF17, Son14b]. Challenging [OPBCM14, RVDWA18]. Chalmers [FL13]. Change [Gib13, Gol17, Him16, Kee13, Ken16, Sut17, Tan17, AGS16, BE14, BCB18, DLSS19, KA17, LU17, MTZ17, Mus16, Web11, Lig14, Mus16]. Changes [LMMD14, FL10]. Changing [BKKS15, Hro14, KdR17, Kwi12, SV15, HGS11, Tup11, WHKB12, ZP17]. Channels [Arz10, DA10, DDT10, Dut10, FdSdS^{+}10]. characteristics [CCR16, JC15, KCS17]. Characterizing [HFNW15]. Charis [Flo14]. charitable [KM19]. Charlie [Les13, Sön14a]. Charting [Bro10, Rya15]. Cheap [Ans14, Hol15, Sch14, Lev15]. chef [For10]. Cheltenham [Alm16, Sti17]. Chemical [Mar13]. Chemicals [Gei15, Aho17, Fer16]. Chicago [Gie17]. Chile [ÁC15, KG17a]. Chilean [TK11]. chimeras [SH14a]. China [ASC16, BLS12, HS16, Klo13, LG16, LT19, LKP16, LSW19, LZ10, LSTC17, LW19, Lv14, SX19, SX20, SFZ16, SFCR18a, SFCR18b, WLYCL15, WW17, ZWL16, vHWW19]. China-born [vHWW19]. Chinese [CL15, LDS19, LZ10, SBL15, YHSH17]. chip [RdMdAFY16]. choices [vR10]. Choreographing [Flo14]. Choreography [Flo14, Tho13]. Chris [Soe10]. Christopher [Lov16, Dos10]. circle [Uln15]. circles [Kas17]. cities [GC13]. Citizen [Cal12, GBH11, SMVdP14, BH11, DT11, DSB^{+}13, HA11]. City [Swa17, SX19, SX20, Wol14, GB15, Kee16]. city-regions [Wol14, Kee16]. civic [DO16]. claims [WM18]. classic [Hal11]. Classifying [KCS12]. Clean [Lev15, Ans14]. cleantech [KBA17]. Clement [Ken17]. clients [LSPC16]. Climate [Gho18, Gib13, Gol17, Kee13, Ken16, Lem16, Sut17, Tan17, AGS16, Bis19, Him18, Löv11, Web11]. Climate-resilient [Gho18]. clinical [Coo14, Rod10, SFCR18a, SFCR18b]. clinician [VG14, vHDdJ17]. clinician-developed [vHDdJ17]. clinician-scientists [VG14]. Cloning [She15, Jen14]. close [Mat10]. Closed [Eng18, Aar17]. Closing [TDPV13]. clothes [FCLG16]. clothing [Hal11]. clouds [BLS12]. clubs [Eng18]. cluster [Lan10a, MAS16, RHJS15, UJ14]. Clusters [Les13, MBD^{+}12, Sön14a, Son18, Wix15, CL15, Col10, Gie17, KD10, Tan13]. Co [DBO13, dMFPP14, Kah18, LKR19, Löv11, PD18, Ekl11, Kas10, Kwo11, MPP^{+}17, Mur13, PRZ^{+}10, TYM^{+}14]. Co-authorship [Kah18, MPP^{+}17]. co-creation [TYM^{+}14]. co-development [Mur13]. co-evolution [Kwo11]. co-existing [Ekl11]. Co-marketing [dMFPP14]. co-ordination [Kas10]. Co-produced [PD18]. Co-producing [LKR19, Löv11]. Co-production [DBO13, PRZ^{+}10]. coal [Gro10]. coalitions [Kin11]. Cognition [Wix11]. Cohesion [Son14b, Cas12, Coz12, Mer12, Ves12]. Colette [FM16]. Collaboration [FFF17, GAM^{+}19, SK18, WHLhC19, BGY^{+}16, Cla10, CAAH18, EBV17, FERV11, Gie17, Gra11, JMSS18a, JMSS18b, JCK14, Kah18, KB17, PJYL15, RHJS15, RRH17, Rya11, SRC17, Sch11, TG11, TWGJ11, TFE11, VR18, WBB11, WSBTM15, Yan18, vHWW19]. collaborations [HSM16, Hem17, HR19, IH11, OPMGCM14, SL11, Uln15, YH14]. Collaborative [Afz13, JC15, Yan18, CBB12, PCS13]. collective [KdLABdP18]. Collider [Hal17]. Collins [Lus16]. Colombo [Art12]. Combating [Gib13, AGS16]. combined [ADM16]. come [LMDU16]. commerce [Has11]. commercial [AG17a, MBHD12]. Commercialising [KBA17]. Commercialization [Hua17, Ben17, CCF16, CLDS18, KVH13, LH11, Lei11]. Commercializing [CS15]. Commission [Him18, Sco14]. commissions [CH17]. commitments [EJB11]. committees [Cro13, Kri15]. commodification [Sev13]. common [vR10]. Commons [Kor14, HP12]. Commonwealth [Rad11]. communicating [NdMSFC12]. Communication [STF17, JCK14, Pre14, SNL15, TM12]. Communities [Gib13, Dun10, RRF^{+}19]. community [AW10, DCJ19, Esc14, MPP^{+}17, VHAHGC12]. companie [Zav18]. companies [AL11]. company [Hem17]. Comparative [CPSL17, Blü18, HA11, HM10, LBS^{+}15, MBM19, PdSA13, TK11, WdJH14, ZLRJ19, ZLL17]. Comparing [ASC16, ASSX16, CHH18, Tsi12, ZLLU16, vHWW19, DA10, HFNW15, MF17, LDB^{+}12]. comparison [Ben17, MPC10, SW19]. comparisons [ASvLW17]. compatible [AGS19]. Competing [Ors14]. Competition [Bee13, Hua17]. competitions [CAAH18]. Competitiveness [Bra13, Shi14, Bon11]. complementarities [HBEvdV10]. complex [CGS19, KD10, MRR17, Smi14]. Complexity [Afz13, Lig14]. Compliance [Sha17]. Compositional [LSKvA17]. computing [RG17]. Conceiving [CETRW12]. concentration [Bee13, LSL15]. concentrations [GB15]. concept [Nie16b]. concepts [ABC^{+}12, GK10]. conceptual [Arz10, Sch11, Egb14]. Conceptualizing [HMH^{+}15, Mot12]. concern [EGI18]. Concise [Wad17]. conditions [Dra14, DNA^{+}18, EBV17, ZP17]. conference [Ano13]. conferences [DT11]. configurations [DCJ19, SvPLK14]. Conflicting [Kin11]. Conflicts [Bas11, AMLL18]. Congress [Bla14]. conjecture [Kim14]. Conjunction [McC13]. connected [MTR13]. connections [LT19]. Consequences [HP12, BK19, MBD^{+}12, TFE11]. considerations [Sch11, YHR12]. consortia [NO16, WdJH14]. constrain [Cro13]. Constraints [Hua17, KHZ^{+}13, ÁC15, CR10, TFB12]. construction [dMCL18]. consultation [dOB17]. contemporary [TFE11]. Content [Les13, AFW10, Phi13]. contentious [Lev12]. Contested [KM13, dJMA12, CC13]. Context [Hei18, Hen19, KCW14, Tho16, Vīt15, dOB17]. contexts [Åm13, EJB11, RRMS12]. contextualised [RGvLR18]. continuity [MTZ17, Uln15]. Contract [RRMS12, Kah19]. contribute [Kam11]. contribution [SvPLK14, Zav18]. contributions [Alm16, CM18]. control [LSL15, Pri19, BL13]. controls [BSBD14]. controversial [EJB11, SCB^{+}16]. controversies [BSBD14, KM13, Mar15]. controversy [Bou19, LKR19]. conundrum [LDL^{+}10]. conventional [Rad11]. Convergence [IR17, LSZ16a, SA16]. cooperating [SMR16]. cooperation [FLCRP19, HR19, KCM^{+}12, MFM18, McC13, Pre14, RPK16, HSM16]. Cooperative [TWGJ11]. Coordinated [TZI13]. Coordination [WBB11, BHPW13, LGGK12, NM15, Tam17, WdJH14, dR18]. Copenhagen [Gie17]. Coping [WHKB12]. Copyright [Kle12]. cord [MBHD12]. Core [Kam11]. Core-HRST [Kam11]. corporate [MPC10, SF19, Tam13, Mot16]. corporations [CPP^{+}13, CR10]. correlation [Tam13]. Corrigendum [Mil14, WH12a]. Cosmopolitan [Kor14]. cost [Liv12]. Costa [OR10, VHAHGC12]. costs [Lau11]. Could [Him16]. Council [Luu14, Fra12, Kro19b]. councils [CDD^{+}19]. countries [AW10, ADLK12, BCSF17a, BCSF17b, Ben17, CZ19, CHH18, CAAH18, DA10, GG14, Hag10, HN15, KCW14, OAE18, Rod10, Vīt15]. country [MBM19, SMR16, SW19, SNL15, NVAM14]. country-level [SNL15]. coupling [BOP18]. course [Bro10]. Courvisanos [SL14]. Cozzens [And15]. Creating [Grø14, SH14a, rL11, LBMM13, VMP10, Bra13]. Creation [DN18, Grø11a, JC15, MAS16, SRC17, SSTM12, TYM^{+}14]. creative [AFW10]. Creativity [Hro14]. Credibility [WWL^{+}10, HMH^{+}15, SNT^{+}14]. criminal [HFWL19]. Crises [SL14]. Crisis [IR17, FG16, KPG10, Kas13, KB11, MSZ18, Val14a]. Cristiano [Lig14]. Cristina [Art12, Cat12]. criteria [VCdS16]. criterion [Kam13]. Critical [CS10, CGS19, SNT^{+}14]. Cronin [Sch15]. Crops [Cla13b]. Cross [BW18, Gra11, TK17, TWGJ11, ASvLW17, CKS15, Cla10, FERV11, Kas17, MBM19, RHJS15, Rya11, SW19, Sch11, TFE11, Wu13, YH14]. Cross-case [BW18]. cross-country [MBM19, SW19]. cross-disciplinary [YH14]. cross-European [Kas17]. cross-industry [CKS15]. cross-national [ASvLW17, RHJS15]. Cross-sector [Gra11, TWGJ11, FERV11, Rya11, Sch11, TFE11]. cross-state [Wu13]. cross-university [Cla10]. crossroads [TWGJ11]. Crowding [DCT18a, DCT18b]. Crowding-out [DCT18a, DCT18b]. Crowdsourcing [MFC12]. Cuba [CP16]. Cultural [Fra12]. Culture [Tra14, CH10]. Cultures [EJB11, Eva15, KdR17, HMB14b]. curious [Wes12]. current [Elz12, FWRF14]. Currents [Hro14]. curves [Har13]. Custodians [DB10]. CWTS [PH18a, PH18b]. cybersecurity [CGS19]. Cycle [Son18, Tas13, TDPV13]. Cycles [SL14]. Czech [LS12b].

D [Art14]. Dabbagh [Art14]. Dachs [Man15]. Dalian [ZWL16]. Danish [AS16, KB11]. Dariusz [Mil13, Mil14]. data [BLKR17, CRS17, FG13, WCK^{+}14, Phi17, YPH17]. databases [HFW^{+}12]. David [Afz13, Bra13, Fit18a, Hoy13, Hro14, Kee16, Lem16, Lev15, Liv12, Ors14]. death [dMCL18]. Debate [She15, Blü18, ET19, Jen14]. debates [BB13, KCS17]. debris [McC13]. Debunking [Haw15, Maz14]. decarbonisation [LP16]. Decentralisation [KA17]. decentralization [BCSFC15]. decide [LMDU16]. Deciphering [Tan10]. decision [HHJ10, NdMSFC12, WWL^{+}10]. decision-makers [NdMSFC12]. decision-making [HHJ10]. Decisions [Fit13, BSBD14, Gro10, O’B13, SSC17]. Decline [Su14]. dedicated [CBB12]. Defence [SW19]. deficit [LTU10]. define [Mus16]. definition [Tam13]. degree [DS18a]. deinstitutionalization [Kwi12]. delegated [DS18a]. delegation [Dun10]. deliberation [DSB^{+}13]. deliver [Car18]. Delivering [Wix15]. Delphi [KH12a, VCLA15]. Delta [LKP16]. Delusion [RNN18]. Demand [BE18, MAS13, AS10, BN15, BCB18, Cho19, EB18, tKBKS18]. Demand-led [MAS13]. demand-oriented [EB18]. demand-side [Cho19]. demands [KB19]. Democracy [Rei15, BH11, Rya15, vdK13, Cla13a]. Democratic [BH11, CH17, Gru19]. Democratising [DP13]. Demortain [BL13]. dependence [DBO13]. deployment [PCS10]. depoldering [SLvP10]. depth [Vic14]. descriptions [Phi13]. descriptive [HDT17]. Design [CPSL17, AVL12, TDPV13]. Designing [DN18]. designs [CM18]. dessert [YZ11]. DESY [HH13]. Determinants [HF17, SSC17, ADR15, AB19, CP15, HFW11, KdLABdP18, KCW14, KWJ17]. determines [MBL16]. developed [MMHV19, MVFG15, vHDdJ17]. Developing [Gal12, LT11, AG17b, AW10, ADLK12, CZ19, Hag10, OAE18, PJYL15, Rod10, WS16]. Development [Ali15, Hal18, Haw12, Lab14, Mat17, SL14, SvTvL15, TK17, Val14a, AAL19, Dag12, Dia11, Dos12, EJB11, ET19, FWRF14, Gho18, GK10, Hem12, HBRUM13, Isk17, Jul13, LSKvA17, MAS13, MVFG15, Mur13, NZ16, Pea16, RB19, SZ11, SKY18a, SKY18b, SPH10, SNL15, TDPV13, Wu13, dBdLHB14, Egb13, Sön14a]. developmental [AGS19]. Developments [BB13, Lus16, CR14]. develops [Tan17]. Dhabi [MAS16]. Dhanaraj [And15]. Dialogue [DN18, DNA^{+}18, Han10, PE13, ABC^{+}18, DS18b, SBY18, Esc14]. dialogues [ÁJT18a, ÁJT18b, ABC^{+}12, SCH^{+}13, SV15, dMCL18]. Diary [Sha18]. Diaz [Del16]. dichotomy [BJR17]. Dickey [CETRW12]. Difference [Tei14]. differences [OAB10]. Different [CHH18, TSK^{+}13, Col10, HFNW15, HR12, JL19, JJ17, LSPC16]. Differential [CKS15, vdBHC19]. differentials [AWD16]. Differentiation [JJ17]. diffusion [CD16, LH11, PCS10, WKGK11, vHDdJ17]. Digital [Hag18, BLKR17]. dilemma [Kas19]. dimensional [HH13]. dimensions [DNA^{+}18]. Dimitri [Val14a]. Dimitris [Sti17]. diplomacy [FS10]. direct [Vic14, VCdS16]. direction [Haw17, tKBKS18]. Directionality [EB18]. Directions [Eva15, HMB14b]. Dirk [Son18]. disaster [DO16]. disciplinary [OMW13, OPBCM14, WSBTM15, YH14]. discipline [Rei15]. disciplines [HSM16, Mur13]. Disclosure [VRL13, CCF16]. Disco [Kor14]. discourse [AR11, LMMD14, PE13, Wag15, WW17]. discourses [KJ18]. discoveries [vHDdJ17]. discovery [Fra12, Jul13]. discursive [NK15]. discussion [Coo14]. disease [dOB17]. diseases [DSTS10, MVFG15]. dish [CHH18]. disparities [SNL15]. Dispatches [BGY^{+}16]. displacement [DCT18a, DCT18b]. Disrupting [Aho17]. Disseminating [DSB^{+}13]. distorted [ML13]. distribution [COC14, Wu13]. distributive [Cas12]. Divergence [IR17, SA16, LSZ16a]. Divergent [SFZ16]. diversification [MAS13]. Diversity [Egb14, OMW13]. divide [BLKR17]. divides [AS10]. Division [HLL14]. Do [AC11, Bey14, CS15, Lor11, Mes14, BFPCB10, Bes15, DRVWA18, Gou11, OAB10, Uln15]. doctorate [BCSFC15]. Doctors [Fit13]. documents [Blü18, Boh14]. Does [Kam11, Kas19, OPBCM16, BPSB16, Bes15, HL11, Lv14]. doesn’t [Son14b]. Doing [VMP10, Bro14]. Dole [WS16]. domains [Par10, vDFL14]. Domestic [HR19]. done [BFPCB10]. Don’t [Åm13]. Dosi [Ali15]. doubts [BSBD14]. down [Aag15]. downstream [Esc14]. downturns [Kim14]. drain [AW10, Ciu10]. driven [DBD^{+}10, LP16, LKP16, PCS10, Yan18]. driver [Lau11]. drivers [GC13, JCK14, vdBHC19]. Driving [Wix15, TDBS11]. Drone [Pri19]. Drug [BG16, Jul13, vHDdJ17]. drugs [CR14]. Dual [EG14, Mar13, EGI18, JS18, WJOS14]. Dual-use [EG14, EGI18, WJOS14]. during [FG16]. Dutch [Bre12, GvEV19, KdR17, SSB19, dOB17]. Duysters [Mot16]. Dynamic [LSW19, SC19a, SC19b, SvPLK14, vDFL14]. Dynamics [Ali15, Cal12, KHZ^{+}13, Swa17, BHPW13, Bee13, CVCBA18, Das17a, Dos12, Har11, KD10, LSKvA17, LBS^{+}15, SSTM12, UG19, YHR12].

e-businesses [Das17b]. E-commerce [Has11]. E-hailing [SX19, SX20]. Early [YPH17, KG17a, LS14, vDH14]. early-career [LS14]. earners [Kwi18]. Earth [Lem16]. ease [OPBCM16]. East [TBvB15]. Eastern [LSL15, Rad11, SK10]. eco [YHR12]. eco-innovation [YHR12]. ecogenomics [RBdCBB10]. Ecological [Lus16, Nef18, Val14a, Nef19]. Ecologist [Ogr17]. Ecology [Wic13a, Wic13a]. econometric [KdLABdP18]. Economic [BL10, Dos12, Egb13, Kee16, Ken17, Lig14, LC18, Sön14a, Sti17, Ano13, Ant18, CKS15, Das17b, Dia11, Gho18, JL19, KPG10, Kim14, LS12a, MAS13, SZ11, SKY18a, SKY18b, Wol14, Ali15]. Economics [Hic12, Das17b, Rei15, SRC17]. Economies [Kee16, CS15, CLDS18, Wol14]. Economists [Woo16]. Economy [Cat12, Hag18, Haw12, Lem18, RNN18, AG17b, BPSB16, Gal12, Joh11, SS10, Smi16, Tan13, WS16, Art12, BK17]. Ecosystem [Fit18b, Ken17, BAP19, KDL19, Fit18a]. Eda [Kor14]. Edited [Son18, And15, Art12, Bey14, Bra13, Del16, Egb14, Eva15, Fit13, FM16, Gib13, Gop14, Hoy13, Hro14, Kee16, Kor14, Lab14, Les13, Lig14, Man15, Mar13, Mil15, Mus16, Rob14, Sch15, Shi14, Sön14a, Son14b, Sri14, Sti17, Su16, Tan13, Tei12, Tei14, Tra14, Val14a, Val14b, Cat12]. editorial [ASSX16]. Edler [Mus16]. Education [BE15, DSP17, Hro14, O’B13, Sha18, Tra18, AB19, Kam11, Kwi12, SKY18a, SKY18b, SSTM12, Wil16, Zav18]. educational [Lor11, ZP17]. Edward [Alm16, Sti17]. effect [BG16, LBS^{+}15, Wha19, WS16, ZLGG19, vdBHC19]. effective [SR11, Van10]. Effectiveness [Rad19, HMH^{+}15, RG17]. effects [ADM16, ÁC15, CKS15, Cla10, FSS18, HGS11, HL11, JL19, MPP^{+}17, ORB19, VJ12, YWH^{+}15]. Efficiency [Cad14, AWD16, CZ19, LSW19]. effort [vdBHC19]. elected [O’B13]. electric [SFZ16]. electromobility [ASC16]. electronic [HFW^{+}12]. Elevator [POD19]. Elgar [Alm16, Sti17]. elite [Uln17]. Elizabeth [FM16]. embedded [JNP16]. Embedding [KH12b, SSB19, AdCBBB15]. Embraer [dAFUFd19]. embryo [CETRW12]. embryonic [LLH13]. Embryos [All10]. emergence [ASC16, FS11, Hen19, YSRF19]. Emerging [And15, BB13, Con17, FERV11, HO19, Mar13, STF17, Sti17, All10, ABC^{+}12, CS15, CLDS18, CT14, DSTS10, DS18b, FTNdB14, FG13, HWC^{+}18a, HWC^{+}18b, Lev12, RM14, SZ11, SCH^{+}13, Smi16, Sta13, TYM^{+}14, vDFL14]. Emperor [FCLG16]. empirical [Bee13, FvDvR16, GK10, RHS12, Sch11, TG11, BdV14]. Employment [BCSFC15, Waa17]. Empowering [SCE19, Cro13, DCJ19]. Empowerment [BW18]. enabled [GC13]. Enabling [FBB14]. encounter [Mus12]. End [Hag18, SL17, Flo14]. endeavor [Ano13]. endocrine [Aho17]. endocrine-disrupting [Aho17]. endorsement [Wha19]. enemies [Zha19]. Energy [Del16, Har13, Lev15, Liv12, Wag15, Ans14, ABC^{+}18, GP13, Isk17, Lan10b, TBvB15]. engaged [RRF^{+}19]. Engagement [GAM^{+}19, Aho17, ACBP16, Bes15, DP13, Esc14, Gro11b, JNP16, KA16, KCS17, LSBOPCM18, MRR17, OMW13, PE13, RMZP12, TWR11, TRGA16, TZ18]. Engaging [RBdCBB10, HLH13]. engendering [HLH13]. Engineering [BGW12, Lem16, Him18, RRF^{+}19]. engineers [HFW11, Wal15]. engines [Boo11]. enhance [dC16]. Enhancing [ADLK12, SBKP17, Gie17]. ensure [Åm13]. Entering [BRA12]. enterprises [BL16, Boo11, JL19, LSPC16, LKP16, LSW19, Nik15]. Entrenchments [SG16]. entrepreneur [Art14]. Entrepreneurial [Art14, Fit18a, Fit18b, Haw15, AAL19, FM16, FL13, KKL17, Maz14, ZLGG19, Mot12]. entrepreneuring [Mot12]. Entrepreneurs [Bey14, DS18b, OAE18, Tan18]. Entrepreneurship [Alm16, Bra13, Egb14, FM16, Hro14, Mil15, Sön14a, Tei14, Wad17, AMLL18, Das15, Das17a, Grø14, Hen19, Hig16, HKS14, Hug13, LHGW14, Men12, Mot12, NGB14, Tan10, TZI13, UJ14, Bey14, Grø14, Hoy13, Tei12]. Environment [Alm16, BKKS15, HGS11, Hug13, MTR13, SDP11, TD12, vDFL14]. environment-monitoring [SDP11]. Environmental [Hug13, JRG19, LH13, Rah15, Sha17, Sut17, CH10, Gal14, HFS^{+}13, KB19, RHJS15, RRH17, Alm16, Su16]. environments [GC13]. Envisioning [Fit13]. Epistemic [Dun10, KdR17, TS15]. epistemology [DO16]. equal [JJ17]. equality [SC19a, SC19b]. equation [AB19]. ERIC [LK17, She15]. Erratum [LSZ16b, SX20, SH14b]. errors [WM18]. espoused [LD15]. ESS [LK17, Agr12]. ESS-ERIC [LK17]. essays [Dos12, Ali15]. establish [Bis19]. Establishing [Art14, HN15]. Estonian [KKL17]. ethanol [Ekl11]. Ethic [MBHD12]. Ethical [Flo14, CETRW12, Rod10, Tho13, WF13]. ethically [Flo14]. Ethics [STF17, Sut17, All10]. ethnographic [LDL^{+}10]. EU [BFG^{+}15, DLV10, IR17, KPG10, LGGK12, LKR19, Lor11, MF17, MBHD12, MPC10, OAB10, PCS10, PCS13, Rya15, Son14b, ZLRJ19]. EU-27 [Lor11]. EU-funded [PCS13]. EU-US [OAB10]. EU27 [CP15]. Europe [ASSX16, AdC12, CR10, FG16, FL10, IG11, JS18, LSZ16a, LSL15, MBD^{+}12, MB12, MBM19, Rad11, SDP11, SA16, SV15, SK10, ZLLU16]. European [Cla13a, Kro19b, Son14b, BHPW13, Bon11, CAK12, Dun10, FvDvR16, FBA19, Hal15, HN15, Him18, KPG10, Kas17, Kas10, KD10, Kwi18, LTU10, LK17, LKR19, Löv11, Luu14, MBHD12, RP17, Ric16, RNN18, Rya15, TS15, Tsi12, Uln17]. Europeanization [SK10]. evaluate [Kro19a, YH14]. evaluating [HMH^{+}15, RdR17, Han10]. Evaluation [KdR17, Lab14, AGS19, ADS10, BH11, Bul15, CVCBA18, DP13, LSW19, LS12a, NSS19, TDPV13, WM18]. everything [Art16]. Evidence [ADR15, ADK16, AG17b, ÁC15, CGS19, FG16, Hem17, KG17b, LT19, MB18, MRR17, PJYL15, Por16, RP17, RHS12, SRC17, SSC17, Smi16, Val14b, vDTC19, AF14, BdV14, DMMPC18, DDT10, FdSdS^{+}10, LP16, MFC12, OAE18, Son17, Tan17, TDBS11, YHSH17, Lv14, SR11, ZP17, vDFL14]. evidence-based [DMMPC18, ZP17]. Evolution [EBV17, NJ18, AAL19, GB15, Kwo11, TBvB15]. Evolutionary [ART10, NZ16]. evolving [DLSS19, HO19, Kin11, PJYL15]. Ex [Bul15, TFB12]. Ex-ante [Bul15]. ex/inclusion [TFB12]. Examination [AS16, CGL13]. examining [AG17a, LTU10]. example [GvEV19]. Excellence [Hel18, LBS^{+}15, LS12a, Son14b, BG19, Bor16, CVCBA18, CHH18, FPVH18, JJ17, KG17a, TKM18, WKGK11]. excellent [JS18]. exception [MF16]. exchange [CC13, HL11, LSBOPCM18, ZLL17]. Executive [LH13]. exercise [BRB18, Bou19]. exist [LMDU16]. existing [Ekl11]. Expansion [Zav18, HH15]. expectations [LU17]. Experience [Mat17, BSHG16, BCSFC15, LZ10, MP12, PRZ^{+}10, Rya11]. Experiences [ACBP16, BW18, BGY^{+}16, KG17a, NZ16, SKY18a, SKY18b]. experiment [Tas14, Lem16]. experimental [GK10, HH15, Tas14]. Experiments [Hol15, LF13, Sch14]. Expert [Bou19, FWRF14, KJ18, LKR19, POD19]. expertise [All10, CH17, Kri15, Par10, SV15]. experts [SCB^{+}16]. explain [OAB10]. Explaining [HM10, MF16, Mus16, BE14, VJ12]. exploration [Bul15, Dog13, GST13, dC17]. exploratory [COC14, HR12]. Exploring [AWD16, CRS17, DS18a, LDL^{+}10, Tan13, TYM^{+}14, HA11, HBEvdV10, Rod10, HL11, Kas19, MJH^{+}16, MVFG15]. extension [MB18]. External [JNP16, HF17, TRGA16]. extra [Phi13]. extra-university [Phi13]. extramural [SH14b, SH14c]. extraordinary [Haw17]. Extremes [Ken16].

F [Fra12]. face [RPK16]. faced [BdCB14]. facilitate [PEV^{+}14]. Facilitators [BK18, BdCB14, FWS19]. Facing [KSC^{+}12, Rad11]. Factors [IR17, LDS19, PEV^{+}14, CS10, FBS16, GRLR17, HFWL19, LS14, LSBOPCM18, TDBS11, WSBTM15, LKP16, OPBCM16]. Factual [WM18, LDB^{+}12]. Faculties [KdR17]. failure [Art12, vHDdJ17]. Failures [Sha17, MF16]. Faïz [Gop14]. Falck [Hoy13]. Fall [Hal17, LMMD14]. familiarity [LDB^{+}12]. Family [FBB14]. Family-friendly [FBB14]. Farewell [Pag11]. Farnham [Lov16]. fast [WHKB12]. fast-changing [WHKB12]. favourable [RRMS12]. Fayolle [Egb14]. Features [Elz12]. Federal [Sco14, SL11, Sen17, Wu13, BKKS15]. Federica [BK17]. feedback [Uln15]. feet [BLS12]. Fernandez [Cat12]. field [FM16, HH15, Lev12, VMP10, WBB11, WLA19]. financial [ÁC15, FG16, Kas13]. Financing [BL16, Cad14, CR10, Wu13]. Findings [BFG^{+}15, TDPV13]. Finland [Gal12, Hen19, RPK16]. Finnish [NK15, TS15]. Firm [SK18, ADS10, FLCRP19, SSTM12, SK11, Wix11, Xu15]. Firms [Val14b, AL11, BN15, Dut10, FdSdS^{+}10, HL19, HR12, KVH13, rL11, NSS19, OR10, OAB10, PR10, SFZ16, SSC17, TDBS11, ZLLU16]. first [AF14, UJ14, WHMS13]. FIS [RRMS12]. FIS/Miguel [RRMS12]. Fiscal [Su16, CP16]. fisheries [Cro13, SvTvL15]. fit [BPSB16, Son14b]. five [GG14]. flows [Har11, Son17]. flu [EG14]. Flynn [Lus16]. Flyverbom [Mus12]. fodder [ADLK12]. Follow [Fit17]. Food [Cla13a, HM10]. foods [Cla13c]. Footprint [Lus16]. forces [BUK16]. forecast [SG16]. foreign [FS10, HR19, KG17a, LSL15, SMR16, Wal15]. foreign-born [Wal15]. foresight [AFW10, CS10, KH12b, NdMSFC12]. forestry [RB19]. form [SCB^{+}16]. formal [TG11]. Formation [HH15, OL13, LS14, NK15, Van10]. forms [FERV11, Rya15]. formulating [VHAHGC12]. Fornahl [Son18]. forward [AVL12]. forward-looking [AVL12]. Fostering [HNRM11, Vil11, CL15]. Foundation [HLL14, Kam13]. Foundations [Tan13]. four [BCSF17a, BCSF17b, DA10, vDFL14]. fracking [PM18]. Fragile [Rei15, WO17, vdK13]. frame [RBdCBB10]. framework [Arz10, DRVWA18, HNRM11, SC19a, SC19b, SvPLK14, Sta13, WH12a, WH12b, OL13]. frameworks [Coo14, Ekl11, YHR12]. Framing [Agr12, Cla13c, HLI^{+}11, Bou19]. framings [KM13]. France [ASC16, Bou19, PD18]. Fraunhofer [CBB12]. Fred [Haw12, Rob14, Sri14]. free [Mar15]. Freeman [Dos10, Soe10]. French [MF16]. Frey [Cad14]. Frieder [Lab14]. friendly [FBB14]. front [BGY^{+}16]. Frontier [BKM10]. FTA [BJS12, CAK12]. Fudge [Gib13]. fuelled [SSTM12]. Fukushima [Hub14, Lov16]. full [ADM16]. fun [Ste10]. function [TYM^{+}14]. Functional [TBvB15, Tan10]. Functions [Ken17]. funded [Bul15, PCS13, SH14b, SH14c]. Funding [BKKS15, CAAH18, Fit17, AS16, BL10, BCSF17a, BCSF17b, BS15, Bor16, CP15, GRLR17, Kim14, KM19, Luu14, LBMM13, MPP^{+}17, SBK^{+}19, SPH10, TZ18, Tõn19, Wha19, Wu13, ZLRJ19]. funds [SH14b, SH14c]. funnel [FLCRP19]. fusion [PCS13]. Future [DSTS10, Bas11, CBB12, DBD^{+}10, HHJ10, Jul12, KSC^{+}12, Mic11, RBdCBB10, VR18, WHKB12, Wil16]. future-oriented [HHJ10, WHKB12]. futures [CCR16].

G [Art12, Das15, Kee13, Les13, Sti17]. Galaz [Rah15]. Gallouj [Gop14]. gamete [CETRW12]. gap [AV18, DLV10, OAB10, WKGK11, Gal14, Rah15, Rah15]. gaps [Kas17, SCB^{+}16]. Gas [Ben13, Wag15]. Gatewood [FM16]. gathering [LP16]. Gault [Haw12, Rob14, Sri14]. geeks [AG17a]. Geert [Mot16]. Geiser [Fer16]. Gender [CFO16, Fra12, WLA19, AV18, SC19a, SC19b, SNL15, Nie16a]. gene [Hem12]. generalizability [SX19, SX20]. generate [DRVWA18]. Generating [RRMS12]. Generation [KR18, BJR17, EB18, Kwi18]. Generic [Tam13, CR14]. Genetically [ML13, Cla13c, VMP10]. genetics [LDL^{+}10]. genomics [Hei13, HM10]. Geoengineering [Him16, Lem16]. geographic [Lev12]. geography [RH15, Son17, TS15]. Georg [Man15]. Gerd [Fit13]. German [Phi13, Sch11]. Germany [ASC16, Ano13, ART10, BL16, HH12, JJ17, ZP17, vDH14]. Get [Woo16]. Getting [CC13]. Geuna [BK17]. Ghana [AAL19]. Ghetto [Nel11]. gift [Tup11]. Gina [Esp13]. Giovanni [Ali15]. Girgerenzer [Fit13]. glimpse [CLDS18]. Global [Cla13b, DSB^{+}13, Gal14, Haw12, HS16, Lev15, Mus12, She15, WHMS13, AAL19, Ans14, BSBD14, BGW12, Coz12, CBB12, Das17a, HBEvdV10, Jen14, KCM^{+}12, Mis11, Ros10, SDP11, Rah15]. globalisation [ZLLU16]. Globalization [Dan17]. globally [MTR13]. GM [Cla13a]. goals [AGS19, Dun10]. goes [Ros10]. Good [Hol15, Tho13, BGY^{+}16, Sch14, Flo14]. goods [SDP11]. Göran [Cal12]. governable [Web11]. Governance [Con17, DSP17, Kee13, Mus16, OL13, Rah15, Rya15, Shi15, Åm13, AF14, Art11, BE14, Cla13c, DT11, EJB11, Gal14, GST13, HA11, HNRM11, HMB14b, KJ18, KCH19, KD10, Lar12, LDL^{+}10, MSZ18, MMGM12, MBHD12, Nie16b, SH14a, SBL15, SFCR18a, SFCR18b, Yan18, Eva15, Val14b]. Governing [BAP19, BCB18, FG13, HLH13, dJMA12, Rod10, SS10, WJOS14, Bru10, Cal12]. Government [Kim14, Wha19, Art12, Boh14, CS10, Cla10, Has11, HFS^{+}13, HO19, IH11, LH13, NO16, PJYL15, SMVdP14, SF19, Sen17, Tam17, WS16, dOB17, Art14]. government-led [CS10]. government-sponsored [NO16]. government-supported [PJYL15]. governmental [OL13, Van10]. governments [Car18, Kri15]. Governpreneurship [Art14]. Grace [Ren13]. gradual [HH12]. graduate [GRLR17]. graduates [Waa17]. Grand [KR18, BJS12, CAK12, CM18, KCM^{+}12, Ric16, MF17]. grant [SSC17]. granting [CDD^{+}19]. gravitational [FSS18]. Gray [Fit13]. greater [DP13]. Greece [PCS10]. Greek [Kas13]. Green [Su16, Bou19, HM10]. Greene [Das15]. Greenhouse [Kee13]. greybeards [Bru11a]. Griffiths [Ken16]. Grilli [Art12]. ground [BLS12]. Group [MBHD12, Van10]. groups [HR19, KKL17]. Growth [Wix15, Dun10, JL19, LBS^{+}15, Men10, Tas13, Xu15, YZ11, Bra13]. Guest [ASSX16]. Guide [Mes14, Wad17, Kee18]. Gusterson [Pri19].

H5N1 [EG14]. Haber [Wil16]. Hail [For10]. hailing [SX19, SX20]. Hallmarks [Van10]. Handbook [Lai17, Hoy13, Lig14, Rob14, Sri14]. Hands [BW18]. Hans [Mot12]. hard [Rob14]. hardback [Gut16, Sti17]. hardcover [Alm16, Lov16]. hardware [Pea16]. Harm [Fer16, Gei15]. Harnessing [Sch15, CS14]. Hart [Liv12]. Hartmut [Mil15]. Harvest [Sea13]. hazard [Bor16]. Head [BLS12]. Health [ACBP16, Fit13, Fit17, SH14b, PM19, Smi14, WBB11, WHMS13, vdWVvdB12, LS18, SH14c]. heartland [JNP16]. Heblich [Hoy13]. Heidemann [Bey14, Grø14]. Heidi [Das15]. Helix [FT12, LC18, JMSS18a, JMSS18b, NZ16, RHJS15, RRH17, Bre12, Vil11, YHR12]. hello [Pag11]. help [LSTC17, WE18]. Henry [FM16]. heritage [Tõn19]. heterogeneity [AF14]. Heterogeneous [ÁC15, BUK16]. High [Tuu11, ZWL16, OAE18]. high-quality [OAE18]. High-tech [Tuu11]. Higher [DSP17, FM15, Sha18, Tra18, AB19, Kwi12, SSTM12, Wil16, Zav18]. Higher-level [FM15]. hindering [dBdLHB14]. Hirsch [Mil15]. Hirsch-Kreinsen [Mil15]. Hisrich [Art14]. Historical [AAL19, Elz12]. history [CVCBA18, Day19, PH18a, PH18b]. hits [RGvLR18]. holders [BCSFC15]. holistic [NVAM14]. Holloway [Fra12]. Hommels [Eva15]. honour [BFE11]. Hope [Su16]. hopes [Tuu11]. Hopper [Ren13]. horizon [ABC^{+}12, KSC^{+}12, vR10, Eng18]. Horizons [Swa17, HLH13]. hormone [Dun10]. hospitals [BdV14, RRMS12]. House [Har14]. HRST [Kam11]. Hubbard [Lov16]. Huggins [Alm16]. Hugh [Pri19]. Human [KWJ17, Flo14, LLH13, MF16, MF17, SH14a, SBC16, YH14]. human/animal [SH14a]. humanities [CVCBA18, Kas10, OPMGCM14]. hybrid [Kri15]. hydrocarbons [LKR19]. hype [Tan13]. Hypothesis [Hol15, Sch14].

Ian [Cat12]. ICT

[LTU10, Dia11, HR19, LTU10]. idea

[Day19, Isk17]. ideas

[Ko15, Sch14, Smi14, Hol15]. identification [ZLGG19]. identified [BdV14]. identify [MSZ18]. Identifying [Pre14, VJ12, vR10]. identities [CS15]. identity [RH19]. if [Tei14]. Ignorance [Gro10]. IIT [Kam11]. Ilan [Sti17]. illusion [Mer12]. Illustrations [NM15, Sch11]. imagery [WO17]. Imagination [Mot12]. Imaginative [Gib13]. Imaging [Esc14]. immersive [WWL^{+}10]. Impact [ADS10, Hel18, AV18, BG19, BFZFMT18, BCSF17a, BCSF17b, CS14, Cro13, DS18a, Dia11, ET19, GC13, GST13, GBH11, HFW^{+}12, HOSV13, Hem12, HDT17, HFWL19, JCK14, JC15, JS18, Kam13, KVH13, LVG11, MBL16, PCS10, RVDWA18, RGvLR18, SF19, Tan18, TG11, VMP10, WBWW15, Wal15, WHLhC19, Wil16, WF13, dJSvD16, vdBHC19, Mes14, Sch15, Sti17, Su16]. Impacts [CR14, HHJ10, BL15, DRVWA18, HA11, Kam13, Mic11, Vic14]. impede [Kas19, PEV^{+}14]. Implementation [Haw12, Lei11, PEV^{+}14, STF17]. implementing [HJS17]. Implications [Con17, BL10, BS15, HHJ10, LW19, RHS12, SCH^{+}13]. importance [HL19, JMSS18a, JMSS18b]. important [Wil16]. improved [Mar11]. Improving [CH10, KH12a, VJ12]. in-depth [Vic14]. inadvertently [Nef19]. incentive [CP16]. Incentives [FvDvR16, Tas14, Aag15, BK19, CAAH18, Nef18, RCRR19, SD15]. Incentivizing [KB17]. Inclusion [Mat17, Mer12, TFB12, Ves12]. inclusive [HWC^{+}18a, HWC^{+}18b, dBdLHB14]. Income [JL19, LSTC17]. Income-level [JL19]. Inconsistencies [DSP17]. incorporation [RRMS12]. Increasing [TDPV13]. incubators [AG17b, SR11]. Incumbent [Cho19]. Independent [Rad11]. Index [Ano10a]. India [ASC16, BB13, DS11, FSK11, Kam11, LT11]. indicator [Aag15]. Indicators [Ano13, BRA12, Rob14, Sch15, Sri14, ASvLW17, CS14, FPVH18, SSK15, SNL15]. indigenous [YZ11]. indirect [Vic14]. Individual [ADK16, KCW14, LSBOPCM18, HFW11]. Industrial [Ali15, Cla10, Col10, Dos12, DMMPC18, FL10, Har11, LSW19, MTZ17, Mur13]. industries [AFW10, DBdPR18, DLV10, HKS14, Esp13, Mil15]. Industry [GAM^{+}19, Aho17, AV10, Bee13, BSBD14, CR14, CKS15, Dan17, Dog13, DA10, DDT10, Dut10, FdSdS^{+}10, dAFUFd19, FBS16, Grø11a, JMSS18a, JMSS18b, MFM18, MHB14, Rad11, SL11, Smi16, Son17, TG11, TDBS11, Tuu11, VR18, Vil11, WS16, YZ11, Zav18, dC17]. inequalities [Smi14]. Inequality [And15, BK19, CT14, Rya10]. inequity [GB19]. infectious [DSTS10]. infernal [Esp13]. influence [JC15, ML13, RdMdAFY16, Waa17]. Influences [GRLR17]. Influencing [tKBKS18, Kin11, LDS19]. Informal [OPMGCM14, JCK14, KG17b]. Informatics [ACBP16]. Information [DBdPR18, Lai17, Mis11, Ren13, STF17, dBdLHB14, BRB18, Bon11, HL11, Pre14]. informed [DP13]. infrastructure [CGS19, LK17]. infrastructures [Aar17, MMGM12, Mis11, LK17]. Ingar [Tan13]. inherently [Gru19]. initiating [ABC^{+}12]. initiatives [Has11, JJ17, Rya11]. Innes [Afz13]. innovate [MJH^{+}16]. Innovating [Hol10, LU17, Wol14, Kee16]. Innovation [AVL12, BL10, Boe18, Bra13, Cad14, Cat12, CPSL17, CT14, Dia11, DBdPR18, Egb13, Gop14, Hag10, Hoy13, IR17, Klo13, KHZ^{+}13, KR18, Lem18, Lem16, LS18, Liv12, Mat17, NJ18, PE13, RCRR19, SDP11, SA16, SL14, Sea13, Sri14, Tan13, Tas14, TK17, UG19, Val14a, Val14b, Vīt15, Wad17, Yaq12, ZI10, Zha19, ASSX16, ÁC15, ÁJT18a, ÁJT18b, AS10, Art11, Art12, ADS10, ADLK12, BKM10, BL16, BRA12, BG16, Bon14, BE18, Boo11, BFE11, Bor11, BE15, BL15, Bru10, CAK12, CL15, CR14, CP16, Cha16, CM18, CKS15, CZ19, cC12, Cla15, Dan17, DS11, Day19, DBD^{+}10, DMMPC18, EB18, FS11, FERV11, FG16, FBA19, FT12, FSK11, FSHM19, FT17, Gal12, Gau12, Gra11, HFNW15, HWC^{+}18a, HWC^{+}18b, Har11, Haw17, Hen19]. innovation [HBEvdV10, Hig16, HNRM11, HO19, HR19, HP12, JMSS18a, JMSS18b, JH12, JC18, JS18, Jul13, KJ18, KVH13, KDL19, Kro19a, KG17b, LSKvA17, LT11, Lar12, Lau11, rL11, LMMD14, LSZ16a, LSL15, LG16, LSW19, LZ10, LSTC17, Lor11, Lv14, MAS13, Mar16, MTR13, MBM19, MHB14, MJH^{+}16, NSS19, Nie16b, NK15, NM15, ORB19, Ors14, OMS12, PCS10, Rad19, RHS12, Ric16, Rob14, RGFLRP^{+}18, Rol13, Rya10, SS10, SMR16, SvPLK14, Sco14, Sen14, Sen17, SX19, SX20, SSTM12, Son14b, Sta13, STF17, SBY18, SK10, Tam17, Tam13, Tan10, TDPV13, TK11, TBvB15, TZI13, Tsi12, VHAHGC12, VG14, Wag15, WW17, WE18, WH12a, WH12b, YHR12, YHSH17, YWH^{+}15, dBdLHB14, tKBKS18, MFC12, Del16, Mot16, Shi14, And15, Haw12, Mar13]. Innovations [RMZP12, HE11, vHDdJ17]. Innovative [Esp13, Art11]. Innovator [Hol15, Sch14]. innovators [KCS12]. input [ASvLW17, CH10]. inquiry [Esc14]. Insights [BdCB14, LH13, PH18a, SM13, MPC10, PH18b]. inspired [Arz10]. Institute [Lin19]. Institutes [LS18, SH14b, SH14c, CPP^{+}13, PJYL15, Phi13, PR10]. institution [DBO13]. Institutional [Coo14, HH12, LC18, Rei15, Wes12, AMLL18, BUK16, GRLR17, HGS11, KA17, KB11, MF16, SFZ16, SPH10, Tõn19]. Institutionalization [TG11]. Institutionalized [GB19]. Institutions [Egb13, Sch11, Åm13, FWS19, Rol13, SSTM12, Cla15]. instrument [AVL12, EB18]. instrumental [BSBD14]. instrumentality [BJR17]. instruments [SL11, WH12a, WH12b]. insurance [FG16]. Integrated [Ben13]. Integrating [BN15, HFW^{+}12, WF13, FBA19]. Integration [FL13, LGGK12, RM14, BHPW13, BSM14, FM15, OMW13, SV15, TS15]. intellectual [AR11, BGW12, CRS17, CKS15, Dos10, Hem12, Nik15, Tam13, Cad14]. intelligence [SCE19]. Intended [KHZ^{+}13]. intensity [OAB10]. Intensive [Bey14, Cat12, CL15, DBdPR18, FSS18, GBH11, Grø14, HKS14, OAB10, Mil15]. Interaction [Bro14, CPP^{+}13, DDT10, Lau11, TG11, VCLA15, Cal12]. Interactions [AV10, Dut10, AG17a, Arz10, DCJ19, DA10, OR10, Vil11]. interactive [CDI16]. interdependence [KDL19]. Interdisciplinarity [DSP17, Hei13, OMW13]. Interdisciplinary [RH15, EBV17, HBRUM13, LF13, LBMM13, Mic11, WSBTM15, dR18]. interest [AMLL18, Bas11]. interface [BJB11, Saa19]. interfaces [HMH^{+}15, SNT^{+}14, Tan13]. Intermediaries [MK13, BSHG16, HDT17, LSPC16, RCRR19, WW17]. intermediary [Åm13]. intermediate [SMR16]. intermediating [KBA17]. intermediation [Dog13, SM13]. International [BFG^{+}15, DMRJ13, FM16, Lan10a, LS15, SRC17, CAAH18, FWS19, HSM16, IH11, JCK14, JCC10, KCM^{+}12, KG17a, LHGW14, LT19, MB18, McC13, Pre14, Uln15, WHLhC19, dC17]. internationalisation [DSZ14, JCC10, Tõn19]. Internationalization [Man15]. Internet [Tra14, Das17b, Mus12, Sev13, Mus12]. interplay [Smi14]. interpretation [ART10]. intersection [FS10]. intra [HSM16]. intra- [HSM16]. intricate [Flo14]. Introduction [Afz13, BFE11, BL13, CDI16, Dut10, EB18, FG13, HHJ10, LH11, Lin14, MK13, MPC10, RVDWA18, RM14, SZ11, TFE11, Ves12, CP16, GBH11]. invented [AC11, CD16]. inventions [KBA17]. Inventor [Hua17, Son17]. inventors [ADK16, CD16, LS14]. Investigating [WLA19]. investigation [CKS15, GK10, LD15, SF19]. investment [BSBD14, DLV10, Pea16]. investments [CR10, FSS18, LF13]. Invisible [Hua17]. involvement [DBD^{+}10, FWRF14, KCS17, SSB19]. Involving [DS18b, PEV^{+}14]. Iran [KM13].

Ireland [Hen19, ZLL17]. Irish [Rya11]. Isabel [Mil15]. ISBN [Alm16, Gut16, Lov16, Sti17, Wil16]. Islamic [HSM16]. Island [Sen17]. isomorphic [BUK16]. Israel [ART10]. issue [BFE11, CDI16, Dut10, GBH11, LH11, MPC10, SZ11, Tho16, TFE11]. issues [ABC^{+}12, EG14, FG13, ML13, MSZ18, SDP11, YSRF19, Ken17]. Italian [AWD16, AF14, CD16, DSP17, VRL13]. Italy [ADR15, MRR17, SR11, VCLA15].

J [Fit13, FM16, Ors14]. Jack [Lem16]. Jackson [Gib13]. Jakob [Mus16]. Janet [Su16]. Japan [LS15, SB11]. Jemelniak [Mil13]. Jemielniak’ [Mil14]. Jensen [She15]. Jerry [SL14]. Jerven [Woo16]. Jessica [Eva15, Mot16]. Job [WLS12]. Johannisson [Tei14]. Johansson [Sön14a]. John [FM16]. Joint [vR10, JCC10]. Jonathan [Mar13, Wil16]. Jorge [Egb13]. journal [HFWL19]. Journalism [She15, Jen14]. journey [LS12b]. Judith [Afz13]. Juma [Sea13]. just [Yaq12]. Justice [Sut17].

Kaderabkova [Son14b]. Kaleckian [SL14]. Karin [Tei14]. Karla [Fra12]. Karlsson [Les13, Sön14a, Tei14]. Kass [Fra12]. Kate [FM16]. Kazakhstan [Smi16]. Kazuhiro [Su16]. Keith [Lem16]. Ken [Fer16]. Kenya [Kin11, PRZ^{+}10]. key [Lv14, SBK^{+}19]. Killewald [Su14]. Kinchy [Cla13b]. Klundert [Rei15]. know [BFPCB10]. Knowledge [Ant18, Art12, Bey14, Cat12, GST13, Hem17, HKS14, KdR17, Mil13, Mil14, NK15, NO16, Tra18, AF14, AS10, Arz10, BG19, BFZFMT18, BJR17, BPSB16, BOP18, CL15, CVCBA18, CD16, Cas12, CRS17, DS18b, FSS18, FSHM19, Gal12, GBH11, Grø14, HFW^{+}12, HFNW15, Har11, HB16, HBRUM13, JC15, KA16, KB19, KB11, Kwo11, LDB^{+}12, LDL^{+}10, LH11, LKR19, LSBOPCM18, Löv11, LBMM13, NZ16, OMW13, OPBCM16, PRZ^{+}10, RB19, RJ14, RGFLRP^{+}18, SRC17, SZ11, Saa19, SS10, SM13, Smi16, Son17, SvTvL15, SD15, Tan13, TS15, TSK^{+}13, Ves12, VR18, Wag15, WHMS13, WKGK11, Yan18, YZ11, ZLL17, ZLLU16, vR10, vdBHC19, Mil15]. knowledge-based [DS18b, FSHM19, Gal12, YZ11]. knowledge-driven [Yan18]. Knowledge-intensive [HKS14, CL15, FSS18, GBH11, Grø14, Mil15]. knowledge-transfer [Kwo11]. Knut [Tan13]. König [Kro19b]. Konisky [Lev15]. Korea [CRS17, Hem17, Ko15, Kwo11, rL11, OL13, YH14]. Korean [PJYL15]. Kranakis [Kor14]. Kreinsen [Mil15]. Kreiser [Su16]. Krsto [Sti17]. Kurt [Ren13]. Kyrö [Egb14].

label [vHDdJ17]. Labor [Esp13]. Laboratories [Ben13, HH12, JCC10, Lv14]. laboratory [Hal11]. Labour [Joh17, Esp13, Lor11]. labs [KdLABdP18, Wes12, SBKP17]. ladder [LVG11]. lagging [NSS19]. Lamastra [Art12]. landscape [BKM10, FTNdB14, Luu14, ML13]. Lanka [AW10]. Laperche [Val14a]. Large [Ben13, Fit17, BRB18, CR10, LF13]. large-scale [LF13]. Larry [Su16]. Lassen [Bey14, Grø14]. last [Uln15]. latecomer [TDBS11]. Latin [Arz10, Dia11, DA10, Dut10, DNA^{+}18, Mat17]. Latvia [HLI^{+}11]. Laura [Alm16, Del16]. Law [KdR17, Kle12, SL17, CGL13, HLH13, Doo18]. lay [SCB^{+}16]. leaders [vdWVvdB12]. leadership [HS16]. leading [rL11, SFZ16]. learn [VMP10]. learned [HDT17]. Learning [Bla14, Tan13, AS10, BFE11, Bor11, Bro14, Han10, Har13, Joh11, Kin11, Klo13, Ko15, KD10, MTZ17, Wil16]. led [CS10, MAS13, SvTvL15]. Lederer [Hoy13]. Lee [Álv13, Su16]. legacy [Bla14]. Legal [Hem12]. legitimacies [PD18]. legitimacy [CH17, HMH^{+}15, SNT^{+}14, WWL^{+}10]. legitimization [Wha19]. Legitimizing [AMLL18]. Leibniz [DLSS19]. Leiden [PH18a, PH18b]. Lenfest [BSHG16]. lens [HBRUM13]. Leon [Vil11]. less [MMHV19]. Lessons [ABC^{+}12, ADLK12, GG14, Gie17, OR10, SFBB10]. Lester [Liv12]. Level [FM16, Coz12, FM15, HH13, JL19, KD10, LHGW14, Lor11, PCS10, RHJS15, SNL15, Vīt15]. Leveling [FM16]. Levidow [Cla13a]. Levratto [Val14a]. Lewis [FM16]. liability [HFWL19]. licensing [WLYCL15]. Lidskog [Cal12]. Life [EGI18, Son18, Gie17, Hro14]. light [SSK15]. Lighting [Ben13]. likely [KCS17]. Limitations [SSK15, Jul12]. Limits [Nie16a, BL10, Cro13]. Line [Sco14]. Linenstein [Bra13]. lines [BGY^{+}16]. link [MB18, Fit18a]. linkages [TDBS11, Zav18]. Linking [Gho18, LU17]. links [FdSdS^{+}10, Rad11, dC17]. Linnenluecke [Ken16]. liquid [TD12]. List [Ano10b]. Literacy [BW18]. literature [MFM18, VR18]. Live [Ogr17]. livestock [ADLK12]. living [BCB18]. Local [Aag15, HR19, Dra14, HM10, LT19, SMR16, Yan18]. locally [POD19, Gib13]. locally-situated [POD19]. Locating [MBD^{+}12, RHS12]. location [FSS18]. Logic [Gol17]. Logics [LC18, Åm13, DLSS19, SFZ16]. Logistics [Wix15]. Long [Doo18, Bon11, SvTvL15]. long-term [Bon11, SvTvL15]. look [AR11, Ant18, Lin14, Mat10, Mic11]. Looking [SSK15, AVL12, LVG11]. looks [Rob14]. loops [Uln15]. Love [Joh17]. Low [Liv12, Lov16, Mil15, ASSX16, HKS14, Hub14, Gib13]. Low-Carbon [Lov16, Liv12, ASSX16, Hub14]. Low-cost [Liv12]. Low-Tech [Mil15, HKS14]. Luca [Art12]. Lucia [Art12]. Luis [Gop14]. Lundvall [BFE11]. Lyme [dOB17]. Lynn [Hro14].

M [Das15, Lev15, Liv12, Tei12]. MA [Gut16, Sti17, Wil16]. macro [TD12]. macro-environment [TD12]. macroalgal [Bou19]. made [JVH14, Li11]. Madras [Kam11]. magician [Por16]. magicians [Wil10]. mainstream [SBD18, SSK15]. maintenance [SM10]. Make [FT17, Tei14, Åm13, KCS17]. makers [NdMSFC12, WH12a, WH12b]. Making [Art11, BE18, Lab14, Odu13, RH19, Web11, AGS19, DNA^{+}18, Esc14, FPVH18, HHJ10, KM13, KSC^{+}12, Kri15, Löv11, PE13, Por16, SPH10, ZP17]. Malaysian [Has11]. manage [Col10]. Management [AdCBBB15, KH12a, PD18, RHJS15, SvTvL15, Tas14, Van10, Haw12]. manager [WWL^{+}10]. managers [IH11]. Managing [HFS^{+}13, IH11, Mar13, Grø14, Sti17]. Manipulating [HFWL19]. manufacturing [Rad19, YHSH17]. many [JVH14]. Mapping [FTNdB14, MF17, MBHD12, RGvLR18]. marginal [SBD18]. Marianna [Haw15]. Market [vHDdJ17, ADK16, Art12, Gau12, MK13, VRL13, WLS12]. marketing [dMFPP14]. Markets [Alm16, CBB12, Hug13, Lor11, Cad14]. Mars [OPBCM14]. Martina [Ken16]. Martinez [Cat12]. Martinez-Fernandez [Cat12]. Mary [Bra13]. Massimo [Art12]. masts [MP12]. matching [ADS10]. material [FWS19, Wes12]. Mathias [Tei14]. matter [Bes15, CS15, HL11, Lor11]. Matters [Sha18]. Matthew [Del16, LBS^{+}15]. Maureen [Bey14, Grø14]. Max [Cla15, Har13]. Mazzucato [Haw15]. McKelvey [Bey14, Grø14]. meaningful [SSB19]. measured [BdV14]. Measurement [HL19, Haw12, Rob14, Rob14, Sri14]. measures [Rad19]. Measuring [ASvLW17, BHPW13, TK11]. mechanisms [LSBOPCM18, Yan18]. Media [Röd15, BRB18, Les13]. medical [AdCBBB15, KWJ17, LMDU16, MVFG15]. medicine [Bas11, MSZ18, SBD18, Wha19]. medium [BL16, Boo11, LSPC16, LKP16, Nik15]. medium- [Boo11]. medium-sized [BL16, LSPC16, LKP16, Nik15]. Meet [Alm16, Hug13, KCM^{+}12, WE18]. meets [LT19]. megascience [JH12]. Mehmet [Val14b]. Member [ZLRJ19, Tsi12]. men [FBB14]. meritocracy [Nie16a]. Mesman [Eva15]. meta [JRG19]. meta-synthesis [JRG19]. metaphor [Yaq12]. Metaphors [Wil10]. method [CC13]. Methodological [Egb14, SS17, YHR12]. methods [ABC^{+}12, WLS12]. Mets [Egb14]. Mexico [DDT10, dJMA12, Nef18, Nef19, Rod10, TDBS11, Vil11, WM18]. Michael [Gib13, Hol15]. micro [Boo11]. micro-sized [Boo11]. microbiology [BFZFMT18]. Microchips [Li11]. Microelectronics [Doo18]. middle [LSTC17]. migrant [SV15]. Miguel [RRMS12]. Mikkel [Mus12]. Miles [Cat12]. Milne [Su16]. Mind [WKGK11]. mines [Gro10]. miniature [dBdLHB14]. mirrors [Wil10]. misconduct [HFWL19, HBM19]. misleading [WM18]. Missing [Gut18]. Mission [Phi13, BUK16, JJ17, Kwi12, LD15, NZ16, TYM^{+}14]. misunderstand [Nef19]. Mitigation [Sha17]. Mix [RP17, Vīt15]. Mobile [BW18, MP12, WHLhC19]. Mobility [Hua17, BFG^{+}15, Bre12, CFO16, GvEV19, IG11, LT19, Son17, WHLhC19]. Mobilizing [Mat10]. Mode [Hlá17, KB11, AF14, CCF16, HFNW15, ZP17]. Mode-2 [KB11]. Model [LC18, All10, Bon14, DLSS19, FLCRP19, FL10, Lau11, MAS16, Vil11, WWL^{+}10]. Modeling [VCdS16]. modelling [AB19, UG19]. Models [Cla13c, Day19, VCLA15, BH11, Dog13, Grø11a, MFC12, OMW13, Wil16]. moderating [ZLGG19]. modern [Kin11]. modes [LGGK12, Wic13a]. modified [Cla13c, ML13, VMP10]. mole [Har14]. Money [Fit17]. monitoring [RdR17, SDP11]. monodisciplinarity [BdCB14]. Montserrat [DBO13]. MOOCs [Hab14, Wil16, Wil16]. Moon [Nel11]. Moore [Doo18]. Moral [Bor16, Dra14]. Morrone [Nef19]. Morten [Woo16]. Morton [Him16]. most [AC11]. motivation [JCK14, JC15]. motivational [vdBHC19]. Moving [ACBP16, BdCB14, ZP17]. Mozambique [Zav18]. Mubadala [MAS16]. much [Mus12]. much-needed [Mus12]. Muir [Fit13]. Multi [FM16, HH13, KD10, LHGW14, Lor11, Vīt15]. Multi-dimensional [HH13]. Multi-Level [FM16, HH13, KD10, LHGW14, Lor11, Vīt15]. Multidimensional [Sch15, CS14]. multilevel [KCW14, LSKvA17]. multinational [CPP^{+}13]. Multiplicity [SFZ16]. municipal [LU17]. museums [DLSS19]. mutineers [Wil10]. mutuality [JMSS18a, JMSS18b]. my [Nef19]. Myths [Rya10, Maz14, Haw15].

N [Fit18a]. Nadine [Val14a]. Nano [Ros10, KCS12, Boh14]. nano-scientists [KCS12]. nanofood [dBdLHB14]. nanomedicine [Jul12]. nanoscale [RRF^{+}19]. nanoscience [Wal15]. Nanoscientists [KCS17]. nanotechnologies [SCH^{+}13]. Nanotechnology [HMB14a, Åm13, BB13, FM15, FL13, Gro11b, HWC^{+}18a, HWC^{+}18b, Hei13, Mic11, SD15, VJ12, YSRF19, ZWL16]. Narayanamurti [Del16]. narrative [BJB11, LP16]. narrative-based [LP16]. National [cC12, Egb13, FS11, HBEvdV10, Lin19, Mat17, SH14b, SH14c, Aag15, AS16, ASvLW17, CPP^{+}13, CZ19, CP15, FS10, Gra11, HH12, LGGK12, Lv14, PCS10, RHJS15, SS17, Wes12, HLL14, Kam13, LS18]. nationalism [SS17]. nations [RMZP12, Val14b]. natural [KH12a, TK11]. nature [Gru19]. Neck [Das15]. need [AFW10, Has11, KVH13, Tas13]. needed [Mus12]. Neff [Esp13, WM18]. neglected [MVFG15]. Negotiated [Kri15]. negotiation [NK15]. Neoliberal [Tra18]. neonicotinoids [PM19]. Nepal [PRZ^{+}10]. Nephrops [SvTvL15]. Netherlands [FvDvR16, Lei11, PM18, SLvP10, vDTC19, vdWVvdB12]. Network [Ben13, DN18, Eng18, Lar12, YH14, FFF17, Isk17, KJ18, LDL^{+}10, LKR19, NZ16, SN13, WW17, YSRF19, Aar17]. Networked [Cap10, MMGM12, Álv13]. networking [Ciu10, PCS13, SZ11, SK11]. Networks [Gop14, Mus12, ADLK12, DSB^{+}13, FWS19, KH12a, Lan10a, PCS10, RGvLR18, SSTM12, SD15, Yaq12, ZLLU16, vHWW19]. neuroimaging [AdCBBB15]. neutral [Åm13]. Newcastle [GB15]. Next [KR18, EB18, HN15]. Next-Generation [KR18]. Nexus [Lem18, HFS^{+}13]. Ng [Kle12]. Ni [Shi14]. Niche [KHZ^{+}13]. Nigeria [OAE18]. Nil [Kor14]. Nine [BW18]. Niosi [Egb13]. No [CC13, Wic13b]. Non [Hua17, BJR17, KBA17, ML13, MVFG15, OPMGCM14, Van10, VRL13, Wag15, WO17]. non- [KBA17, Wag15]. non-academic [OPMGCM14]. Non-Competition [Hua17]. non-governmental [Van10]. non-instrumentality [BJR17]. non-market [VRL13]. non-profit [MVFG15]. non-safety [ML13]. non-state [WO17]. Noordwaard [SLvP10]. Nordic [BCSF17a, BCSF17b, TK17]. norms [SM10]. North [BKM10, SvTvL15]. Northampton [Sti17]. northeast [AG17b]. Norway [Bor16, CH17, KA17]. Norwegian [GK10, SK18]. note [Löv11]. novel [Hen19]. novelty [MMHV19]. Nuclear [Lov16, Hub14, PD18]. Nuevo [Vil11]. nutshell [Das15].

Objective [Jac12]. objectives [Tuu11]. Objects [BK18, WWL^{+}10]. obscuring [AMLL18]. Observatory [DBO13]. obtained [KDL19]. Ocean [BSHG16]. off [RdMdAFY16, vHDdJ17]. off-label [vHDdJ17]. offer [YHR12]. Office [Har14, Ben17, Bla14]. Officers [Lai17]. official [MBHD12]. officials [O’B13]. offs [RdMdAFY16, RGFLRP^{+}18, SR11, SNT^{+}14]. old [RdMdAFY16]. Oliver [Him16, Hoy13]. One [BPSB16, Son14b, Odu13]. online [Bes15]. Ontario [Tam17]. Ontological [dCM12]. open [BLKR17, BG16, HBEvdV10, Lau11, Pea16, YWH^{+}15]. opencast [Gro10]. Openness [CP15, Lv14, Por16]. Operating [Álv13]. operational [DRVWA18]. operationalizations [LDB^{+}12]. Ophuls [Wic13a]. Opinion [MBHD12]. Opportunities [Kam13, Klo13, HMH^{+}15, HN15, Mar11, McC13, RH19, RM14, TFB12]. opportunity [dMCL18, Val14a]. opposition [Dra14]. optimists [KCS12]. options [AdCBBB15]. Orchestrating [BCB18]. order [Han10, MSZ18]. ordinary [Haw17]. ordination [Kas10]. organisation [Dos12, Van10]. Organisational [HBM19, Gök10]. organisations [AV10, AC11, DA10, DDT10, Dut10, OR10, SSC17]. Organization [Ali15, Ken16, Bre12, SFBB10, SF19, HSM16]. Organizational [HFW11, Bor11, CS15, HH12, HH15, Han10, HJS17, KDL19, MJH^{+}16]. organizations [FFF17, Han10, JRG19, SFBB10, SC19a, SC19b]. organized [Sen17]. Organizing [Mot16, Mus12]. orientation [SBC16, ZLGG19, vdWVvdB12]. oriented [EB18, HHJ10, WHKB12]. Orienting [CAK12, KCM^{+}12]. origins [LLH13]. Orphan [BG16]. Oshri [Sti17]. other [GG14]. outcome [ZLGG19]. outcomes [LDL^{+}10, Lor11, NdMSFC12, WLS12]. Outdated [RNN18]. Output [KdR17, ASvLW17, MRR17]. outreach [Kas19]. overall [OAB10]. overcome [LSTC17]. Overcoming [RH15, RRH17].

Overview [AdC12, Bon14, Ste10]. own [AC11, Bla14]. owned [CD16]. Ownership [Hag18]. Oxford [Pag11].

P [Hro14]. Pacific [DSTS10]. Pager [Tra14]. pages [Alm16, Gut16, Lov16, Sti17, Wil16]. Pakistan [BKM10, BRA12]. Pandza [Sti17]. paperback [Wil16]. paradigm [BRA12, KPG10]. Paradox [RNN18, dJSvD16, RP17]. parks [SR11]. Parliamentary [PD18]. part [AF14]. partial [SM10]. participate [Bhu15, FvDvR16]. participation [BH11, Eng18, HA11, KDL19, Mer12, RG17, SCE19]. Participatory [DT11, KCH19, LVG11, RMZP12, Saa19, OMW13]. particle [HH13]. Particular [TRGA16]. Particularities [NJ18]. parties [VRL13]. partners [CDI16, OPMGCM14, SMR16]. partnership [FT12]. partnerships [MVFG15]. party [LKP16]. party-promoted [LKP16]. Patent [Hua17, ZLLU16, AR11, CD16, Dan17, HFW11, LKP16, LW19, SF19]. patent-based [Dan17]. patenting [AL11, HF17, LSL15, vDTC19]. patents [ADK16, AC11, CD16, Hem12, MB18, SF19]. paternalism [DT11]. Path [rL11, SL14]. Path-creating [rL11]. Paths [Tõn19, ASSX16, AdCBBB15, LSZ16a, MTZ17, SA16]. pathways [ASC16, BUK16, DSB^{+}13]. patient [PEV^{+}14, SSB19, dOB17]. Patients [Fit13, PEV^{+}14]. Patricia [Das15]. Patterns [CPP^{+}13, BL16, CFO16, HB16, Kwi18, TG11, WHLhC19]. Paul [Gop14]. Paula [Egb14, Hic12]. pays [Fra12]. Pearl [LKP16]. peculiar [LS12b]. peer [DP13, DS18a, Han10]. Penetrating [Tei14]. Pengfei [Shi14]. people [Li11, Zha19]. Perceived [LDB^{+}12, Waa17, BCSF17a, BCSF17b, RH19]. perception [BL16, KCW14]. Perceptions [VHAHGC12, AG17a, BFZFMT18, RB19, Saa19, SCH^{+}13, TKM18]. percolation [BFZFMT18]. Perfecting [POD19]. Performance [LS12a, ZLRJ19, AS16, ADM16, AB19, ADS10, BK19, BCSFC15, Bon11, CKS15, Kas19, KB17, LDS19, MPP^{+}17, MF16, MHB14, RMZP12, RCRR19, SS17, SK11, TKM18, Tsi12, YHSH17, vdBHC19]. Performance-based [ZLRJ19, BK19, KB17, RCRR19]. Performing [Kea13]. persistence [FG16, HH12, JH12, Wes12]. person [Soe10]. Personal [Hag18]. Perspective [Bhu15, BSHG16, Dut10, Elz12, FFF17, FSHM19, Gra11, HE11, HM10, Isk17, Kas13, LVG11, Phi17, VRL13, Zav18, Son14b, Son18]. Perspectives [HFS^{+}13, BCSF17a, BCSF17b, BPSB16, BCSFC15, EGI18, LSPC16, LH11, PEV^{+}14]. Peruvian [RB19]. pervasiveness [ÁJT18a, ÁJT18b]. pesticides [PdSA13]. Peters [Gib13]. Pharma [Joh17]. Pharmaceutical [Jul13, CR14, cC12]. PhD [FBS16, Waa17]. Phillips [Hro14]. phone [MP12]. Phones [BW18, MP12]. Photon [Wes12, HH13]. physics [HH13, Vic14]. Picard [Les13]. picture [KD10]. Pied [FCLG16]. Piper [FCLG16]. Piscitello [Art12]. Pitch [POD19]. Place [Mot12]. places [Hug13, Alm16]. plan [SvTvL15, Lar12]. Planet [Him16, Ogr17]. planning [Web11, Afz13]. Plans [SKY18a, SKY18b]. plant [SBD18]. Platforms [Har11, LSKvA17, Ric16]. Plato [Wic13a]. play [CAK12, MMHV19, TK11]. playing [FM16]. Plenary [SBKP17]. Plundered [Ogr17]. Poland [KB19]. poles [RCRR19]. Policies [Bra13, DN18, Fer16, IR17, Kas10, TKM18, TK17, TM12, vDTC19, ÁJT18a, ÁJT18b, AS10, Bor11, BL15, Cas12, Cho19, CHH18, FBB14, FS10, dAFUFd19, Gei15, HM10, Kwi12, LGGK12, Lar12, LH11, Lei11, LSL15, LLH13, LW19, dCM12, MTR13, SX19, SX20, TD12, Tup11, WW17, YWH^{+}15, ZI10, dJSvD16, Ors14]. Policy [Art12, Bor11, Cad14, Cal12, CPSL17, DSB^{+}13, Fra12, Gut16, HLL14, Hen19, Ko15, KR18, Lab14, Lem18, LP16, Lus16, Mes14, MTZ17, NJ18, RP17, SK18, Sha17, Son18, Tam17, AFW10, AG17a, AVL12, ABC^{+}12, AR11, Ant18, ABC^{+}18, ADS10, ART10, BG19, BSHG16, BJB11, Bhu15, BRA12, Blü18, BJS12, BE18, BFE11, BE15, Bru11b, CR14, Car18, CP16, Cla10, CH10, DCT18a, DCT18b, DMRJ13, DMMPC18, DNA^{+}18, EB18, EJB11, Ekl11, Elz12, Esc14, FWRF14, FBA19, FT12, Gal12, GvEV19, GBH11, HHJ10, Hal11, Hal15, HLI^{+}11, Has11, HMH^{+}15, HBEvdV10, HFS^{+}13, Isk17, JH12, JC18, JRG19, Jul13, KPG10, KM13, Kas13, Kin11, KKL17, Klo13, KA17, KB19, Kri15, KCH19, LH13, LDL^{+}10, LMMD14, Lev12, Lin14, LSTC17, Löv11]. policy [MAS13, MSZ18, Mer12, MK13, MF16, NdMSFC12, NK15, Nik15, NM15, O’B13, ORB19, Par10, PE13, Por16, PCS10, Rad11, RB19, Röd15, Rya15, SL11, Saa19, SNT^{+}14, SCH^{+}13, SV15, SvPLK14, Smi14, Son14b, SCB^{+}16, SD15, SK10, Tan17, Tas14, TDPV13, Tho16, TFB12, TFE11, Tuu11, UG19, VCLA15, VHAHGC12, VG14, Vīt15, WBB11, WH12a, WH12b, WJOS14, YHSH17, ZI10, ZP17, ZLL17, dC16, dJSvD16, dMCL18, dOB17, Har14, Afz13]. policy-based [Tan17]. Policy-driven [LP16, PCS10]. Policy-Making [Lab14, DNA^{+}18, Esc14, KM13, Kri15, PE13, ZP17]. policymaker [KB19]. policymaking [DS18b]. Polish [AWD16, Hei18, Kwi12]. Political [Boh14, CV12, KCS17, SCE19, VHAHGC12]. political-scientific [VHAHGC12]. Politicization [SV15, LS12b]. politicized [Gou11]. Politics [Cla13b, DMRJ13, KS17, KdR17, Rah15, Wic13a, AGS16, Gal14, dCM12, Mus12, Smi14, SPH10]. pollution [BSBD14]. poor [BGW12]. population [AL11]. Portfolio [HFW^{+}12, Hlá17, Bis19]. portrait [Art14]. Portugal [Lar12, SSC17]. Portuguese [SRC17]. Positioning [Tei14]. Post [Hei18, CETRW12]. post-Dickey-Wicker [CETRW12]. Post-Socialist [Hei18]. postdoctoral [GB19, WLS12]. potential [BSM14, RBdCBB10]. potentials [ADK16]. Power [Lov16, Mus12, ABC^{+}18, CGL13, Cho19, Hub14, LSZ16a]. PPI [dC16]. Practical [Bis19, Gut16]. Practice [Boe18, Con17, Das15, Bre12, FWRF14, Kee18, NGB14, Sch11, WBB11, Lus16]. Practice-based [Das15, NGB14]. Practices [Boe18, HGS11, LH11, TM12, Tup11, VG14]. Practise [Lab14]. practitioner [BSHG16]. precautionary [PM19]. Prediction [Hlá17]. Predictors [WCK^{+}14]. preferences [Lev12]. Press [Gut16, Wil16, ET19, Wag15, Pag11]. prestige [Kwi18]. price [Fra12]. principle [PM19]. principles [GST13]. priorities [HJS17, Him18, LP16]. prioritization [HJS17]. priority [BLS12, FM15]. privacy [FG13, Sev13, Sta13, SN13, WF13]. Private [Fra12, Gop14, Haw15, Arz10, FT12, GG14, LKP16, Maz14, SN13, VCdS16]. proactive [BLS16]. probable [Art11]. problems [RG17, SH14a, VHAHGC12, WH12a, WH12b]. process [BFZFMT18, HLI^{+}11, Kin11, Kro19a, LDS19, LVG11, SSB19]. process-based [Kro19a]. Processes [DN18, Das17b, DNA^{+}18, JRG19, JC15, MJH^{+}16, NZ16, Nie16a, RB19, SvPLK14, tKBKS18]. procurement [CM18, Cla15, LG16, Rad19, Rol13, TZI13, WE18]. produced [PD18]. producer [Joh11, Lau11]. producing [Aar17, LKR19, Löv11]. product [MVFG15]. production [BG19, BJR17, BCSF17a, BCSF17b, CVCBA18, Dan17, DBO13, HFNW15, HB16, HFW11, KdLABdP18, KB11, LDL^{+}10, LH11, NZ16, OMW13, PRZ^{+}10, Saa19, SvTvL15, Wha19]. productive [WS16]. productivity [AV18, ASvLW17, DBdPR18, FBB14, FCLG16, HL11, Kwi18, vHWW19]. products [MVFG15]. Professionalization [PH18a, PH18b]. professor [Vic14]. professors [ADM16]. profiles [TRGA16]. profit [MVFG15]. Program [SK18, Bre12, CP16, HFW^{+}12, BSHG16, LS12a, LS18, Sco14, Tas14, TWGJ11]. Programme [BBH18, KHZ^{+}13, BJR17, ACBP16, RRMS12]. Programmes [RP17, CP15, SBK^{+}19, SSC17]. programming [KH12b, PEV^{+}14]. programs [FvDvR16, OL13, RdR17]. Progress [Kle12, dAFUFd19]. progressive [Dos10]. Projecting [Aar17]. projects [Bul15, CBB12, FBS16, PCS13, VCdS16]. proliferation [Bou19]. Promise [Gol17, Car18]. Promises [Jul12]. promote [Han10]. promoted [LKP16]. promoters [Dun10]. Promoting [Tei14]. promotion [Agr12, LW19, Nie16a, SBY18]. properties [Ant18]. Property [Cad14, Hag18, AR11, BGW12, CRS17, CKS15, Hem12, Nik15, Tam13]. Proposal [CP16, Kro19a, SSB19]. proposals [Him18]. Proposition [Dra14]. Prospecting [AdCBBB15]. prospects [Agr12, Waa17]. protection [FG13]. PROTEE [VMP10]. providers [LSPC16]. Province [BKM10]. provision [SDP11]. proximity [HFNW15, RH15]. proxy [Kah18]. Public [Afz13, Cad14, CM18, Fra12, Gro11b, Haw15, Hro14, Lai17, LS12b, NSS19, Rol13, She15, WE18, AG17a, Agr12, ACBP16, Art14, Arz10, AV10, AC11, BN15, Bes15, BH11, CV12, CPP^{+}13, Cla15, Coo14, DCT18a, DCT18b, DP13, DA10, DDT10, Dut10, DS18b, Esc14, ET19, FT12, FFF17, dAFUFd19, GP13, HLI^{+}11, HFS^{+}13, Hol10, Jen14, JCC10, Jul13, Kas19, LH13, Mar11, dCM12, Maz14, MSZ18, Nik15, O’B13, OR10, PE13, PR10, Rad19, RMZP12, Röd15, SDP11, SFBB10, SN13, TD12, TZ18, TZI13, Tõn19, TM12, VHAHGC12, WBB11, dC16, Cla15, Gop14]. public-private [Arz10, FT12]. public-sector [Art14, GP13]. Publication [Nef18, ASvLW17, EG14]. Publications [Lun11, AL11]. publicly [Bul15]. publics [EJB11, HLH13, LKR19]. Pure [KS17, Saa19]. Pursuing [Kee18].

Qiongjie [Shi14]. qualitative [Phi13].

Quality

[Lab14, OR10, RdR17, LS12a, LW19, OAE18]. Quantifying [KdR17]. quantile [JL19]. Quantitative [CCR16]. Quebec [MBL16]. question [AGS16]. questions [Wil16, vR10].

R [Sch15, Sön14a]. R&D

[ASvLW17, BL16, BKKS15, BdV14, Bul15, Car18, CGL13, CP16, CPSL17, CR10, DSZ14, DLV10, FvDvR16, FLCRP19, FL10, GRLR17, HLL14, Hal18, KPG10, Kas13, Kim14, KG17a, LSPC16, LTU10, MPC10, NO16, OL13, ORB19, OAB10, RP17, SK18, SSC17, VJ12, VCdS16, Man15]. R&D-based [FvDvR16, CPSL17]. Rabe [Kee13]. Race [Fra12]. radical [Sen14]. Radosevic [Son14b]. Raggamby [Lab14]. Rainie [Álv13]. Raising [Wil16, BL10]. Rämö [Mot12]. rational [Tan17, Gut16]. rationale [TFE11]. rationales [CM18]. Rationality [Afz13]. RDI [HJS17]. reality [Boo11, Rad11, Tuu11, BL13]. reason [Gut16]. recipe [Odu13, Wic13b]. Recipes [BL13, CHH18]. Reconciling [Haw17]. recruitment [Nie16a]. Red [Jac12]. reducing [Mus16]. reduction [DO16]. referees [Ano10b]. reference [Coo14]. referral [WLS12]. Reflecting [BOP18]. reflection [SvPLK14]. Reflections [Boe18, NZ16]. reflexivity [Mar11]. Reforms [Su16, SBL15]. Reframing [SM10]. Regan [Alm16]. regarding [ABC^{+}18]. regenerative [MSZ18]. regime [Elz12, LDL^{+}10, RJ14]. Region [TK17, AG17b, Vīt15]. Regional [Egb13, FT12, MTR13, NM15, NJ18, SBC16, Sön14a, Swa17, CL15, FBA19, LSL15, MTZ17, NVAM14, RCRR19, TK11, UJ14]. regionalisation [KA17]. regionally [JNP16]. regionally-embedded [JNP16]. regions [MMHV19, NSS19, Tan10, TK11, Wol14, Kee16]. Regulating [CGS19]. Regulation [BL13, PdSA13, Val14b, Aho17, BFPCB10, EGI18, HMB14a, HM10, MP12, PM18]. regulator [All10]. regulators [KCS12]. Regulatory [SFCR18a, SFCR18b, FSK11, Kin11, ML13]. reimagine [FPVH18]. reinstitutionalization [Kwi12]. reintegration [LDS19]. related [AdC12, MAS13, OPBCM16]. relation [Art16]. relations [Joh11, KB19, Mar11, WS16]. Relationship [Hua17, KDL19, MRR17, SNL15]. relative [FG16]. relevance [BG19, HMH^{+}15, SNT^{+}14]. Remade [Him16]. Remarks [VCLA15]. Remote [Pri19]. renewable [Isk17, TBvB15]. Renewal [Hei18, HH13, Uln15]. rent [NO16]. repertoires [TSK^{+}13]. replace [TZ18]. representations [Boh14, LU17, WO17]. reproduction [SD15]. required [Wic13a]. Research [AS16, BHPW13, BL10, BLS12, BK18, BGY^{+}16, BBH18, CAAH18, Eva15, FPVH18, FM16, Fit17, HLL14, Hoy13, IR17, Kas10, KHZ^{+}13, Kro19b, LKP16, LK17, LS18, Luu14, PEV^{+}14, PCS13, SBK^{+}19, SK18, SLvP10, TKM18, TS15, TK17, Wha19, WSBTM15, YSRF19, Aar17, ADM16, All10, AdC12, AGS19, AV10, AC11, BK19, BLS16, BFPCB10, Bee13, BCSF17a, BCSF17b, BOP18, BCSFC15, BRB18, BS15, Blü18, Bor16, BAP19, CV12, CETRW12, CVCBA18, CS15, CLDS18, CPP^{+}13, CH10, CBB12, CP15, DLSS19, DP13, DMMPC18, DA10, DDT10, Dut10, EG14, EGI18, FvDvR16, FM15, FBA19, Flo14, FL13, FWS19, Gra11, Gro10, GK10, Han10, Hem12, HGS11, HDT17, Hig16, Him18, HMB14b, HL11, JH12, JCK14, JC15, JCC10, JS18, JL19, KdLABdP18]. research [Kas19, KB17, KKL17, KH12a, Ko15, KH12b, KM19, Kwi12, Kwo11, LSKvA17, LGGK12, LS15, LMMD14, LLH13, LHGW14, LS12b, MPP^{+}17, MFM18, MMGM12, MBM19, MJH^{+}16, MRR17, Nef18, Nef19, Nie16b, OPBCM14, OR10, OMS12, PJYL15, PH18a, PH18b, Phi13, PRZ^{+}10, PR10, Pre14, PCS10, RRF^{+}19, RVDWA18, RHS12, RH15, RRMS12, Rod10, RG17, RRH17, RdR17, Rya11, Rya15, SL11, SFBB10, SR11, SW19, Sch11, SKY18a, SKY18b, SC19a, SC19b, SV15, SvPLK14, Sco14, Sen14, SB11, SBKP17, SFCR18a, SFCR18b, Smi14, SS17, Sta13, STF17, SPH10, TWR11, TDPV13, Tho13, Tõn19, TWGJ11, TFE11, Uln15, VMP10, VRL13, VR18, Vic14, VG14, WBWW15, Wal15, WdJH14, WBB11, WKGK11, WJOS14, Wu13, YH14, YPH17, ZLRJ19, ZP17, dBdLHB14, dR18, dOB17]. research [vdWVvdB12, BG19, Kwi18, Egb14]. Research-Based [SK18]. Research-driven [LKP16]. research-policy [SV15]. Researcher [PRZ^{+}10, Kas19, Mar11, MBL16, WM18]. researchers [AV18, BFG^{+}15, CFO16, DRVWA18, Dut10, FdSdS^{+}10, GB19, Hem17, IG11, JVH14, MBL16, OPMGCM14, RRMS12, Saa19]. Researching [Men12, Wic13b]. Resilience [Ken16, Shi15]. Resilient [DO16, Ken16, Gho18]. resist [Zha19]. resistances [CV12]. resource [KH12a, TK11]. resource-based [TK11]. Resources [Kor14, JCK14]. responding [BJS12, vDFL14]. response [Cho19, dJSvD16]. responses [HBM19, SFZ16]. Responsible [Lem16, MBM19, OMS12, Sta13, dBdLHB14, AdCBBB15, FBA19, STF17]. responsiveness [FM15, Nie16b]. Restivo [Jac12]. restoring [Gro11b]. Restricted [SX19, SX20]. restricting [EG14]. Results [GAM^{+}19, KH12a, DSB^{+}13, YWH^{+}15]. retention [WLA19]. retrospective [Lin14]. Return [Pea16, Tan18]. returnee [LDS19]. returns [CGL13, HLL14]. Revenge [Wic13a]. Review [Afz13, Ali15, Alm16, Álv13, And15, Art12, Art14, Bey14, BK17, Bra13, Cad14, Cal12, Cat12, Cla13a, Cla13b, Cla15, Das15, Del16, Egb13, Egb14, Esp13, Eva15, Fer16, Fit13, Fit18a, Fra12, Gib13, Gop14, Grø14, Gut16, Hal17, Haw12, Haw15, Hic12, Him16, Hol15, Hoy13, Hro14, Jac12, Kee13, Kee16, Ken16, Ken17, Kle12, Kor14, Kro19b, Lab14, Lem16, Les13, Lev15, Lig14, Liv12, Lov16, Lus16, Man15, Mar13, Mes14, Mil13, Mil15, Mot12, Mot16, Mus12, Ors14, Pri19, Rah15, Ren13, Rob14, Sch15, SL14, Sea13, Sha17, She15, Shi14, Sön14a, Son18, Sri14, Sti17, Su14, Su16, Tan13, Tei12, Tra14, Val14a, Val14b, Wic13a, Woo16, CM18, CCR16, DP13, DS18a, FWRF14, Han10, MFM18, UG19, VR18, WE18]. review [WJOS14, Wil16, vDH14, KPG10]. Reviews [Tei14, BL13]. Revisited [KS17, Nel11]. Revisiting [DCT18a, DCT18b]. revival [SBD18]. revolution [Lan10b]. reward [KB17]. Rhetoric [HMB14a, KKL17]. Rica [OR10, VHAHGC12]. Rich [Fra12]. Richard [Liv12]. Richness [Egb14]. Rights [Cad14, CD16, CRS17, CKS15, Hem12]. RIS3 [FBA19]. Rise [Hal17, Swa17, LMMD14]. Risk [BL13, D’S13, Esp13, Tan17, Agr12, Cla13c, DO16, Hal15, HLH13, HMB14a, MP12, PM18, Web11, Wic13b]. Risks [Kor14, Mar13, Arz10, CGS19, GC13]. River [LKP16, SLvP10]. roadmapping [AVL12]. roadmaps [GP13]. roads [HA11]. Robert [Art14, Les13, Man15]. robustness [Gro11b, KH12a]. Roger [Sön14a]. Role [Cla15, NJ18, AGS16, BJS12, BGW12, Cas12, EB18, FWS19, HWC^{+}18a, HWC^{+}18b, KM19, LS14, LBMM13, MMHV19, MBHD12, PJYL15, RGFLRP^{+}18, Rol13, SL11, SZ11, SK11, Sta13, SD15, VHAHGC12, ZLGG19]. roles [CAK12, CL15, Coz12, FL13, HO19, LSPC16, PRZ^{+}10, TK11, TSK^{+}13]. Rolf [Cal12]. Rolfstam [Cla15]. Room [SLvP10]. Rossi [Art12, BK17]. Rossi-Lamastra [Art12]. Route [Hei18]. RRI [FBA19]. Rubalcaba [Gop14]. Rubik [Lab14]. rural [Cha16]. Russia [FT17, Klo13, TZ18].

s [SL17]. S&T [FS10, Mer12, SSC17]. safety [ML13]. Saharan [CDD^{+}19]. Sal [Jac12]. salary [Kwi18]. Sale [Tra18]. salience [WWL^{+}10]. same [CHH18]. satellite [WO17]. scale [Isk17, LF13]. Scaling [Ben13]. Scandinavia [Agr12]. Scandinavian [Ben17]. Scanning [KSC^{+}12, ABC^{+}12, vR10]. scenario [SG16]. Scherdin [Tei12]. scholar [LDS19, WH12a]. Scholarly [Sch15, CS14, SNL15]. scholars [WLS12, WH12b]. Schrage [Hol15]. Schumpeterian [SL14, Son14b, Kas13]. Schwartz [Tei14]. Schwinge [Mil15]. Science [Ano13, BSHG16, Bhu15, Cal12, CDD^{+}19, Cla13b, Doo18, FS10, Flo14, HLL14, Har14, Jac12, KS17, Kam13, Kle12, Lai17, Lem18, MB12, OL13, PH18a, PH18b, Rad11, RNN18, She15, Su14, Swa17, TFB12, WM18, ÁJT18a, ÁJT18b, AdC12, Art16, AMLL18, BRA12, BH11, Bis19, Blü18, Bro14, Bru11b, BAP19, Cap10, Coz12, Dag12, Dog13, DBO13, DO16, DBdPR18, Ekl11, Elz12, ET19, FFF17, FTNdB14, GB15, Gie17, Gou11, Hal11, HH13, HH15, HA11, HLH13, HSM16, HMH^{+}15, HGS11, HFS^{+}13, HL11, IH11, JH12, JVH14, JRG19, Jen14, JS18, Kah19, Kee18, KCM^{+}12, KCW14, KJ18, KA17, KB19, KB11, LH13, LDL^{+}10, LMMD14, Lin14, LS12b, LZ10, Löv11, Mar11, Mat10, dJMA12, MBD^{+}12, MK13, Nef18, OMS12, Par10, PE13, RRF^{+}19]. science [RM14, RG17, Saa19, SS10, SH14a, SR11, SMVdP14, SNT^{+}14, Sch11, SM13, SS17, SCB^{+}16, TD12, Tan17, Tho13, Tsi12, TSK^{+}13, UJ14, VCLA15, VHAHGC12, VG14, Yan18, YPH17, Hal15, Röd15, Flo14, Hic12, Art12]. science-and-technology [DBdPR18]. Science-industry [Rad11, Dog13]. Science-policy [BSHG16, Ekl11, KB19, VCLA15]. science-public [Mar11]. science/policy [LDL^{+}10]. Sciences [Gut16, HDT17, Kas10, KWJ17, OPMGCM14, WSBTM15, Hei18, LZ10]. Scientific [AL11, BFZFMT18, CCF16, KdLABdP18, AV18, AW10, Bon11, COC14, CHH18, Cro13, DRVWA18, FCLG16, HFW^{+}12, HBM19, HL11, IH11, KJ18, LDB^{+}12, MPP^{+}17, Mar15, Men12, MBL16, MJH^{+}16, OPBCM16, PJYL15, Pea16, RB19, RVDWA18, Rya15, Tei11, Tho16, Uln17, VHAHGC12, Wha19, vDFL14, HFW^{+}12]. Scientist [D’S13, LSBOPCM18, dJSvD16, BK19, Dos10, EGI18, TM12, WLS12]. scientists [BdCB14, Bes15, Bhu15, FBB14, HF17, Him18, HFW11, KCS12, KCS17, Lev12, O’B13, SH14b, SH14c, VG14, Wal15, vHWW19, BL13]. Scientization [KCH19]. Scotland [ART10, Gal12]. Scottish [ACBP16]. scrutinising [KA17]. scrutinizing [Xu15]. Sea [SvTvL15, Vīt15, Rei15]. Sean [Tra14]. search [MFC12, WLS12]. Searching [Bul15, KCW14]. section [EB18, FG13, HHJ10, Lin14, MK13, RVDWA18, RM14, Ves12]. Sector [Hal18, Haw15, Art14, Bee13, DBdPR18, dMFPP14, FERV11, FFF17, GP13, Gra11, Hem12, Hol10, JCC10, KKL17, LTU10, Maz14, Rya11, Sch11, TZ18, TWGJ11, TFE11, VCdS16, Waa17, vDTC19]. Sectoral [SBY18, cC12]. sectors [Cho19, Mur13, Rad19]. secularity [dJMA12]. Security [Mar13, HOSV13, WBWW15, WO17]. Seeds [Cla13b]. select [BRB18]. Selected [Ali15, Dos12]. selection [Bor16, CCF16, VCdS16]. Selee [Mes14]. self [Phi13]. self-descriptions [Phi13]. Semantic [KJ18]. senior [IH11]. seniority [ADM16]. sense [BE18, Bro14, KSC^{+}12, RH19]. sense-making [KSC^{+}12]. sensegiving [HBM19]. sensehiding [HBM19]. Sensemaking [HBM19]. Sensing [YH14]. September [Ano13]. serial [LS14]. series [Swa17]. servants [HFS^{+}13]. Servet [RRMS12]. Service [Cat12, LSPC16, Rad19]. Services [JC18, BdV14, BAP19, DLV10, Hen19, Hol10, KVH13, LSPC16, SDP11, Gop14]. set [BL13]. Setting [SC19a, SC19b, BLS12, FM15, dR18, dOB17]. settings [KG17b]. setup [CV12]. seven [HN15]. sexy [Åm13]. shadow [tKBKS18]. Shale [Wag15]. Shane [Gib13]. shape [KB19]. shaped [dAFUFd19]. Shapes [Hic12]. Sharing [Kor14, SB11]. Shawn [Alm16]. Shenzhen [LC18]. shift [KPG10]. Shifting [Boe18, RHJS15, SCH^{+}13, TS15]. shifts [KCH19]. shipping [BSBD14]. shops [SM13]. Should [Mes14, Bhu15]. side [Cho19, Mat10, Rad19]. sides [BN15]. Sifting [Bru11a]. Signaling [LG16]. silences [BB13]. simulating [YWH^{+}15]. simulation [WWL^{+}10]. Singapore [Aar17, Gie17]. sites [SN13]. siting [MP12]. situated [BLKR17, dCM12, POD19]. Situating [Kea13, SM13]. size [BPSB16, BS15, HL11, MJH^{+}16, OAB10, Son14b]. sized [BL16, Boo11, LSPC16, LKP16, Nik15]. sketchy [Art14]. Skills [Cat12, FG16, BE15]. SLAC [HH13]. Slavo [Son14b]. Slovakia [NSS19]. Small [BL10, LS18, Nik15, BL16, Boo11, LSPC16, LKP16, MPP^{+}17]. Smart [SL17, GC13, GvEV19, HB16, HO19, MTZ17]. SME [NZ16]. SMEs [CPSL17, FvDvR16, Has11, RP17, YZ11].

SNI [Nef19]. Social [Álv13, Cas12, Coz12, KA16, LK17, Mat17, Mer12, NJ18, SN13, BFZFMT18, BPSB16, Dag12, DBO13, Dos10, ET19, FG16, HLH13, HOSV13, HDT17, Kas10, KCW14, KWJ17, KH12a, KB11, LS12a, Mat10, OPMGCM14, OPBCM14, RRF^{+}19, RG17, TFB12, Ves12, WSBTM15, Yan18, YZ11, YPH17, Tei14]. Socialist [Hei18]. Societal [Con17, YSRF19, AdCBBB15, DRVWA18, DS18a, OPBCM16, RVDWA18, Ric16, RGvLR18, RG17, SBK^{+}19, WBWW15, WE18, dJSvD16, vdWVvdB12, Tei14]. societies [BCB18, HDT17]. Society [Hlá17, Lem18, AdC12, Kah19, MBD^{+}12, MB12, OMS12, Tsi12, TSK^{+}13, Wag15, YSRF19, SSB19]. Socio [LU17, Mus16, Sti17, Ano13, BE14, BCB18, SZ11, FM15]. Socio-Economic [Sti17, Ano13, SZ11]. Socio-Technical [Mus16, LU17, BE14, BCB18, FM15]. Sociology [Jac12, KB11, Tei11]. sociotechnical [CGS19]. solar [Cho19]. solid [LU17]. solution [EG14]. solutions [Hen19]. some [CRS17]. Soocheol [Su16]. Source [Hal15, Pea16, Wes12]. sourcing [HR12]. South [AAL19, AS10, Boo11, CRS17, HWC^{+}18a, HWC^{+}18b, Hem17, Kah19, Ko15, KG17b, Kwo11, RHS12, UJ14]. Southern [NM15]. Space [DCJ19, McC13, Mot12]. spacecraft [YHSH17]. Spaces [MMHV19, SN13]. Spain [ÁJT18a, ÁJT18b, CVCBA18, DBdPR18, MB18, SBC16]. Spallation [Hal15]. Spanish [FERV11, HR12]. spanning [JRG19]. Spatial [COC14, LSL15, Les13, Swa17]. speaking [CAAH18]. Special [LC18, BFE11, CDI16, Dut10, EB18, FG13, GBH11, HHJ10, LH11, Lin14, MK13, MPC10, RVDWA18, RM14, SZ11, TFE11, Ves12]. specialisation [HB16]. specialization [MTZ17]. specific [WCK^{+}14]. Specificity [ÁJT18a, ÁJT18b]. speech [Mar15]. spending [SMVdP14, SW19]. sphere [Jen14, YZ11, She15]. spillovers [NO16, RGFLRP^{+}18]. spin [RdMdAFY16, RGFLRP^{+}18, SR11]. spin-offs [RdMdAFY16, RGFLRP^{+}18, SR11]. Spirit [Art14]. sponsored [NO16]. spotlight [BSM14]. SPP [Ano10a]. squaring [Kas17]. Sri [AW10]. SSH [OPBCM14]. stabilizations [KB11]. staff [AC11]. stages [SBK^{+}19]. stake [Bhu15]. Stakeholder [RB19, SvTvL15, AG17a, OMW13]. Stakeholder-led [SvTvL15]. stakeholders [DS18b]. standardisation [HO19]. Standardising [BL13]. standardization [Tam13]. Standards [D’S13, Lin19, CGS19, Wic13b, BL13]. State [Haw15, Lev12, BSM14, GP13, Hig16, Maz14, WBWW15, WO17, Wu13]. state-of-the-art [GP13]. statements [JJ17, Phi13]. States [ZLRJ19, HSM16, LLH13, Tsi12, HFW11, Lei11, Rad11, Tõn19]. Statistical [Kam13]. steers [Nef19]. Stehrer [Man15]. Stem [Mar11, BFPCB10, Dra14, Flo14, Lev12, LLH13, dJMA12, SS10, SFCR18a, SFCR18b, Tho13, OPBCM14]. Stephan [Hoy13, Hic12]. Stephen [Lev15]. stereotypes [OPBCM14]. STI [BJS12, DN18, DNA^{+}18, GvEV19, dMCL18]. Stilgoe [Lem16]. Still [Sha18, GK10]. stimulate [TZI13, vDTC19]. stimulating [EBV17]. Stop [LVG11]. storage [Ric16, WCK^{+}14]. story [AS16, Ekl11, Wes12]. Stough [Sön14a]. Strategic [KKL17, Mes14, dMFPP14, WdJH14, vR10]. Strategies [BKKS15, AR11, BUK16, Ben17, Haw12, Kro19a, NVAM14, Par10, SFZ16]. Strategy [Hel18, Ors14, Jul13, KVH13, NdMSFC12, Tas13, YHSH17]. streams [BCSF17a, BCSF17b]. street [SSK15]. strengthened [WM18]. strengthening [Uln17]. Structural [FL10, HE11, AB19, KCW14, SC19a, SC19b]. structure [CR14, CH10, WS16]. structured [Kro19a]. structures [CD16, TG11]. Struggle [Cla13b]. STS [Mus12]. students [GRLR17]. Studies [Bey14, PH18a, PH18b, Sön14a, SS17, ADLK12, Hag10, Haw17, JRG19, KG17b, Mar16, RdMdAFY16]. study [AL11, BFZFMT18, Dut10, FT17, Hal11, HDT17, HR12, Kah19, Kam11, KBA17, KH12a, LBS^{+}15, LDL^{+}10, MBL16, NdMSFC12, RB19, RCRR19, SBC16, SKY18a, SKY18b, TG11, TK11, VCLA15, Vic14, WBB11, WE18, ZWL16, ZLL17]. Studying [Åm13]. Sub [CDD^{+}19]. Sub-Saharan [CDD^{+}19]. Subnetwork [CD16]. subsidiaries [HR19, SMR16]. subsidies [AV18, VCdS16]. subsidy [NSS19, SSC17]. substantial [KA17]. subsystem [Kin11]. Success [Ekl11, FBS16, AS16, CS10, IH11]. successes [LF13]. Sugimoto [Sch15]. Sundqvist [Cal12]. Super [Hal17]. Superconducting [Hal17]. supply [BN15, OAE18, Rad19]. supply-side [Rad19]. support [BN15, Cha16, HFW^{+}12, KCS17, KVH13, LSPC16, Nef19, NSS19, O’B13, SMVdP14, ZWL16]. supported [PJYL15]. supporting [TM12]. surprise [SG16]. surrogate [MAS16]. surveillance [GC13]. surveillance-enabled [GC13]. Survey [LK17, BFG^{+}15]. Susan [And15, Cla13a]. Susana [Mus16]. Sustainability [Shi15, Kas19, Kee18, KBA17, PRZ^{+}10, TWR11, TYM^{+}14]. Sustainable [BK18, Fer16, SL14, Val14a, Bre12, Gei15, HNRM11, KH12a, Lab14]. Sustaining [Esp13, Grø14, Ken17]. Sweden [Bor16, FT12, Hal15, Hal18, HLI^{+}11, NM15, VJ12, WKGK11]. Swedish [Boh14, Ekl11]. Swiss [ADS10]. Switzerland [PRZ^{+}10]. synthesis [JC18, JRG19, SL14]. synthesizing [KSC^{+}12]. Synthetic [MHB14, vDH14, Kea13]. System [Ben13, DSP17, Aag15, AB19, BCSF17a, BCSF17b, FERV11, FSHM19, Gra11, HO19, Klo13, OL13, UG19, WM18, Álv13, Liv12]. systematic [FWRF14, MFM18, UG19, VR18]. systematically [Nef19]. Systemic [WH12a, WH12b, AVL12]. Systems [Egb13, Mat17, Mus16, NJ18, ASvLW17, BFE11, BE14, CAK12, CL15, CZ19, cC12, CGS19, DS11, FS11, FSHM19, HFNW15, HNRM11, KB17, Lor11, TK11, TBvB15, WHKB12].

Tacit [RJ14]. Tackling [Gol17]. Taiwanese [cC12]. take [OPBCM16]. Takeover [Tra18]. tale [Ekl11, Kah18]. Tamara [Cat12]. tango [OPBCM16]. tank [Bla14]. Tanks [Mes14]. Tapping [HSM16]. targets [Car18]. Taxation [Su16]. taxonomical [GP13]. Teaching [NGB14, Das15, Das15]. team [JC15]. Tech [Mil15, HKS14, Tuu11]. Technical [Mus16, BSBD14, BE14, BCB18, FM15, Kam11, LU17]. techniques [CCR16]. techno [vDFL14]. techno-scientific [vDFL14]. Technological [BSBD14, Cad14, Eva15, Lig14, ASC16, Bru10, cC12, CBB12, dAFUFd19, FSHM19, HMB14b, LSW19, Mus16, SFZ16, TDPV13, TBvB15, Lar12]. Technologies [And15, Con17, SL17, STF17, Sti17, BB13, CT14, EJB11, FTNdB14, FG13, HWC^{+}18a, HWC^{+}18b, HOSV13, HNRM11, Lan10b, LMDU16, Pre14, RBdCBB10, SCH^{+}13, Zha19]. Technology [Ano13, Bla14, Boe18, Col10, Har14, HN15, HR12, Lem18, Lin19, Ors14, PH18a, PH18b, PR10, Rah15, Ric16, Tas14, TZ18, Wad17, WLYCL15, YHSH17, AG17a, ÁJT18a, ÁJT18b, AdCBBB15, ADS10, Ben17, BN15, Bhu15, BRA12, BH11, Bon11, BL15, CCF16, CCR16, Coz12, Dag12, DSTS10, DBD^{+}10, DT11, DBdPR18, FFF17, Gal14, GP13, Gru19, HHJ10, Har13, HSM16, HS16, KA16, Kas17, KCS12, KCW14, KVH13, LG16, Lin14, LVG11, MB18, PD18, Par10, PCS13, RM14, Rya10, SM10, SMVdP14, SSK15, Tas13, TFB12, VHAHGC12, VG14, WHKB12, WS16, FL13, Lin19, OL13]. technology-based [BN15, KVH13, Tas13]. technoscience [BJB11]. Teece [Ors14]. tempered [Dra14]. temporary [CFO16]. Ten [SBK^{+}19]. Tensions [CV12, SBL15, BOP18, CDD^{+}19, DS18b]. term [Bon11, SvTvL15]. terms [Coo14]. Terry [Alm16]. Testing [Cla13a, Kim14]. Texts [Fra12]. Thai [YZ11]. Thailand [NZ16]. Thakur [And15]. theater [WWL^{+}10]. their [AR11, AC11, Bes15, DDT10, KA16, Lei11, MMGM12, SCH^{+}13, Waa17]. Thematic [DN18, HJS17]. Theo [Rei15]. Theoretical [YHR12]. theories [Xu15]. Theory [Lab14, Su16, Val14b, BH11, Isk17, Wix11]. Therapeutic [She15, Jen14]. there [ORB19]. thermal [LU17]. Think [Lev15, Mes14, Ans14, Bes15, Bla14]. thinking [Car18]. Third [BUK16, Hal18, LD15, NZ16, TYM^{+}14, VRL13]. Thomas [Gut16, Kro19b]. Thompson [Flo14]. thought [Wic13a]. thousand [WLS12]. threats [RH19]. three [BL13, HGS11]. threshold [BL10]. Thriving [Art14]. Tim [Gib13]. Time [BJB11, Mot12, FPVH18, Odu13]. Timing [Mot12, BJB11]. tips [Bis19]. Tisdell [Ken17]. Tissue [Aar17]. Tõnis [Egb14]. Too [MJH^{+}16, vDH14]. tools [SB11]. top [Kwi18]. Totally [Sen14]. Tracing [KCH19]. track [LDS19]. Trade [Sco14, SNT^{+}14]. trade-offs [SNT^{+}14]. tradeoffs [HBEvdV10]. training [BE15, Lor11, TM12, WLA19]. trajectories [RHS12, ZLLU16]. Transdisciplinarity [BBH18, BdCB14]. Transfer [Dog13, Arz10, Ben17, CCF16, KA16, Kwo11, Lin19, LSBOPCM18, MB18, PR10, SM10, SSK15, Smi16, TZ18, VR18]. transferred [PR10]. Transformation [Hal18, Kee16, OL13, SBD18, SC19a, SC19b, Wes12, Wol14]. transformations [AFW10, SBK^{+}19]. transformative [Bru10, GvEV19, Kro19a, Sen14, Sen17]. Transforming [GvEV19, Del16]. Transgenic [Cla13b]. Transition [BBH18, Klo13, RdR17, SPH10]. transitioning [SKY18a, SKY18b]. translation [HBRUM13, PEV^{+}14]. Translational [Ano13, Blü18, KWJ17, RRMS12, RdR17, VG14]. transnational [KH12b, SKY18a, SKY18b, Tra14]. transparency [WO17]. trap [LSTC17]. treatment [LU17]. tree [VMP10, Pag11]. Trends [BS15, CDD^{+}19, SFBB10, BE18]. Trial [Cla13a, VMP10]. trickle [Aag15]. Triple [LC18, JMSS18a, JMSS18b, NZ16, RHJS15, RRH17, Bre12, FT12, Vil11, YHR12]. trust [ACBP16, BFPCB10, Gro11b, HLH13, Mar11, BFPCB10]. trustworthiness [ACBP16]. Tucker [Mar13]. Tunnel [Hal17]. turbine [ZLLU16]. Turn [KCH19]. Turning [Ciu10]. turns [Ano13]. Twenty [GG14, Mar16, WHMS13]. Twenty-first [WHMS13]. twists [Ano13]. Twitter [BFZFMT18]. Two [Dog13, CD16, Dun10, OPBCM16, WdJH14]. type [SF19]. Types [Yan18]. typology [FS10].

U.S [Del16]. Ueta [Su16]. Ugur [Val14b]. UK [Alm16, Lov16, Sti17, AR11, ART10, BJB11, EJB11, FM15, GG14, Gro11b, HDT17, HP12, LP16, MP12, PM18, PE13, Tan17, VR18, Web11, ZLL17]. unbundling [TZI13]. uncertainty [Cla13c]. unconventional [LKR19]. undergraduates [Ste10]. undermine [Nef18]. Understanding [BOP18, Cla15, Das15, Das17a, Das17b, Hen19, KHZ^{+}13, LK17, SCH^{+}13, Aho17, LDB^{+}12, Mis11, Par10]. Unequal [And15, CT14]. Unexpected [Alm16, Hug13]. Unfolding [WW17]. unified [HE11]. unintended [BK19, BG16].

Union [Dun10]. unit [HL11]. United [HFW11, Lei11]. Universitie [FSS18, JJ17, Kwo11]. Universities [BRB18, CDI16, Fit18b, KG17b, TRGA16, ZLL17, AWD16, AR11, CL15, CD16, CPP^{+}13, FLCRP19, Kwi12, Kwi18, LD15, MMHV19, NZ16, PR10, SZ11, TK11, Fit18a]. universities/public [CPP^{+}13]. University [BKKS15, BK17, FL13, GAM^{+}19, JMSS18a, JMSS18b, Pag11, SL11, Zav18, Ben17, BPSB16, BdV14, CD16, CCF16, Cha16, Cla10, DCJ19, FBS16, HF17, Hem17, JNP16, KKL17, KBA17, KM19, MB18, Phi13, RPK16, RdMdAFY16, RGFLRP^{+}18, SBC16, SF19, SBL15, TG11, TDBS11, TYM^{+}14, VRL13, VR18, Vil11, WS16, dC17, BLS16, CLDS18, DS11, KBA17, Lei11, MFM18, Smi16]. university-community [DCJ19]. University-industry [SL11, TG11, TDBS11, Vil11]. university-invented [CD16]. university-owned [CD16]. Unlocking [Liv12]. Unpreparedness [Hal15]. Unraveling [KVH13]. Uppsala [BLS16]. Upstream [Esc14]. uptake [OPBCM16]. Urban [Kee16, Shi15, Swa17, ET19, Wol14, Shi14]. Urgency [Gol17]. Uruguayan [ABC^{+}18]. Urve [Egb14]. US$109.95 [Lov16]. US$15.95 [Wil16]. US$150.00 [Sti17]. US$38.00 [Gut16]. US-born [vHWW19]. USA [CETRW12, DSB^{+}13, GB19, Gra11, HH12, SS10, Wu13]. usable [KB19]. Use [LG16, Mar13, Aag15, BSBD14, EG14, EGI18, Kri15, Ves12, WJOS14]. Useful [Kle12, GK10, JVH14, Löv11]. usefulnes [Aar17]. User [DBD^{+}10, Gau12, Lau11, TWR11, BOP18, FT17, Joh11]. User-driven [DBD^{+}10]. User-producer [Lau11, Joh11]. users [LSBOPCM18, RBdCBB10]. uses [CCR16]. Using [RGvLR18, CGL13, KB17, VMP10, WCK^{+}14]. utility [Nef18]. utilization [BLS16, BOP18, TDPV13]. Uzunidis [Val14a].

V [FM16]. valleys [dMCL18]. valorisation [vdBHC19]. valuation [Art16]. Value [Wix15, AFW10, ADLK12, Grø11a, HDT17, LS12a, OPBCM14, Tho16]. values [LD15]. Variation [SF19, HM10]. Varieties [WdJH14, Kim14]. variety [LSBOPCM18]. vehicle [SFZ16]. Venesaar [Egb14]. Venezuela [CV12]. Venezuelan [Mer12]. Venkatesh [Del16]. Venture [MF16, ART10, DCT18a, DCT18b, Esp13, LMDU16, Esp13]. Venturing [Mot16]. Venus [OPBCM14]. versus [LGGK12, Rad19, SKY18a, SKY18b, SG16, Zav18]. Victor [Rah15]. view [Him18, Kas10]. Viewed [Elz12]. Views [DSB^{+}13, EGI18, IH11, KJ18, WCK^{+}14]. Virtuous [Uln15]. Vision [Ogr17]. Visions [Hal17, VCLA15]. visits [LS15]. visual [WO17]. Volcano [DBO13]. Voting [Dra14]. vs [Haw15, KCW14, Maz14]. Vulnerability [HMB14b, Eva15].

W [Ren13]. Waiting [Gol17]. Walshok [Bra13]. warfare [Pri19]. warming [Ans14, Mis11, DSB^{+}13, Lev15]. waste [LU17, PD18, Tup11]. Water [BBH18, DB10, MHB14, WWL^{+}10]. Watson [FM16]. wave [HN15]. way [TZI13]. ways [JVH14]. Weak [AS10]. weaknesses [FSHM19]. weapons [RJ14, Mar13]. Weather [Ken16]. welfare [KB11]. Wellman [Álv13]. Wentworth [Van10]. were [SH14b, SH14c]. West [BKM10]. Westeren [Tan13]. Western [FT12, ORB19]. Weyman [Cat12]. wheat [GG14, ML13]. Which [KCS17, SH14c, LMDU16, SH14b]. White [Har14]. Who [Bru10, Fra12]. whole [AL11]. Wicker [CETRW12]. Wide [DSB^{+}13]. Wiebe [Eva15]. William [Gut16, Wic13a]. Williams [Nef19]. wind [ABC^{+}18, Cho19, Ric16, ZLLU16, LSZ16a].

Windows [dMCL18]. Windrum [Gop14]. wisdom [Bru11a, Rad11]. Wishful [Car18]. within [CV12, CGS19, FM15, Grø11a, JMSS18a, JMSS18b, KKL17, SSTM12, SD15]. without [Fer16, Gei15, Mus16]. Wolfe [Kee16]. Women [BW18, Joh17, LHGW14, FBB14, SW19, FM16]. Wonks [AG17a]. Work [Cat12, Esp13, FBB14, FT17, POD19, PM19, SM10, Tei11, Wha19]. workers [AMLL18, Mil14, Mil13]. workforce [WLA19]. workshop [SG16]. World [Ogr17, CT14, Gei15, Hub14, Odu13, WCK^{+}14, WHKB12, DSB^{+}13, And15, Fer16, Him16, Lov16]. worlds [Kea13]. Worth [Hol15, Sch14, WLS12]. Wrong [Woo16].

xenotransplantation [Coo14, EJB11, HLI^{+}11]. Xie [Su14].

years [AL11, GG14]. Yu [Su14].

Zahradnik [Man15]. Zander [Tei12]. Zealand [HL19, LMMD14, Men12]. Zheng [Shi14]. zombie [SS17]. Zone [LC18].

References

Aagaard:2015:HIT

[Aag15] Kaare Aagaard. How incentives trickle down: Local use of a national bibliometric indicator system. Science and Public Policy, 42 (5):725–737, October 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/42/5/725.

Amankwah-Amoah:2019:HEE

[AAL19] Joseph Amankwah-Amoah and Yingfa Lu. Historical evolution of entrepreneurial development in the global south: The case of Ghana, 1957– 2010. Science and Public Policy, 46(2):161–172, April 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/2/161/5055851.

Aarden:2017:PPU

[Aar17]  Erik Aarden.  Projecting

and producing ‘usefulness’ of biomedical research infrastructures; or why the Singapore Tissue Network closed. Science and Public Policy, 44(6):753–762, December 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

//academic.oup.com/spp/
article/44/6/753/3192392.

Agasisti:2019:ADH

[AB19] Tommaso Agasisti and Alice Bertoletti. Analysing the determinants of higher education systems’ performance — a structural equation modelling approach. Science and Public Policy, 46(6):834–852, December 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/6/834/5543869.

Amanatidou:2012:CMH

[ABC^{+}12] Effie Amanatidou, Maurits Butter, Vicente Carabias, Totti Könnölä, Miriam Leis, Ozcan Saritas, Petra Schaper-Rinkel, and Victor van Rij. On concepts and methods in horizon scanning: Lessons from initiating policy dialogues on emerging issues. Science and Public Policy, 39(2):208–221, March 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ardanche:2018:PWA

[ABC^{+}18] Melissa Ardanche, Mariela Bianco, Claudia Cohanoff, Soledad Contreras, María Goñi, Lucía Simón, and Judith Sutz. The power of wind: an analysis of a Uruguayan dialogue regarding an energy policy. Science and Public Policy, 45(3):351–360, June 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/3/351/4091601.

Azagra-Caro:2011:DPR

[AC11] Joaquín M. Azagra-Caro. Do public research organisations own most patents invented by their staff? Science and Public Policy, 38(3):237–250, April 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Alvarez:2015:HEF

[ÁC15] Roberto Álvarez and Gustavo A. Crespi. Heterogeneous effects of financial constraints on innovation: Evidence from Chile. Science and Public Policy, 42(5):711–724, October 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /spp.oxfordjournals.org/
                                        content/42/5/711.

Aitken:2016:MTT

[ACBP16] Mhairi Aitken, Sarah Cunningham-Burley, and Claudia Pagliari. Moving from trust to trustworthiness: Experiences of public engagement in the Scottish Health Informatics Programme. Science and Public Policy, 43(5):713–??, October 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/43/
                                        5/713/2525452/Moving-from-
                                        trust-to-trustworthiness-
                                        Experiences.

Anichini:2012:ORR

[AdC12] Giulia Anichini and Suzanne de Cheveigné. Overview of research related to science in society in Europe. Science and Public Policy, 39(6):701–709, December 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Arentshorst:2015:PRT

[AdCBBB15] Marlous E. Arentshorst, Tjard de Cock Buning, Wouter P. C. Boon, and Jacqueline E. W. Broerse. Prospecting responsible technology paths: Management options for an appropriate societal embedding of medical neuroimaging. Science and Public Policy, 42 (6):775–788, December 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Agiakloglou:2016:IIM

[ADK16] Christos Agiakloglou, Kyriakos Drivas, and Dimitris Karamanis. Individual inventors and market potentials: Evidence from US patents. Science and Public Policy, 43 (2):147–156, April 2016. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ayele:2012:EIL

[ADLK12] Seife Ayele, Alan Duncan, Asamoah Larbi, and Truong Tan Khanh. Enhancing innovation in livestock value chains through networks: Lessons from fodder innovation case studies in developing countries. Science and Public Policy, 39(3):333– 346, June 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Abramo:2016:CEA

[ADM16] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Gianluca Murgia. The combined effects of age and seniority on research performance of full professors. Science and Public Policy, 43(3):301–319, June 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Abramo:2015:DAC

[ADR15] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Francesco Rosati. The determinants of academic career advancement: Evidence from Italy. Science and Public Policy, 42(6):761– 774, December 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Arvanitis:2010:IST

[ADS10] Spyros Arvanitis, Laurent Donzé, and Nora Sydow. Impact of Swiss technology policy on firm innovation performance: an evaluation based on a matching approach. Science and Public Policy, 37(1): 63–78, February 2010. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Antonelli:2014:HKA

[AF14] Cristiano Antonelli and Claudio Fassio. The heterogeneity of knowledge and the academic mode of knowledge governance: Italian evidence in the first part of the 20th century. Science and Public Policy, 41(1):15–28, February 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Abadie:2010:AFC

[AFW10] Fabienne Abadie, Michael Friedewald, and K. Matthias Weber. Adaptive foresight in the creative content industries: anticipating value chain transformations and need for policy action. Science and Public Policy, 37(1):19–30, February 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Afzalan:2013:BRB

[Afz13] Nader Afzalan. Book review:

Planning with Complexity: an Introduction to Collaborative Rationality for Public Policy by Judith E. Innes and David E. Booher. Science and Public Policy, 40(6):821–822, December 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ablon:2017:WGE

[AG17a] Lillian Ablon and Andrea A.

Golay. Wonks and geeks: examining commercial technology stakeholders’ perceptions of and interactions with public policy. Science and Public Policy, 44(4):556– ??, August 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
4/556/2907803/Wonks-and-
geeks-examining-commercial-
technology.

Alon:2017:BID

[AG17b] Ido Alon and Manuel Mira Godinho. Business incubators in a developing economy: Evidence from Brazil’s northeast region. Science and Public Policy, 44(1):13– ??, February 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/44/
                                        1/13/2525490/Business-incubators-
                                        in-a-developing-economy.

Agrell:2012:FPR

[Agr12] Wilhelm Agrell. Framing prospects and risk in the public promotion of ESS Scandinavia. Science and Public Policy, 39(4):429–438, August 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Aerni:2016:RBC

[AGS16] Philipp Aerni, Florabelle Gagalac, and Joachim Scholderer. The role of biotechnology in combating climate change: a question of politics? Science and Public Policy, 43 (1):13–28, February 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /spp.oxfordjournals.org/
                                        content/43/1/13.

Arocena:2019:TMR

[AGS19] Rodrigo Arocena, Bo Göransson, and Judith Sutz. Towards making research evaluation more compatible with developmental goals. Science and Public Policy, 46(2):210–218, April 2019. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/2/210/5066365.

Aho:2017:DRU

[Aho17] Brett Aho. Disrupting regulation: understanding industry engagement on endocrinedisrupting chemicals. Science and Public Policy, 44 (5):698–706, October 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

//academic.oup.com/spp/
article/44/5/698/2996781.

Alvarez:2018:SPDa

[ÁJT18a] Isabel Álvarez, Myrna Juan, and Celia Torrecillas. Specificity and pervasiveness of dialogues in science, technology, and innovation policies in Spain. Science and Public Policy, 45(1):141, February 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/1/141/4106303.

Alvarez:2018:SPDb

[ÁJT18b] Isabel Álvarez, Myrna Juan, and Celia Torrecillas. Specificity and pervasiveness of dialogues in science, technology, and innovation policies in Spain. Science and Public Policy, 45(3):329–337, June 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/3/329/4091603.

Archambault:2011:SPP

[AL11] Éric Archambault and Vincent Larivière. Scientific publications and patenting by companies: a study of the whole population of Canadian firms over 25 years. Science and Public Policy, 38(4):269– 278, May 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Alic:2015:EOI

[Ali15] John A. Alic. Book review: Economic Organization, Industrial Dynamics and Development: Selected Essays by Giovanni Dosi. Science and Public Policy, 42(1):141–143, February 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Allyse:2010:EEE

[All10]  Megan Allyse.  Embryos,

ethics and expertise: the emerging model of the research ethics regulator. Science and Public Policy, 37(8): 597–609, October 2010. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Almeida:2016:EEM

[Alm16] Mariza Almeida. Book review: Environmental Entrepreneurship: Markets Meet the Environment in Unexpected Places, By Laura E. Huggins (with contributions from Shawn Regan and Terry Anderson). Edward Elgar, Cheltenham, UK, 2013, 215 pages, £73.00 (hardcover), ISBN 978-1-78195-396-9. Science and Public Policy, 43(1):144–145, February 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/43/1/144.

Alvarez:2013:BRB

[Álv13] J. Francisco Álvarez. Book review: Networked: The New Social Operating System by Lee Rainie and Barry Wellman. Science and Public Policy, 40(6):823–824, December 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Aam:2013:DMN

[Åm13] Heidrun Åm. ‘Don’t make nanotechnology sexy, ensure its benefits, and be neutral’: Studying the logics of new intermediary institutions in ambiguous governance contexts. Science and Public Policy, 40(4):466–478, August 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Axler:2018:IWB

[AMLL18] Renata E. Axler, Fiona A.

Miller, Pascale Lehoux, and Trudo Lemmens. The institutional workers of biomedical science: Legitimizing academic entrepreneurship and obscuring conflicts of interest. Science and Public Policy, 45(3):404–415, June 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/3/404/4596656.

Anderson:2015:BRB

[And15] Eric Anderson. Book review: Innovation and Inequality: Emerging Technologies in an Unequal World edited by Susan Cozzens and Dhanaraj Thakur. Science and Public Policy, 42(2):289–290, April 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /spp.oxfordjournals.org/
                                        content/42/2/289.

Anonymous:2010:IS

[Ano10a] Anonymous. Index, SPP 2010. Science and Public Policy, 37 (10):804–808, December 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:2010:LR

[Ano10b] Anonymous. List of referees 2009. Science and Public Policy, 37(1):83–84, February 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:2013:STI

[Ano13] Anonymous. 18th Science and Technology Indicators conference: “Translational twists and turns: Science as a socio-economic endeavor’, 4– 6th September, Berlin, Germany. Science and Public Policy, 40(1):142, February 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ansolabehere:2014:CCH

[Ans14] Stephen Ansolabehere. Cheap and clean: how Americans think about energy in the age of global warming. MIT Press, Cambridge, MA, USA, 2014. ISBN 0-262-02762-3 (hardcover). x + 261 pp. LCCN HD9502.5.C543 U624 2014.

Antonelli:2018:KPE

[Ant18] Cristiano Antonelli. Knowledge properties and economic policy: a new look. Science and Public Policy, 45 (2):151–158, April 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/2/151/4992964.

Andersen:2011:UUL

[AR11] Professor Birgitte Andersen and Federica Rossi. UK universities look beyond the patent policy discourse in their intellectual property strategies. Science and Public Policy, 38(4):254–268, May 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Avnimelech:2010:EIV

[ART10] Dr. Gil Avnimelech, Dr. Alessandro Rosiello, and Morris Teubal. Evolutionary interpretation of venture capital policy in Israel, Germany, UK and Scotland. Science and Public Policy, 37(2):101–112, March 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Artmann:2011:MIM

[Art11] Dr. Stefan Artmann. Making innovation more probable through innovative governance. Science and Public Policy, 38(10):818–820, December 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Artmann:2012:AGH

[Art12] Stefan Artmann. Book review: Advising government how to address market failure in innovation — and what about government failure? Science and Innovation Policy for the New Knowledge Economy edited by Massimo G. Colombo, Luca Grilli, Lucia Piscitello and Cristina Rossi-Lamastra. Science and Public Policy, 39(5): 690–691, October 2012. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Artmann:2014:BRS

[Art14]  Stefan Artmann.  Book

review: A sketchy portrait of the public-sector entrepreneur. Governpreneurship: Establishing a Thriving Entrepreneurial Spirit in Government by Robert D. Hisrich and Amr Al-Dabbagh. Science and Public Policy, 41(1):132– 133, February 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Artmann:2016:ABV

[Art16] Stefan Artmann. The assetbased valuation of everything and its relation to science. Science and Public Policy, 43(6):881–??, December 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/
6/881/2525543/The-asset-
based-valuation-of-everything-
and-its.

Arza:2010:CBR

[Arz10] Valeria Arza. Channels, benefits and risks of public-private interactions for knowledge transfer: conceptual framework inspired by Latin America. Science and Public Policy, 37(7):473–484, August 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Arocena:2010:WKD

[AS10] Rodrigo Arocena and Judith Sutz. Weak knowledge demand in the South: learning divides and innovation policies. Science and Public Policy, 37(8):571–582, October 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Aagaard:2016:RFN

[AS16] Kaare Aagaard and Jesper W. Schneider. Research funding and national academic performance: Examination of a Danish success story. Science and Public Policy, 43(4):518– 531, August 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Altenburg:2016:EEC

[ASC16] Tilman Altenburg, Eike W. Schamp, and Ankur Chaudhary. The emergence of electromobility: Comparing technological pathways in France, Germany, China and India. Science and Public Policy, 43(4):464–475, August 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Altenburg:2016:GEC

[ASSX16] Tilman Altenburg, Ambuj Sagar, Hubert Schmitz, and Lan Xue. Guest editorial: Comparing low-carbon innovation paths in Asia and Europe. Science and Public Policy, 43(4):451–453, August 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Aksnes:2017:MPN

[ASvLW17] Dag. W. Aksnes, Gunnar Sivertsen, Thed N. van Leeuwen, and Kaja K. Wendt. Measuring the productivity of national R&D systems: Challenges in crossnational comparisons of R&D input and publication output indicators. Science and Public Policy, 44(2):246–??, April 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
2/246/2525560/Measuring-
the-productivity-of-national-
R-amp-D.

Arza:2010:IBP

[AV10] Valeria Arza and Claudia Vazquez. Interactions between public research organisations and industry in Argentina. Science and Public Policy, 37(7):499–511, August 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Aboal:2018:ISR

[AV18] Diego Aboal and Maren Vairo. The impact of subsidies for researchers on the gender scientific productivity gap. Science and Public Policy, 45(4):515–532, August 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/4/515/4663003.

Ahlqvist:2012:IPR

[AVL12] Toni Ahlqvist, Ville Valovirta, and Torsti Loikkanen. Innovation policy roadmapping as a systemic instrument for forward-looking policy design. Science and Public Policy, 39 (2):178–190, March 2012. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anas:2010:BDS

[AW10] M. U. M. Anas and Seetha I. Wickremasinghe. Brain drain of the scientific community of developing countries: the case of Sri Lanka. Science and Public Policy, 37(5):381–388, June 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Agasisti:2016:EED

[AWD16] Tommaso Agasisti and Joanna Wolszczak-Derlacz. Exploring efficiency differentials between Italian and Polish universities, 2001–11. Science and Public Policy, 43(1):128–142, February 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /spp.oxfordjournals.org/
                                        content/43/1/128.

Brunet:2019:GRT

[BAP19] Lucas Brunet, Isabelle Arpin, and Taru Peltola. Governing research through affects: The case of ecosystem services science. Science and Public Policy, 46(6):866–875, December 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/6/866/5543534.

Bassett:2011:CIF

[Bas11] Deborah Bassett. Conflicts of interest and the future of medicine. Science and Public Policy, 38(9):739–740, November 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Beumer:2013:ETI

[BB13] Koen Beumer and Sujit Bhattacharya. Emerging technologies in India: Developments, debates and silences about nanotechnology. Science and Public Policy, 40(5):628–643, October 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brouwer:2018:TTW

[BBH18] Stijn Brouwer, Chris Büscher, and Laurens K Hessels. Towards transdisciplinarity: a water research programme in transition. Science and Public Policy, 45(2):211–220, April 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/2/211/4540815.

Bugge:2018:GST

[BCB18] Markus M. Bugge, Lars Coenen, and Are Branstad. Governing socio-technical change:

Orchestrating demand for assisted living in ageing societies. Science and Public Policy, 45(4):468–479, August 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/4/468/4915391.

Begin-Caouette:2017:PIFa

[BCSF17a] Olivier Bégin-Caouette, Evanthia Kalpazidou Schmidt, and Cynthia C. Field. The perceived impact of four funding streams on academic research production in Nordic countries: the perspectives of system actors. Science and Public Policy, 44(4):598– ??, August 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
4/598/3861499/The-perceived-
impact-of-four-funding-
streams-on.

Begin-Caouette:2017:PIFb

[BCSF17b] Olivier Bégin-Caouette, Evanthia Kalpazidou Schmidt, and Cynthia C Field. The perceived impact of four funding streams on academic research production in Nordic countries: the perspectives of system actors. Science and Public Policy, 44 (6):789–801, December 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

//academic.oup.com/spp/
article/44/6/789/3603522.

Bin:2015:ERP

[BCSFC15] Adriana Bin, Ana M. Carneiro, Sergio Salles-Filho, and Fernando A. B. Colugnati. Employment, research performance and decentralization: The experience and perspectives of doctorate holders in Brazil. Science and Public Policy, 42(5):646–660, October 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/42/5/646.

Benard:2014:MMT

[BdCB14] Marianne Benard and Tjard de Cock-Buning. Moving from monodisciplinarity towards transdisciplinarity: Insights into the barriers and facilitators that scientists faced. Science and Public Policy, 41 (6):720–733, December 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bohrer:2014:CRI

[BdV14] Cariza Teixeira Bohrer and Eduardo Raupp de Vargas. Can R&D be identified and measured in services? Empirical evidence from university hospitals. Science and Public Policy, 41(6):793–808, December 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Borras:2014:GST

[BE14] Susana Borrás and Jakob Edler, editors. The governance of socio-technical systems: explaining change, EU-SPRI Forum on Science, technology and innovation policy series. Edward Elgar, Cheltenham, UK, 2014. ISBN 1-78471-018-0 (hardcover). LCCN HC79.T4 G669 2014.

Borras:2015:ETS

[BE15] Susana Borrás and Charles Edquist. Education, training and skills in innovation policy. Science and Public Policy, 42(2):215–227, April 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /spp.oxfordjournals.org/
                                        content/42/2/215.

Boon:2018:DCI

[BE18] Wouter Boon and Jakob Edler. Demand, challenges, and innovation. making sense of new trends in innovation policy. Science and Public Policy, 45(4):435–447, August 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/4/435/4915393.

Beerkens:2013:CCA

[Bee13] Maarja Beerkens. Competition and concentration in the academic research industry: an empirical analysis of the sector dynamics in Australia 1990–2008. Science and Public Policy, 40(2):157–170, April 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bennertz:2013:SGL

[Ben13] Rafael Bennertz. Scaling up gas lighting: From the laboratories to the large integrated network system. Science and Public Policy, 40(1):140–141, February 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bengtsson:2017:CUT

[Ben17] Lars Bengtsson. A comparison of university technology transfer offices’ commercialization strategies in the Scandinavian countries. Science and Public Policy, 44(4):565– ??, August 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
4/565/2888428/A-comparison-
of-university-technology-
transfer.

Besley:2015:WDS

[Bes15] John C. Besley. What do scientists think about the public and does it matter to their online engagement? Science and Public Policy, 42(2):201–214, April 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/42/2/201.

Beyerlein:2014:HED

[Bey14]  Michael Beyerlein.  Book

review: How Entrepreneurs Do What They Do: Case Studies in Knowledge Intensive Entrepreneurship edited by Maureen McKelvey and Astrid Heidemann Lassen. Science and Public Policy, 41(2):263–265, April 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /spp.oxfordjournals.org/
                                        content/41/2/263.full.pdf+
                                        html.

Borras:2011:ISI

[BFE11] Susana Borrás, Jan Fagerberg, and Charles Edquist. Introduction to special issue on learning, innovation systems and policy in honour of Bengt-Åke Lundvall. Science and Public Policy, 38(9):666–668, November 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Boring:2015:IMF

[BFG^{+}15] Pål Børing, Kieron Flanagan, Dimitri Gagliardi, Aris Kaloudis, and Aikaterini Karakasidou. International mobility: Findings from a survey of researchers in the EU. Science and Public Policy, 42 (6):811–826, December 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bates:2010:HDW

[BFPCB10] Stephen R. Bates, Wendy Faulkner, Sarah Parry, and Sarah Cunningham-Burley. ‘How do we know it’s not been done yet?!’ Trust, trust building and regulation in stem cell research. Science and Public Policy, 37(9):703–718, November 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Barros:2018:SKP

[BFZFMT18] Beatriz Barros, Ana Fernández-Zubieta, Raul Fidalgo-Merino, and Francisco Triguero. Scientific knowledge percolation process and social impact: a case study on the biotechnology and microbiology perceptions on Twitter. Science and Public Policy, 45(6):804–814, December 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/6/804/4925527.

Bigliardi:2016:UEO

[BG16] Barbara Bigliardi and Francesco Galati. The unintended effect of the Orphan Drug Act on the adoption of open innovation. Science and Public Policy, 43(6):840–??, December 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/
6/840/2525464/The-unintended-

                                        effect-of-the-Orphan-Drug-
                                        Act-on.

Bandola-Gill:2019:BRE

[BG19] Justyna Bandola-Gill. Between relevance and excellence? Research impact agenda and the production of policy knowledge. Science and Public Policy, 46(6):895–905, December 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/46/6/895/5542640.

Booker:2012:EGP

[BGW12] Kayje M. Booker, Ashok J. Gadgil, and David E. Winickoff. Engineering for the global poor: The role of intellectual property. Science and Public Policy, 39(6):775–786, December 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bozeman:2016:RCE

[BGY^{+}16] Barry Bozeman, Monica Gaughan, Jan Youtie, Catherine P. Slade, and Heather Rimes. Research collaboration experiences, good and bad: Dispatches from the front lines. Science and Public Policy, 43(2):226–244, April 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Biegelbauer:2011:DTC

[BH11] Peter Biegelbauer and Janus Hansen. Democratic theory and citizen participation: democracy models in the evaluation of public participation in science and technology. Science and Public Policy, 38(8): 589–597, October 2011. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Barre:2013:MIC

[BHPW13] Remi Barré, Luisa Henriques, Dimitrios Pontikakis, and K. Matthias Weber. Measuring the integration and coordination dynamics of the European Research Area. Science and Public Policy, 40(2):187– 205, April 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bhushan:2015:PST

[Bhu15] Bharat Bhushan. Perspective:

Science and technology policy – what is at stake and why should scientists participate? Science and Public Policy, 42 (6):887–900, December 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bisbal:2019:PTE

[Bis19] Gustavo A. Bisbal. Practical tips to establish an actionable science portfolio for climate adaptation. Science and Public Policy, 46(1):148–153, February 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/1/148/5224983.

Beynon-Jones:2011:TTN

[BJB11] Siân M. Beynon-Jones and Nik Brown. Time, timing and narrative at the interface between UK technoscience and policy. Science and Public Policy, 38(8):639–648, October 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Basu:2017:BDI

[BJR17] Soutrik Basu, Joost Jongerden, and Guido Ruivenkamp. Beyond the dichotomy of instrumentality and noninstrumentality of knowledge production: The case of generation challenge programme. Science and Public Policy, 44(4):583–??, August 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/44/
                                        4/583/3609987/Beyond-the-
                                        dichotomy-of-instrumentality-
                                        and-non.

Boden:2012:RFR

[BJS12] Mark Boden, Ron Johnston, and Fabiana Scapolo. The role of FTA in responding to grand challenges: a new approach for STI policy? Science and Public Policy, 39(2):135–139, March 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Blume-Kohout:2017:BRU

[BK17] Margaret E. Blume-Kohout.

Book review: The University and the Economy by Aldo Geuna and Federica Rossi. Science and Public Policy, 44(1):147–??, February 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
1/147/2525544/The-University-
and-the-Economyby-Aldo-
Geuna-and.

Berker:2018:BOF

[BK18] Thomas Berker and Ann Kristin Kvellheim. Boundary objects as facilitators in sustainable building research. Science and Public Policy, 45(2):202–210, April 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/2/202/4759457.

Bak:2019:UCP

[BK19] Hee-Je Bak and Do Han Kim. The unintended consequences of performancebased incentives on inequality in scientists’ research performance. Science and Public Policy, 46(2):219–231, April 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/2/219/5070425.

Blume-Kohout:2015:URF

[BKKS15] Margaret E. Blume-Kohout, Krishna B. Kumar, and Neeraj Sood. University R&D funding strategies in a changing Federal funding environment. Science and Public Policy, 42(3):355–368, June 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /spp.oxfordjournals.org/
                                        content/42/3/355.

Bashir:2010:ILP

[BKM10] Dr. Tariq Bashir, Khalid Khan, and Dr. Khaleel Malik. The innovation landscape of Pakistan’s North West Frontier Province. Science and Public Policy, 37(3):181–191, April 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bearse:2010:EIR

[BL10] Peter M. Bearse and Albert N. Link. Economic implications of raising the threshold funding limits on US small business innovation research awards. Science and Public Policy, 37(10):731–735, December 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brierley:2013:IST

[BL13] Meaghan Brierley and Cooper H. Langford. [introduction to set of three reviews on Standards: Recipes for Reality (Busch, 2011), and Scientists and the Regulation of Risk: Standardising Control (Demortain, 2011)]. Science and Public Policy, 40(5):698, October 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bozeman:2015:TAI

[BL15] Barry Bozeman and Albert N.

Link. Toward an assessment of impacts from US technology and innovation policies. Science and Public Policy, 42(3):369–376, June 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/42/3/369.

Belitz:2016:FPR

[BL16] Heike Belitz and Anna Lejpras.

Financing patterns of R&D in small and medium-sized enterprises and the perception of innovation barriers in Germany. Science and Public Policy, 43(2):245–261, April 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Blair:2014:COT

[Bla14] Peter D. Blair. Congress’s own think tank: Learning from the legacy of the Office of Technology Assessment (1972–95). Science and Public Policy, 41 (4):449–457, July 2014. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bezuidenhout:2017:BDD

[BLKR17] Louise M. Bezuidenhout, Sabina Leonelli, Ann H. Kelly, and Brian Rappert. Beyond the digital divide: Towards a situated approach to open data. Science and Public Policy, 44(4):464–??, August 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
4/464/3959654/Beyond-the-
digital-divide-Towards-
a-situated.

Benner:2012:HCF

[BLS12] Mats Benner, Li Liu, and Sylvia Schwaag Serger. Head in the clouds and feet on the ground: Research priority setting in China. Science and Public Policy, 39(2):258–270, March 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Baraldi:2016:PAU

[BLS16] Enrico Baraldi, Marcus Lindahl, and Kristofer Severinsson. A proactive approach to the utilization of academic research: The case of Uppsala University’s AIMday. Science and Public Policy, 43(5):613– ??, October 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/
5/613/2525533/A-proactive-
approach-to-the-utilization-
of.

Blumel:2018:TRS

[Blü18] Clemens Blümel. Translational research in the science policy debate: a comparative analysis of documents. Science and Public Policy, 45(1):24–35, February 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/1/24/4093159.

Bergek:2015:ISD

[BN15] Anna Bergek and Charlotte Norrman. Integrating the supply and demand sides of public support to new technologybased firms. Science and Public Policy, 42(4):514–529, August 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Boeva:2018:SPR

[Boe18] Yana Boeva. Shifting practices: Reflections on technology, practice, and innovation. Science and Public Policy, 45(2):283–284, April 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/2/283/4094885.

Boholm:2014:PRN

[Boh14] Max Boholm. Political representations of nano in Swedish government documents. Science and Public Policy, 41(5): 575–596, October 2014. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bonaccorsi:2011:ECI

[Bon11] Andrea Bonaccorsi. European competitiveness in information technology and long-term scientific performance. Science and Public Policy, 38(7): 521–540, August 2011. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bonvillian:2014:NMI

[Bon14] William B. Bonvillian. The new model innovation agencies: an overview. Science and Public Policy, 41(4):425– 437, July 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Booyens:2011:SMM

[Boo11] Irma Booyens. Are small, medium- and micro-sized enterprises engines of innovation? The reality in South Africa. Science and Public Policy, 38(1):67–78, February 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Benneworth:2018:RTR

[BOP18] Paul Benneworth and Julia Olmos-Peñuela. Reflecting on the tensions of research utilization: Understanding the coupling of academic and user knowledge. Science and Public Policy, 45(6):764–774, December 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/6/764/4925384.

Borras:2011:PLO

[Bor11] Susana Borrás. Policy learning and organizational capacities in innovation policies. Science and Public Policy, 38 (9):725–734, November 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Borlaug:2016:MHA

[Bor16] Siri Brorstad Borlaug. Moral hazard and adverse selection in research funding: Centres of excellence in Norway and Sweden. Science and Public Policy, 43(3):352–362, June 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bourblanc:2019:EAF

[Bou19] Magalie Bourblanc. Expert assessment as a framing exercise: The controversy over green macroalgal blooms’ proliferation in France. Science and Public Policy, 46(2):264– 274, April 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/2/264/5123559.

Benneworth:2016:OSD

[BPSB16] Paul Benneworth, Rómulo Pinheiro, and Mabel Sánchez-Barrioluengo. One size does not fit all! New perspectives on the university in the social knowledge economy. Science and Public Policy, 43(6):731– ??, December 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/
6/731/2525519/One-size-
does-not-fit-all-New-perspectives-
on-the.

Bhutto:2012:IST

[BRA12] Arabella Bhutto, Pir Irfanullah Rashdi, and Qazi Moinuddin Abro. Indicators for science and technology policy in Pakistan: Entering the science, technology and innovation paradigm. Science and Public Policy, 39(1):1–12, February 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Branscomb:2013:BRB

[Bra13] Lewis M. Branscomb. Book review: Creating Competitiveness: Entrepreneurship and Innovation Policies for Growth edited by David B. Audretch and Mary Linenstein Walshok. Science and Public Policy, 40(6):828–829, December 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Blasi:2018:UCH

[BRB18] Brigida Blasi, Sandra Romagnosi, and Andrea Bonaccorsi. Universities as celebrities? how the media select information from a large research assessment exercise. Science and Public Policy, 45(4):503–514, August 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/4/503/4638488.

Bressers:2012:THO

[Bre12] Nanny Bressers. The Triple Helix organization in practice: Assessment of the Triple Helix in a Dutch sustainable mobility program. Science and Public Policy, 39(5):669–679, October 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brown:2010:CCB

[Bro10] Mark B. Brown. Charting a course for bioethics. Science and Public Policy, 37(9):724– 725, November 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brostrom:2014:ISW

[Bro14] Anders Broström. Interaction with science: In what sense a case of learning by doing? Science and Public Policy, 41(2):141–150, April 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/41/2/141.full.pdf+
html.

Bruggeman:2010:WGT

[Bru10]  David Bruggeman.  Who

is governing transformative technological innovation? Science and Public Policy, 37 (10):797–798, December 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bruggeman:2011:SWG

[Bru11a] David Bruggeman. Sifting the wisdom of greybeards. Science and Public Policy, 38(5):416– 417, June 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bruggeman:2011:SSP

[Bru11b] Dr. David Bruggeman. The science of science policy. Science and Public Policy, 38(8): 662–664, October 2011. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bloch:2015:SRF

[BS15] Carter Bloch and Mads P. Sørensen. The size of research funding: Trends and implications. Science and Public Policy, 42(1):30–43, February 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bloor:2014:IUT

[BSBD14] Michael Bloor, Helen Sampson, Susan Baker, and Katrin Dahlgren. The instrumental use of technical doubts: Technological controversies, investment decisions and air pollution controls in the global shipping industry. Science and Public Policy, 41(2):234–244, April 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/41/2/234.full.pdf+
html.

Bednarek:2016:SPI

[BSHG16] Angela T. Bednarek, Ben Shouse, Charlotte G. Hudson, and Rebecca Goldburg. Science-policy intermediaries from a practitioner’s perspective: The Lenfest Ocean Program experience. Science and Public Policy, 43(2):291–300, April 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Boucher:2014:BUS

[BSM14] Philip Boucher, Robert Smith, and Kate Millar. Biofuels under the spotlight: The state of assessment and potential for integration. Science and Public Policy, 41(3):283–293, June 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Barrioluengo:2016:TMI

[BUK16] Mabel Sánchez Barrioluengo, Elvira Uyarra, and Fumi Kitagawa. Third mission as institutional strategies: Between isomorphic forces and heterogeneous pathways. Science and Public Policy, 43(6):736– ??, December 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/

spp/article-abstract/43/
6/736/2525512/Third-mission-
as-institutional-strategies-
Between.

Bulathsinhala:2015:EAE

[Bul15] Nadika A. Bulathsinhala. Exante evaluation of publicly funded R&D projects: Searching for exploration. Science and Public Policy, 42(2): 162–175, April 2015. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http://

spp.oxfordjournals.org/
content/42/2/162.

Brierley:2018:MPL

[BW18] Meaghan Brierley and Melanie Walker. Mobile phones & literacy: Empowerment in women’s hands: a crosscase analysis of nine experiences. Science and Public Policy, 45(1):134–135, February 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/1/134/4049582.

Currie-Alder:2018:RAS

[CAAH18] Bruce Currie-Alder, Rigas Arvanitis, and Sari Hanafi. Research in Arabic-speaking countries: Funding competitions, international collaboration, and career incentives. Science and Public Policy, 45(1):74–82, February 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/1/74/4107903.

Cadwalader:2014:BRB

[Cad14] Erin L. Cadwalader. Book review: Intellectual Property Rights and the Financing of Technological Innovation: Public Policy and the Efficiency of the Markets by Carl Benedikt Frey. Science and Public Policy, 41(2):258–260, April 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/41/2/258.full.pdf+
html.

Cagnin:2012:OEI

[CAK12] Cristiano Cagnin, Effie Amanatidou, and Michael Keenan. Orienting European innovation systems towards grand challenges and the roles that FTA can play. Science and Public Policy, 39(2):140–152, March 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Callison:2012:BRB

[Cal12] Candis Callison. Book review: Governing the Air: The Dynamics of Science, Policy, and Citizen Interaction by Rolf Lidskog and Göran Sundqvist. Science and Public Policy, 39(4):542–543, August 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Caplan:2010:NSN

[Cap10] Avi Caplan. Networked science for a networked century. Science and Public Policy, 37 (2):162–163, March 2010. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Carvalho:2018:WTA

[Car18]  Adão Carvalho.  Wishful

thinking about R&D policy targets: what governments promise and what they actually deliver. Science and Public Policy, 45(3):373–391, June 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/3/373/4600797.

Casas:2012:SCD

[Cas12] Rosalba Casas. Social cohesion in distributive policies and the role of knowledge. Science and Public Policy, 39(5): 562–569, October 2012. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Catey:2012:BRB

[Cat12] Andrew Scott Catey. Book review: The Knowledge Economy at Work: Skills and Innovation in Knowledge Intensive Service Activities edited by Cristina Martinez-Fernandez, Ian Miles and Tamara Weyman. Science and Public Policy, 39(6):837–838, December 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cuhls:2012:FFM

[CBB12] Kerstin Cuhls, Alexander Bunkowski, and Lothar Behlau. Fraunhofer future markets: From global challenges to dedicated, technological, collaborative research projects. Science and Public Policy, 39(2): 232–244, March 2012. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Chung:2012:NST

[cC12]  Chao chen Chung.  Na-

tional, sectoral and technological innovation systems: The case of Taiwanese pharmaceutical biotechnology and agricultural biotechnology innovation systems (1945–2000). Science and Public Policy, 39(2): 271–281, March 2012. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Castle:2013:GNM

[CC13] David Castle and Keith Culver. Getting to ‘no’: The method of contested exchange. Science and Public Policy, 40(1):34–42, February 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Chang:2016:SDC

[CCF16] Xuhua Chang, Qiang Chen, and Patrick S. W. Fong. Scientific disclosure and commercialization mode selection for university technology transfer. Science and Public Policy, 43(1):85–101, February 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/43/1/85.

Ciarli:2016:QAT

[CCR16] Tommaso Ciarli, Alex Coad, and Ismael Rafols. Quantitative analysis of technology futures: a review of techniques, uses and characteristics. Science and Public Policy, 43(5):630–??, October 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/
5/630/2725282/Quantitative-
analysis-of-technology-
futures-A.

Capellari:2016:AIA

[CD16] Saveria Capellari and Domenico De Stefano. Academic inventors, allocation of patent rights and knowledge diffusion: Subnetwork structures in university-owned and university-invented patents in two Italian universities. Science and Public Policy, 43(5):585–??, October 2016. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/
5/585/2770708/Academic-
inventors-allocation-of-
patent-rights-and.

Chataway:2019:SGC

[CDD^{+}19] Joanna Chataway, Charlie Dobson, Chux Daniels, Rob Byrne, Rebecca Hanlin, and Aschalew Tigabu. Science granting councils in Sub-Saharan Africa: Trends and tensions. Science and Public Policy, 46(4):620–631, August 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/4/620/5416172.

CarvalhodeMello:2016:ISI

[CDI16] José Manoel Carvalho de Mello, Claudia De Fuentes, and Donato Iacobucci. Introduction to the special issue: Universities as interactive partners. Science and Public Policy, 43(5):581– ??, October 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/
5/581/2525534/Introduction-
to-the-special-issue-Universities-
as.

Campo-Engelstein:2012:CEG

[CETRW12] Lisa Campo-Engelstein, Candace Tingen, Sarah Rodriguez, and Teresa K.

Woodruff. Conceiving ethical gamete and embryo research in a post-Dickey-Wicker USA. Science and Public Policy, 39 (1):129–132, February 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Canibano:2016:GPT

[CFO16] Carolina Cañibano, Mary Frank Fox, and F. Javier Otamendi. Gender and patterns of temporary mobility among researchers. Science and Public Policy, 43(3):320–331, June 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Casault:2013:EBR

[CGL13] Sébastien Casault, Aard J. Groen, and Jonathan D. Linton. Examination of the behavior of R&D returns using a power law. Science and Public Policy, 40(2):219–228, April 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Clark-Ginsberg:2019:RRW

[CGS19] Aaron Clark-Ginsberg and Rebecca Slayton. Regulating risks within complex sociotechnical systems: Evidence from critical infrastructure cybersecurity standards. Science and Public Policy, 46(3):339–346, June 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/3/339/5184558.

Clark:2010:IIR

[CH10] Rebecca Clark and John Holmes. Improving input from research to environmental policy: challenges of structure and culture. Science and Public Policy, 37(10):751–764, December 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Christensen:2017:ACA

[CH17] Johan Christensen and Cathrine Holst. Advisory commissions, academic expertise and democratic legitimacy: the case of Norway. Science and Public Policy, 44 (6):821–833, December 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

//academic.oup.com/spp/
article/44/6/821/3817193.

Charles:2016:RUC

[Cha16] David Charles. The rural university campus and support for rural innovation. Science and Public Policy, 43(6):763– ??, December 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/
6/763/2525518/The-rural-
university-campus-and-support-
for-rural.

Cremonini:2018:DRS

[CHH18] Leon Cremonini, Edwin Horlings, and Laurens K Hessels. Different recipes for the same dish: Comparing policies for scientific excellence across different countries. Science and Public Policy, 45(2):232–245, April 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/2/232/4562547.

Choi:2019:IRD

[Cho19] Hyundo Choi. Incumbents’ response to demand-side policies: The case of solar and wind power sectors. Science and Public Policy, 46 (5):648–660, October 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/46/5/648/5485749.

Ciumasu:2010:TBD

[Ciu10] Dr. Ioan M. Ciumasu. Turning brain drain into brain networking. Science and Public Policy, 37(2):135–146, March 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cho:2015:DEI

[CKS15] Kyungchul Cho, Changseok Kim, and Juneseuk Shin. Differential effects of intellectual property rights on innovation and economic performance: a cross-industry investigation. Science and Public Policy, 42(6):827–840, December 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cai:2015:RUF

[CL15] Yuzhuo Cai and Cui Liu. The roles of universities in fostering knowledge-intensive clusters in Chinese regional innovation systems. Science and Public Policy, 42(1):15– 29, February 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Clark:2010:EGA

[Cla10] Benjamin Y. Clark. The effects of government, academic and industrial policy on crossuniversity collaboration. Science and Public Policy, 37(5): 314–330, June 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Clark:2013:BRBa

[Cla13a] Lisa F. Clark. Book review:

GM Food on Trial: Testing European Democracy by Les Levidow and Susan Carr. Science and Public Policy, 40 (6):818–819, December 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Clark:2013:BRBb

[Cla13b] Lisa F. Clark. Book review:

Seeds, Science, and Struggle: The Global Politics of Transgenic Crops by Abby Kinchy.

Science and Public Policy, 40 (6):820–821, December 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Clark:2013:FUR

[Cla13c] Lisa F. Clark. Framing the uncertainty of risk: Models of governance for genetically modified foods. Science and Public Policy, 40(4):479–491, August 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Clayton:2015:BRU

[Cla15] Paige A. Clayton. Book review: Understanding the public procurement of innovation: Public Procurement and Innovation: The Role of Institutions by Max Rolfstam. Science and Public Policy, 42(5):738–739, October 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/42/5/738.

Chatterjee:2018:URC

[CLDS18] Debabrata Chatterjee, Liudvika Leisyte, S Dasappa, and Balram Sankaran. University research commercialization in emerging economies: a glimpse into the ‘black box’. Science and Public Policy, 45(3):361–372, June 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/3/361/4557533.

Chicot:2018:PPI

[CM18] Julien Chicot and Mireille Matt. Public procurement of innovation: a review of rationales, designs, and contributions to grand challenges. Science and Public Policy, 45(4):480–492, August 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/4/480/4913820.

Chiarini:2014:SDS

[COC14] Tulio Chiarini, Vanessa Parreiras Oliveira, and Fabio Chaves do Couto e Silva Neto. Spatial distribution of scientific activities: an exploratory analysis of Brazil, 2000–10. Science and Public Policy, 41(5):625–640, October 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Coletti:2010:TIC

[Col10] Michele Coletti. Technology and industrial clusters: how different are they to manage? Science and Public Policy, 37 (9):679–688, November 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Conley:2017:ASI

[Con17] Shannon N. Conley. Assessing the societal implications of emerging technologies: Anticipatory governance in practice. Science and Public Policy, 44(6):877–879, December 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

//academic.oup.com/spp/
article/44/6/877/3860941.

Cook:2014:IFT

[Coo14] Peta S. Cook. Institutional frameworks and terms of reference: The public discussion on clinical xenotransplantation in Australia. Science and Public Policy, 41(5):673–684, October 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cozzens:2012:SCG

[Coz12] Susan E. Cozzens. Social cohesion at the global level: The roles of science and technology. Science and Public Policy, 39(5):557–561, October 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cuntz:2015:ODN

[CP15] Alexander Cuntz and Jan Peuckert. Openness determinants of national research funding programmes in EU27. Science and Public Policy, 42(4):474–486, August 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Castellacci:2016:TNI

[CP16] Fulvio Castellacci and Saira Pons. Towards a new innovation policy in Cuba: Proposal for the introduction of a R&D fiscal incentive program. Science and Public Policy, 43(5):680– ??, October 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/
5/680/2725284/Towards-a-
new-innovation-policy-in-
Cuba-Proposal.

ChavesdoCoutoeSilvaNeto:2013:PIB

[CPP^{+}13] Fábio Chaves do Couto e Silva Neto, Ulisses Pereira dos Santos, Vanessa Parreiras Oliveira, Priscila Gomes de Castro, Luiza Teixeira de Melo Franco, and Fernanda de Negri. Patterns of interaction between national and multinational corporations and Brazilian universities/public research institutes. Science and Public Policy, 40(3):281–292, June 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cho:2017:CAR

[CPSL17] Chanwoo Cho, Sun Young Park, Jong Ku Son, and Sungjoo Lee. Comparative analysis of R&D-based innovation capabilities in SMEs to design innovation policy. Science and Public Policy, 44(3):

403–??, June 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/44/
                                        3/403/2725290/Comparative-
                                        Analysis-of-R-amp-D-Based-
                                        Innovation.

Cincera:2010:FCR

[CR10] Michele Cincera and Julien Ravet. Financing constraints and R&D investments of large corporations in Europe and the US. Science and Public Policy, 37(6):455–466, July 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Caliari:2014:BPI

[CR14] Thiago Caliari and Ricardo Machado Ruiz. Brazilian pharmaceutical industry and generic drugs policy: Impacts on structure and innovation and recent developments. Science and Public Policy, 41(2):245–256, April 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /spp.oxfordjournals.org/
                                        content/41/2/245.full.pdf+
                                        html.

Crosson:2013:IES

[Cro13] Scott Crosson. The impact of empowering scientific advisory committees to constrain catch limits in US fisheries. Science and Public Policy, 40(2):261– 273, April 2013. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Chiarini:2017:AKC

[CRS17] Tulio Chiarini, Marcia Siqueira Rapini, and Leandro Alves Silva. Access to knowledge and catch-up: Exploring some intellectual property rights data from Brazil and South Korea. Science and Public Policy, 44(1):95– ??, February 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
1/95/2525541/Access-to-
knowledge-and-catch-up-
Exploring-some.

Calof:2010:CSF

[CS10] Jonathan Calof and Jack E.

Smith. Critical success factors for government-led foresight. Science and Public Policy, 37(1):31–40, February 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cronin:2014:BBH

[CS14] Blaise Cronin and Cassidy R.

Sugimoto, editors. Beyond bibliometrics: harnessing multidimensional indicators of scholarly impact. MIT Press, Cambridge, MA, USA, 2014. ISBN 0-262-52551-8 (paperback), 0-262-02679-1 (hardcover). viii + 466 pp. LCCN Z669.8 .B49 2014.

Chatterjee:2015:CAR

[CS15] Debabrata Chatterjee and Balram Sankaran. Commercializing academic research in emerging economies: Do organizational identities matter? Science and Public Policy, 42(5):599–613, October 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /spp.oxfordjournals.org/
                                        content/42/5/599.

Cozzens:2014:IIE

[CT14] Susan E. Cozzens and Dhanaraj Thakur, editors. Innovation and inequality: emerging technologies in an unequal world. Edward Elgar, Cheltenham, UK, 2014. ISBN 1-78195-166-7 (hardcover), 1-78195-167-5 (e-book). xi + 344 pp. LCCN HC79.T4 I56 2014.

Cadenas:2012:TRP

[CV12] María Sonsiré López Cadenas and Hebe Vessuri. Tensions and resistances in the political alignment of public research within Venezuela’s new political setup. Science and Public Policy, 39(5):602–612, October 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Canibano:2018:ERE

[CVCBA18] Carolina Cañibano, Immaculada Vilardell, Carmen Corona, and Carlos Benito-Amat. The evaluation of research excellence and the dynamics of knowledge production in the humanities: The case of history in Spain. Science and Public Policy, 45(6):775–789, December 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/6/775/4925395.

Choi:2019:AEN

[CZ19] Hyeri Choi and Hangjung Zo. Assessing the efficiency of national innovation systems in developing countries. Science and Public Policy, 46(4):530–540, August 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/4/530/5372390.

Dutrenit:2010:CBI

[DA10] Gabriela Dutrénit and Valeria Arza. Channels and benefits of interactions between public research organisations and industry: comparing four Latin American countries. Science and Public Policy, 37(7):541– 553, August 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Francelino:2019:HPP

[dAFUFd19] Josiane de Araújo Francelino, Ligia Maria Soto Urbina, André Tosi Furtado, and Milton de Freitas Chagas, Jr. How public policies have shaped the technological progress in the Brazilian aeronautics industry: Embraer case. Science and Public Policy, 46(6):787–804, December 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/46/6/787/5549866.

Dagnino:2012:WST

[Dag12] Renato Dagnino. Why science and technology capacity building for social development? Science and Public Policy, 39(5):548–556, October 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Danguy:2017:GIP

[Dan17] Jérôme Danguy. Globalization of innovation production: a patent-based industry analysis. Science and Public Policy, 44(1):75–??, February 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/44/
                                        1/75/2525531/Globalization-
                                        of-innovation-production-
                                        A-patent.

Das:2015:BRU

[Das15] Anup Kumar Das. Book review: Understanding entrepreneurship teaching in a nutshell Teaching Entrepreneurship: a Practicebased Approach by Heidi M.

Neck, Patricia G. Greene and Candida G. Brush. Science and Public Policy, 42 (5):739–741, October 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/42/5/739.

Das:2017:UDG

[Das17a] Anup Kumar Das. Understanding the dynamics of global entrepreneurship. Science and Public Policy, 44(2): 300–??, April 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
2/300/2525555/Understanding-
the-dynamics-of-global.

Das:2017:UEI

[Das17b] Anup Kumar Das. Understanding the economics of the Internet and the economic processes of the e-businesses. Science and Public Policy, 44(5):723–725, October 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

//academic.oup.com/spp/
article/44/5/723/3814578.

Daye:2019:MIH

[Day19] Christian Dayé. Models of innovation: The history of an idea. Science and Public Policy, 46(4):633–634, August 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/46/4/633/5418569.

Dore:2010:CW

[DB10] Mohammed H. I. Dore and Geoff Black. Custodians of water. Science and Public Policy, 37(7):554–556, August 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

DeMoor:2010:UDI

[DBD^{+}10] Katrien De Moor, Katrien Berte, Lieven De Marez, Wout Joseph, Tom Deryckere, and Luc Martens. User-driven innovation? Challenges of user involvement in future technology analysis. Science and Public Policy, 37(1):51–61, February 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

deBakker:2014:RRI

[dBdLHB14] Erik de Bakker, Carolien de Lauwere, Anne-Charlotte Hoes, and Volkert Beekman. Responsible research and innovation in miniature: Information asymmetries hindering a more inclusive ‘nanofood’ development. Science and Public Policy, 41(3):294–305, June 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Duch-Brown:2018:IPS

[DBdPR18] Néstor Duch-Brown, Andrea de Panizza, and Ibrahim Kholilul Rohman. Innovation and productivity in a scienceand-technology intensive sector: Information industries in Spain. Science and Public Policy, 45(2):175–190, April 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/2/175/4992966.

Donovan:2013:CPI

[DBO13] Amy R. Donovan, Michael Bravo, and Clive Oppenheimer. Co-production of an institution: Montserrat Volcano Observatory and social dependence on science. Science and Public Policy, 40(2): 171–186, April 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

daCosta:2016:CPE

[dC16] Janaina Oliveira Pamplona da Costa. Can PPI enhance public policy? Science and Public Policy, 43(6):879–??, December 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/
6/879/2525540/Can-PPI-enhance-
public-policy.

deCampos:2017:IEU

[dC17] André Luiz Sica de Campos. An international exploration of university–industry links. Science and Public Policy, 44(5):726–727, October 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

                                        //academic.oup.com/spp/
                                        article/44/5/726/3835593.

Dorland:2019:SCE

[DCJ19] Jens Dorland, Christian Clausen, and Michael Søgaard Jørgensen. Space configurations for empowering university-community interactions. Science and Public Policy, 46(5):689–701, October 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/46/5/689/5489340.

Marques:2012:OPS

[dCM12] Ivan da Costa Marques. Ontological politics and situated public policies. Science and Public Policy, 39(5):570–578, October 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dahaj:2018:RCPa

[DCT18a] Arash Soleimani Dahaj, Brian P. Cozzarin, and Kambiz Talebi. Revisiting the Canadian public policy towards venture capital: Crowding-out or displacement. Science and Public Policy, 45(5):719–730, October 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/5/719/4911086.

Dahaj:2018:RCPb

[DCT18b] Arash Soleimani Dahaj, Brian P.

Cozzarin, and Kambiz Talebi. Revisiting the Canadian public policy towards venture capital: Crowding-out or displacement. Science and Public Policy, 45(5):747, October 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/5/747/5060714.

Dutrenit:2010:CIB

[DDT10] Gabriela Dutrénit, Claudia De Fuentes, and Arturo Torres. Channels of interaction between public research organisations and industry and their benefits: evidence from Mexico. Science and Public Policy, 37(7):513–526, August 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Delina:2016:BRT

[Del16] Laurence L. Delina. Book review: Transforming U.S. Energy Innovation edited by Laura Diaz Anadon, Matthew Bunn and Venkatesh Narayanamurti. Science and Public Policy, 43(5):728–??, October 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/
5/728/2525502/Transforming-
U-S-Energy-Innovation-edited-
by-Laura.

Dias:2011:IED

[Dia11] Dr. Rafael Dias. Innovation and economic development: the impact of ICT in Latin America. Science and Public Policy, 38(10):820–821, December 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Medina-Arellano:2012:CSG

[dJMA12] María de Jesús Medina-Arellano. Contested secularity: Governing stem cell science in Mexico. Science and Public Policy, 39(3):386–402, June 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

deJong:2016:SRS

[dJSvD16] Stefan P. L. de Jong, Jorrit Smit, and Leonie van Drooge. Scientists’ response to societal impact policies: a policy paradox. Science and Public Policy, 43(1):102–114, February 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /spp.oxfordjournals.org/
                                        content/43/1/102.

Decker-Lange:2019:BEL

[DLSS19] Carolin Decker-Lange, Marie-Noëlle Singer, and Florian Schrader. Balancing evolving logics: Business model change in the Leibniz research museums. Science and Public Policy, 46(3):430–440, June 2019. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/3/430/5203409.

Duchene:2010:RSI

[DLV10] Vincent Duchêne, Elissavet Lykogianni, and Arnold Verbeek. R&D in services industries and the EU–US R&D investment gap. Science and Public Policy, 37(6):443–453, July 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

deMatos:2018:WOD

[dMCL18] Marcelo Pessoa de Matos, José Eduardo Cassiolato, and Helena M. M. Lastres. Windows of opportunity and death valleys in STI policy construction: Policy dialogues in Brazil. Science and Public Policy, 45(3):318–328, June 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/3/318/4096829.

Fagundes:2014:CMA

[dMFPP14] Paloma de Mattos Fagundes, Ana Claudia Machado Padilha, and Antonio Domingos Padula. Co-marketing alliances in the Brazilian biofuel sector: an analysis based on strategic alliances. Science and Public Policy, 41(5):553– 560, October 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dosso:2018:TEB

[DMMPC18] Mafini Dosso, Ben R. Martin, and Pietro Moncada-Paternò-Castello. Towards evidencebased industrial research and innovation policy. Science and Public Policy, 45(2):143–150, April 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/2/143/4992967.

Dominique:2013:IBP

[DMRJ13] Kathleen C. Dominique, Ammar Anees Malik, and Valerie Remoquillo-Jenni. International benchmarking: Politics and policy. Science and Public Policy, 40(4):504–513, August 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dutrenit:2018:DPD

[DN18] Gabriela Dutrénit and José Miguel Natera. Dialogue processes for designing STI policies: The creation of a thematic network. Science and Public Policy, 45(3):291–292, June 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/3/291/4102812.

Dutrenit:2018:DPS

[DNA^{+}18] Gabriela Dutrénit, José Miguel Natera, Martín Puchet Anyul, Alexandre O. Vera-Cruz, and Arturo Torres. Dialogue processes on STI policy-making in Latin America and the Caribbean: dimensions and conditions. Science and Public Policy, 45(3):293–308, June 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/3/293/4139722.

Donovan:2016:RSC

[DO16] Amy Donovan and Clive Oppenheimer. Resilient science: The civic epistemology of disaster risk reduction. Science and Public Policy, 43(3):363– 374, June 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

denOudendammer:2017:LDD

[dOB17] Willemijn M. den Oudendammer and Jacqueline E. W. Broerse. Lyme disease in the Dutch policy context: patient consultation in government research agenda setting. Science and Public Policy, 44(3): 303–??, June 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
3/303/3807113/Lyme-disease-
in-the-Dutch-policy-context-
patient.

Doganova:2013:TET

[Dog13] Liliana Doganova. Transfer and exploration: Two models of science-industry intermediation. Science and Public Policy, 40(4):442–452, August 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Doom:2018:LAM

[Doo18] Travis Doom. The long arm of Moore’s Law: Microelectronics and American science. Science and Public Policy, 45(2):289–290, April 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/2/289/4772230.

Dosi:2010:CFS

[Dos10] Giovanni Dosi. Christopher Freeman: a social scientist and a progressive intellectual. Science and Public Policy, 37(8):631–635, October 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dosi:2012:EOI

[Dos12] Giovanni Dosi. Economic organisation, industrial dynamics and development: selected essays. Edward Elgar, Cheltenham, UK, 2012. ISBN 1-84720-913-0 (hardcover). lvi + 763 pp. LCCN HD30.22 .D67 2012.

Derrick:2013:DRE

[DP13] Gemma E. Derrick and Vincenzo Pavone. Democratising research evaluation: Achieving greater public engagement with bibliometrics-informed peer review. Science and Public Policy, 40(5):563–575, October 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

deRaymond:2018:AAC

[dR18] Antoine Bernard de Raymond. ‘Aligning activities’: coordination, boundary activities, and agenda setting in interdisciplinary research. Science and Public Policy, 45(5):621–633, October 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/5/621/4794724.

Dragojlovic:2014:VSC

[Dra14] Nick Dragojlovic. Voting for stem cells: How local conditions tempered moral opposition to Proposition 71. Science and Public Policy, 41(3):359– 369, June 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

DEste:2018:HDR

[DRVWA18] Pablo D’Este, Irene Ramos-Vielba, Richard Woolley, and Nabil Amara. How do researchers generate scientific and societal impacts? Toward an analytical and operational framework. Science and Public Policy, 45(6):752–763, December 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/6/752/4925528.

Datta:2011:UIS

[DS11] Dr. Surja Datta and Professor Mohammed Saad. University and innovation systems: the case of India. Science and Public Policy, 38(1):7–17, February 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

DSilva:2013:SRS

[D’S13] Joel D’Silva. Standards, risk and the scientist. Science and Public Policy, 40(5):701–703, October 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Derrick:2018:EDD

[DS18a] G. E. Derrick and G. S. Samuel. Exploring the degree of delegated authority for the peer review of societal impact. Science and Public Policy, 45(5):673–682, October 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/5/673/4819248.

Dutrenit:2018:ISP

[DS18b] Gabriela Dutrénit and Marcela Suárez. Involving stakeholders in policymaking: tensions emerging from a public dialogue with knowledge-based entrepreneurs. Science and Public Policy, 45(3):338–350, June 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/3/338/4111196.

Delborne:2013:PPP

[DSB^{+}13] Jason Delborne, Jen Schneider, Ravtosh Bal, Susan Cozzens, and Richard Worthington. Policy pathways, policy networks, and citizen deliberation: Disseminating the results of World Wide Views on Global Warming in the USA. Science and Public Policy, 40(3):378–392, June 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Donina:2017:IGI

[DSP17] Davide Donina, Marco Seeber, and Stefano Paleari. Inconsistencies in the governance of interdisciplinarity: the case of the Italian higher education system. Science and Public Policy, 44(6):865–875, December 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

//academic.oup.com/spp/
article/44/6/865/3752473.

Damrongchai:2010:FTA

[DSTS10] Nares Damrongchai, Ponpiboon Satangput, Greg Tegart, and Chatri Sripaipan. Future technology analysis for biosecurity and emerging infectious diseases in Asia–Pacific. Science and Public Policy, 37(1): 41–50, February 2010. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dachs:2014:IBR

[DSZ14] Bernhard Dachs, Robert Stehrer, and Georg Zahradnik, editors. The internationalisation of business R&D. New perspectives on the modern corporation. Edward Elgar, Cheltenham, UK, 2014. ISBN 1-78347-089-5 (hardcover), 1-78347-090-9 (ebook). xiv + 204 pp. LCCN HD30.4 .I68 2014.

Degelsegger:2011:PPC

[DT11] Alexander Degelsegger and Helge Torgersen. Participatory paternalism: citizens’ conferences in Austrian technology governance. Science and Public Policy, 38(5):391– 402, June 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dunlop:2010:ECT

[Dun10] Claire A. Dunlop. Epistemic communities and two goals of delegation: hormone growth promoters in the European Union. Science and Public Policy, 37(3):205–217, April 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dutrenit:2010:ISI

[Dut10] Gabriela Dutrénit. Introduction to special issue: Interactions between public research organisations and industry in Latin America: a study on channels and benefits from the perspective of firms and researchers. Science and Public Policy, 37(7):471–472, August 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Edler:2018:NGI

[EB18] Jakob Edler and Wouter P.

Boon. ‘the next generation of innovation policy: Directionality and the role of demand-oriented instruments’

— introduction to the special section. Science and Public Policy, 45(4):433–434, August 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/4/433/4915398.

Edelenbos:2017:EIC

[EBV17] Jurian Edelenbos, Nanny Bressers, and Lieselot Vandenbussche. Evolution of interdisciplinary collaboration: what are stimulating conditions? Science and Public Policy, 44(4):451–??, August 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
4/451/4035807/Evolution-
of-interdisciplinary-collaboration-
what.

Engel-Glatter:2014:DUR

[EG14] Sabrina Engel-Glatter. Dualuse research and the H5N1 bird flu: Is restricting publication the solution to biosecurity issues? Science and Public Policy, 41(3):370–383, June 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Egbetokun:2013:BRB

[Egb13] Abiodun Egbetokun. Book review: Building National and Regional Innovation Systems: Institutions for Economic Development by Jorge Niosi. Science and Public Policy, 40(2): 279–280, April 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Egbetokun:2014:CRM

[Egb14] Abiodun Egbetokun. Book review: Conceptual Richness and Methodological Diversity in Entrepreneurship Research edited by Alain Fayolle, Paula Kyrö, Tõnis Mets and Urve Venesaar. Science and Public Policy, 41(4):549–551, July 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Engel-Glatter:2018:LSV

[EGI18] Sabrina Engel-Glatter and Marcello Ienca. Life scientists’ views and perspectives on the regulation of dual-use research of concern. Science and Public Policy, 45(1):92–102, February 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/1/92/4158227.

Einsiedel:2011:CCC

[EJB11] Edna F. Einsiedel, Mavis Jones, and Meaghan Brierley. Cultures, contexts and commitments in the governance of controversial technologies: US, UK and Canadian publics and xenotransplantation policy development. Science and Public Policy, 38(8):619–628, October 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Eklof:2011:SSC

[Ekl11] Jenny Eklöf. Success story or cautionary tale? Swedish ethanol in co-existing sciencepolicy frameworks. Science and Public Policy, 38(10):795– 806, December 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Elzinga:2012:FCS

[Elz12] Aant Elzinga. Features of the current science policy regime: Viewed in historical perspective. Science and Public Policy, 39(4):416–428, August 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Enger:2018:CCN

[Eng18] Simen G. Enger. Closed clubs:

Network centrality and participation in Horizon 2020. Science and Public Policy, 45(6):884–896, December 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/6/884/4982651.

Escobar:2014:UPE

[Esc14] Oliver Escobar. Upstream public engagement, downstream policy-making? The Brain Imaging Dialogue as a community of inquiry. Science and Public Policy, 41(4):480– 492, July 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Espinosa:2013:BRA

[Esp13] Juan Espinosa. Book review: Accepting infernal alternatives: Sustaining venture capital on venture labour. Venture Labor: Work and the Burden of Risk in Innovative Industries by Gina Neff. Science and Public Policy, 40(4): 560–561, August 2013. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Esko:2019:ASI

[ET19] Terhi Esko and Juha Tuunainen. Achieving the social impact of science: an analysis of public press debate on urban development. Science and Public Policy, 46(3):404– 414, June 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/3/404/5203408.

Evans:2015:BRB

[Eva15] Sam Weiss Evans. Book review: Vulnerability in Technological Cultures: New Directions in Research and Governance edited by Anique Hommels, Jessica Mesman and Wiebe E. Bijker. Science and Public Policy, 42(4):594–595, August 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fitjar:2019:TRR

[FBA19] Rune Dahl Fitjar, Paul Benneworth, and Bjørn Terje Asheim. Towards regional responsible research and innovation? Integrating RRI and RIS3 in European innovation policy. Science and Public Policy, 46(5):772–783, October 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/5/772/5491609.

Feeney:2014:EWF

[FBB14] Mary K. Feeney, Margarita Bernal, and Lauren Bowman. Enabling work? Familyfriendly policies and academic productivity for men and women scientists. Science and Public Policy, 41(6):750–764, December 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Frenken:2016:SFU

[FBS16] Koen Frenken, Rudi Bekkers, and Negin Salimi. Success factors in university–industry PhD projects. Science and Public Policy, 43(6):812–??, December 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/
6/812/2525454/Success-factors-
in-university-industry-
PhD.

Fernandez-Carro:2016:ECP

[FCLG16] Remo Fernández-Carro and Víctor Lapuente-Giné. The Emperor’s clothes and the Pied Piper: Bureaucracy and scientific productivity. Science and Public Policy, 43(4):546– 561, August 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fernandes:2010:AIL

[FdSdS^{+}10] A. C. Fernandes, B. Campello de Souza, A. Stamford da Silva, W. Suzigan, C. V. Chaves, and E. Albuquerque. Academy-industry links in Brazil: evidence about channels and benefits for firms and researchers. Science and Public Policy, 37(7):485–498, August 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ferguson:2016:BRC

[Fer16] Sean Ferguson. Book review: Chemicals without Harm: Policies for a Sustainable World by Ken Geiser. Science and Public Policy, 43 (3):430–431, June 2016. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fernandez-Esquinas:2011:EFC

[FERV11] Dr. Manuel Fernández-Esquinas and Irene Ramos-Vielba. Emerging forms of crosssector collaboration in the Spanish innovation system. Science and Public Policy, 38(2):135–146, March 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fonseca:2017:CST

[FFF17] Bruna P. F. Fonseca, Elton Fernandes, and Marcus V. A. Fonseca. Collaboration in science and technology organizations of the public sector: a network perspective. Science and Public Policy, 44(1):37– ??, February 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
1/37/2525503/Collaboration-
in-science-and-technology.

Friedewald:2013:ISS

[FG13] Michael Friedewald and Serge Gutwirth. Introduction to special section: Governing privacy and data protection issues of emerging technologies. Science and Public Policy, 40 (6):705–707, December 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Filippetti:2016:SSI

[FG16] Andrea Filippetti and Frederick Guy. Skills and social insurance: Evidence from the relative persistence of innovation during the financial crisis in Europe. Science and Public Policy, 43(4):505–517, August 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fitzpatrick:2013:BRB

[Fit13] Susan M. Fitzpatrick. Book review: Better Doctors, Better Patients, Better Decisions: Envisioning Health Care 2020 edited by Gerd Girgerenzer and J. A. Muir Gray. Science and Public Policy, 40(6):830– 831, December 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fitzpatrick:2017:FMF

[Fit17] Susan M. Fitzpatrick. Follow the money: Funding research in a large academic health center. Science and Public Policy, 44(5):722–723, October 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

                                        //academic.oup.com/spp/
                                        article/44/5/722/3071670.

Fitzpatrick:2018:BRU

[Fit18a] Susan M. Fitzpatrick. Book review: Universities and the Entrepreneurial Ecosystem by David B. Audretsch and Albert N. Link. Science and Public Policy, 45(4):591–592, August 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/4/591/4855969.

Fitzpatrick:2018:UEEb

[Fit18b] Susan M. Fitzpatrick. Universities and the entrepreneurial ecosystem. Science and Public Policy, 45(4):595, August 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/4/595/4984519.

Foray:2010:SCI

[FL10] Dominique Foray and Stéphane Lhuillery. Structural changes in industrial R&D in Europe and the US: towards a new model? Science and Public Policy, 37(6):401–412, July 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fogelberg:2013:IAE

[FL13] Hans Fogelberg and Mats A.

Lundqvist. Integration of academic and entrepreneurial roles: The case of nanotechnology research at Chalmers University of Technology. Science and Public Policy, 40 (1):127–139, February 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fernandez-Lopez:2019:FMF

[FLCRP19] Sara Fernández-López, Nuria Calvo, and David Rodeiro-Pazos. The funnel model of firms’ R&D cooperation with universities. Science and Public Policy, 46(1):45–54, February 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/46/1/45/4973763.

Flower:2014:CEB

[Flo14] Michael Flower. Choreographing the end of the beginning of human stem cell research: an ethically intricate affair. Good Science: The Ethical Choreography of Stem Cell Science by Charis Thompson. Science and Public Policy, 41(6):838– 839, December 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fisher:2015:HLR

[FM15] E. Fisher and G. Maricle. Higher-level responsiveness? Socio-technical integration within US and UK nanotechnology research priority setting. Science and Public Policy, 42(1):72–85, February 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fitzpatrick:2016:LEP

[FM16] Susan M. Fitzpatrick and James S. McDonnell. Leveling the entrepreneurial playing field: Women’s Entrepreneurship in the 21st Century: an International Multi-Level Research Analysis edited by Kate V. Lewis, Colette Henry, Elizabeth J. Gatewood and John Watson. Science and Public Policy, 43(2):289–290, April 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Forlano:2010:HC

[For10] Laura Forlano. Hail to the chef. Science and Public Policy, 37(2):163–164, March 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ferretti:2018:REI

[FPVH18] Federico Ferretti, Ângela Guimarães Pereira, Dániel Vértesy, and Sjoerd Hardeman. Research excellence indicators: time to reimagine the ‘making of’? Science and Public Policy, 45(5):731–741, October 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/5/731/4858431.

Frank:2012:WPW

[Fra12] Arthur W. Frank. Book review: Who pays what price for bioscientific discovery? A Rich Bioethics: Public Policy, Biotechnology, and the Kass Council by Adam Briggle Private Bodies, Public Texts: Race, Gender, and a Cultural Bioethics by Karla F. C. Holloway. Science and Public Policy, 39(3):404–407, June 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Flink:2010:SDI

[FS10] Tim Flink and Ulrich Schreiterer. Science diplomacy at the intersection of S&T policies and foreign affairs: toward a typology of national approaches. Science and Public Policy, 37(9):665–677, November 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fagerberg:2011:NIS

[FS11] Jan Fagerberg and Koson Sapprasert. National innovation systems: the emergence of a new approach. Science and Public Policy, 38(9):669–679, November 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Frishammar:2019:KBP

[FSHM19] Johan Frishammar, Patrik Söderholm, Hans Hellsmark, and Johanna Mossberg. A knowledge-based perspective on system weaknesses in technological innovation systems. Science and Public Policy, 46(1):55–70, February 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/46/1/55/4989362.

Freeman:2011:ABR

[FSK11] Julia Freeman, Terre Satterfield, and Milind Kandlikar. Agricultural biotechnology and regulatory innovation in India. Science and Public Policy, 38(4):319–331, May 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fischer:2018:UGE

[FSS18] Bruno Brandão Fischer, Paola Rücker Schaeffer, and Julia Phaiffer Silveira. Universities’ gravitational effects on the location of knowledge-intensive investments in Brazil. Science and Public Policy, 45 (5):692–707, October 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/5/692/4818299.

Fogelberg:2012:RIP

[FT12] Hans Fogelberg and Stefan Thorpenberg. Regional innovation policy and publicprivate partnership: The case of Triple Helix Arenas in Western Sweden. Science and Public Policy, 39(3):347–356, June 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fursov:2017:MIW

[FT17] Konstantin Fursov and Thomas Wolfgang Thurner. Make it work!

— a study of user innovation in Russia. Science and Public Policy, 44(3):392– ??, June 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
3/392/2730505/Make-it-work-
a-study-of-user-innovation-
in-Russia.

Forsberg:2014:AES

[FTNdB14] Ellen-Marie Forsberg, Erik Thorstensen, Rasmus Øjvind Nielsen, and Erik de Bakker. Assessments of emerging science and technologies: Mapping the landscape. Science and Public Policy, 41(3):306– 316, June 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Faber:2016:IBR

[FvDvR16] Jan Faber, Jaco van Dijk, and Frank van Rijnsoever. Incentives and barriers for R&Dbased SMEs to participate in European research programs: an empirical assessment for The Netherlands. Science and Public Policy, 43(3):414–428, June 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fischer:2014:EIP

[FWRF14] Arnout R. H. Fischer, Meike T. A. Wentholt, Gene Rowe, and Lynn J. Frewer. Expert involvement in policy development: a systematic review of current practice. Science and Public Policy, 41(3):332– 343, June 2014. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Fusi:2019:BFA

[FWS19] Federica Fusi, Eric W. Welch, and Michael Siciliano. Barriers and facilitators of access to biological material for international research: The role of institutions and networks. Science and Public Policy, 46(2):275–289, April 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/2/275/5114417.

Galabova:2012:DKB

[Gal12] Lidia P. Galabova. Developing a knowledge-based economy through innovation policy: The cases of Bulgaria, Finland and Scotland. Science and Public Policy, 39(6):802– 814, December 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Galaz:2014:GEG

[Gal14] Victor Galaz. Global environmental governance, technology and politics: the Anthropocene Gap. Edward Elgar, Cheltenham, UK, 2014. ISBN 1-78195-554-9 (hardcover), 1-78195-555-7 (e-book). xiv + 189 pp. LCCN GE149 .G35 2014.

Garcia:2019:HBR

[GAM^{+}19] Renato Garcia, Veneziano Araújo, Suelene Mascarini, Emerson G. Santos, and Ariana R. Costa. How the benefits, results and barriers of collaboration affect university engagement with industry. Science and Public Policy, 46(3):347–357, June 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/3/347/5145037.

Gault:2012:UIM

[Gau12] Fred Gault. User innovation and the market. Science and Public Policy, 39(1):118–128, February 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gertner:2015:ESC

[GB15] Drew Gertner and Bart A. G. Bossink. The evolution of science concentrations: The case of Newcastle Science City. Science and Public Policy, 42 (1):121–138, February 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gaughan:2019:IIU

[GB19] Monica Gaughan and Barry Bozeman. Institutionalized inequity in the USA: The case of postdoctoral researchers. Science and Public Policy, 46(3):358–368, June 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/3/358/5232675.

Griessler:2011:CIK

[GBH11] Erich Griessler, Peter Biegelbauer, and Janus Hansen. Citizens’ impact on knowledgeintensive policy: introduction to a special issue. Science and Public Policy, 38(8):583–588, October 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Galdon-Clavell:2013:SCD

[GC13] Gemma Galdon-Clavell. (not so) smart cities?: The drivers, impact and risks of surveillance-enabled smart environments. Science and Public Policy, 40(6):717–723, December 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Geiser:2015:CHP

[Gei15] Ken Geiser. Chemicals without harm: policies for a sustainable world. Unban and industrial environments. MIT Press, Cambridge, MA, USA, 2015. ISBN 0-262-01252-9 (hardcover), 0-262-51206-8 (paperback). xii + 444 pp. LCCN TP155.2.E58 G324 2015.

Galushko:2014:TFY

[GG14] Viktoriya Galushko and Richard Gray. Twenty five years of private wheat breeding in the UK: Lessons for other countries. Science and Public Policy, 41(6):765–779, December 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ghosh:2018:CRD

[Gho18]  Ritwick Ghosh.  Climate-

resilient development: Linking climate adaptation and economic development. Science and Public Policy, 45 (2):286–287, April 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/2/286/4191366.

Gibbs:2013:BRB

[Gib13] Beverley Gibbs. Book review: Low Carbon Communities: Imaginative Approaches to Combating Climate Change Locally edited by Michael Peters, Shane Fudge and Tim Jackson. Science and Public Policy, 40(6):815–816, December 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Giest:2017:CEC

[Gie17]  Sarah Giest.  The chal-

lenges of enhancing collaboration in life science clusters: Lessons from Chicago, Copenhagen and Singapore. Science and Public Policy, 44(2): 163–??, April 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
2/163/2525549/The-challenges-

of-enhancing-collaboration-
in-life.

Gulbrandsen:2010:CBR

[GK10] Magnus Gulbrandsen and Svein Kyvik. Are the concepts basic research, applied research and experimental development still useful? An empirical investigation among Norwegian academics. Science and Public Policy, 37(5):343– 353, June 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gok:2010:BBO

[Gök10]  Abdullah Gök.  Building

blocks of organisational behaviour. Science and Public Policy, 37(4):311–312, May 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldstein:2017:WWW

[Gol17] Anna P. Goldstein. Why are we waiting?: The logic, urgency, and promise of tackling climate change. Science and Public Policy, 44 (6):880–882, December 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

//academic.oup.com/spp/
article/44/6/880/4058410.

Gopakumar:2014:BRB

[Gop14] Govind Gopakumar. Book review: Public–Private Innovation Networks in Services edited by Faïz Gallouj, Luis Rubalcaba and Paul Windrum. Science and Public Policy, 41(3):401–402, June 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gouyon:2011:WDP

[Gou11] Jean-Baptiste Gouyon. What to do with politicized science? Science and Public Policy, 38 (5):417–419, June 2011. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Geum:2013:SAP

[GP13] Youngjung Geum and Yongtae Park. The state-of-the-art of public-sector technology roadmaps: a taxonomical approach to energy technology roadmaps. Science and Public Policy, 40(3):327–339, June 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gray:2011:CSR

[Gra11] Dr. Denis O. Gray. Crosssector research collaboration in the USA: a national innovation system perspective. Science and Public Policy, 38(2): 123–133, March 2011. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Graddy-Reed:2017:IAI

[GRLR17] Alexandra Graddy-Reed, Lauren Lanahan, and Nicole M. V. Ross. Influences of academic institutional factors on R&D funding for graduate students. Science and Public Policy, 44 (6):834–854, December 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

//academic.oup.com/spp/
article/44/6/834/3769305.

Gross:2010:IRD

[Gro10] Matthias Gross. Ignorance, research and decisions about abandoned opencast coal mines. Science and Public Policy, 37(2):125–134, March 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gronning:2011:BMV

[Grø11a] Terje Grønning. Business models and value creation within the biopharmaceutical industry. Science and Public Policy, 38(6):504–508, July 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Groves:2011:PEN

[Gro11b] Christopher Groves. Public engagement and nanotechnology in the UK: restoring trust or building robustness? Science and Public Policy, 38 (10):783–793, December 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gronning:2014:BRC

[Grø14] Terje Grønning. Book review: Creating and sustaining knowledge-intensive entrepreneurship. Managing Knowledge Intensive Entrepreneurship by Maureen McKelvey and Astrid Heidemann Lassen. Science and Public Policy, 41(1):137–138, February 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Grunwald:2019:IDN

[Gru19] Armin Grunwald. The inherently democratic nature of technology assessment. Science and Public Policy, 46 (5):702–709, October 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/46/5/702/5488504.

Gerritsen:2013:KGE

[GST13] Alwin L. Gerritsen, Marian Stuiver, and Catrien J. A. M. Termeer. Knowledge governance: an exploration of principles, impact, and barriers. Science and Public Policy, 40(5):604–615, October 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gutzeit:2016:BRP

[Gut16] Stephan Gutzeit. Book review: Practical reason: Rational Action: The Sciences of Policy in Britain and America, 1940–1960 by William Thomas, MIT Press, Cambridge, MA, 2015, 399 pages, US$38.00 (hardback), ISBN 978-0-262-02850-9. Science and Public Policy, 43(4):575– 576, August 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gutzeit:2018:MAA

[Gut18] Stephan Gutzeit. Missing in action: Agency. Science and Public Policy, 45(1): 135–137, February 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/1/135/4036275.

Girones:2019:TMD

[GvEV19] Edgar Salas Gironés, Rinie van Est, and Geert Verbong. Transforming mobility: The Dutch smart mobility policy as an example of a transformative STI policy. Science and Public Policy, 46(6):820–833, December 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/6/820/5542646.

Hansen:2011:AIC

[HA11] Janus Hansen and Agnes Allansdottir. Assessing the impacts of citizen participation in science governance: exploring new roads in comparative analysis. Science and Public Policy, 38(8):609–617, October 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Haber:2014:M

[Hab14] Jonathan Haber. MOOCs. MIT Press, Cambridge, MA, USA, 2014. ISBN 0-262-52691-3 (paperback). xiv + 227 pp. LCCN LB1044.87 .H27 2014.

Hagendijk:2010:ISD

[Hag10] Rob Hagendijk. Innovation studies and developing countries. Science and Public Policy, 37(10):798–801, December 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hagen:2018:EOP

[Hag18] Ryan Hagen. The end of ownership: Personal property in the digital economy. Science and Public Policy, 45(1):137–139, February 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/1/137/4065203.

Hallonsten:2011:CLS

[Hal11] Dr. Olof Hallonsten. A classic laboratory study in science policy clothing. Science and Public Policy, 38(1):79– 80, February 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hallonsten:2015:URB

[Hal15] Olof Hallonsten. Unpreparedness and risk in Big Science policy: Sweden and the European spallation source. Science and Public Policy, 42(3):415–426, June 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/42/3/415.

Hallonsten:2017:BRT

[Hal17] Olof Hallonsten. Book review:

Tunnel Visions: The Rise and Fall of the Superconducting Super Collider. Science and Public Policy, 44(1):149– ??, February 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
1/149/2566747/Tunnel-Visions-
The-Rise-and-Fall-of-the.

Hallonsten:2018:DTT

[Hal18] Olof Hallonsten. Development and transformation of the third sector of R&D in Sweden, 1942–2017. Science and Public Policy, 45 (5):634–644, October 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/5/634/4745839.

Hansson:2010:DPR

[Han10] Dr. Finn Hansson. Dialogue in or with the peer review?

Evaluating research organizations in order to promote organizational learning. Science and Public Policy, 37(4):239– 251, May 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Harvey:2011:PID

[Har11] Dr. Olivia Harvey. Platforms of innovation: dynamics of new industrial knowledge flows. Science and Public Policy, 38(10):817–818, December 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hart:2013:ETL

[Har13] David M. Hart. Energy technology learning curves: To the max. Science and Public Policy, 40(2):275–276, April 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hart:2014:AMA

[Har14] David M. Hart. An agent, not a mole: Assessing the White House Office of Science and Technology Policy. Science and Public Policy, 41(4):411– 418, July 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hashim:2011:CGP

[Has11] Noor Azuan Hashim. Ecommerce and government policy initiatives for Malaysian SMEs: the need for assessment. Science and Public Policy, 38(10):807–816, December 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hawkins:2012:BRB

[Haw12] Richard Hawkins. Book review: Innovation strategies for a Global Economy: Development, Implementation, Measurement and Management by Fred Gault. Science and Public Policy, 39(1):133– 134, February 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hawkins:2015:MME

[Haw15] Richard Hawkins. Book review: Marianna Mazzucato, The Entrepreneurial State: Debunking Public vs Private Sector Myths. Science and Public Policy, 42(1):143–145, February 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hawkins:2017:NDI

[Haw17] Richard Hawkins. A new direction for innovation studies? Reconciling the ordinary and the extraordinary. Science and Public Policy, 44(2): 294–??, April 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
2/294/3065357/A-new-direction-
for-innovation-studies-
Reconciling.

Heimeriks:2016:HSS

[HB16] Gaston Heimeriks and Pierre-Alexandre Balland. How smart is specialisation? An analysis of specialisation patterns in knowledge production. Science and Public Policy, 43(4):562–574, August 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Herstad:2010:NIP

[HBEvdV10] Dr. Sverre J. Herstad, Dr. Carter Bloch, Bernd Ebersberger, and Dr. Els van de Velde. National innovation policy and global open innovation: exploring balances, tradeoffs and complementarities. Science and Public Policy, 37(2):113– 124, March 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Horbach:2019:ORA

[HBM19] Serge P. J. M. Horbach, Eric Breit, and Svenn-Erik Mamelund. Organisational responses to alleged scientific misconduct: Sensemaking, sensegiving, and sensehiding. Science and Public Policy, 46(3):415–429, June 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/46/3/415/5230895.

Huzair:2013:IDL

[HBRUM13] Farah Huzair, Alexander Borda-Rodriguez, Mary Upton, and Julius T. Mugwagwa. An interdisciplinary and development lens on knowledge translation. Science and Public Policy, 40(1):43–50, February 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hewitt:2017:MTR

[HDT17] Michael Hewitt, Robert Dingwall, and Ilke Turkmendag. More than research intermediaries: a descriptive study of the impact and value of learned societies in the UK social sciences. Science and Public Policy, 44 (6):775–788, December 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

//academic.oup.com/spp/
article/44/6/775/3748296.

Howells:2011:SIT

[HE11] Professor Jeremy Howells and Professor Jakob Edler. Structural innovations: towards a unified perspective? Science and Public Policy, 38(2):157– 167, March 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Heimeriks:2013:IBG

[Hei13] Gaston Heimeriks. Interdisciplinarity in biotechnology, genomics and nanotechnology. Science and Public Policy, 40(1):97–112, February 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Heinecke:2018:RTR

[Hei18] Steffi Heinecke. On the route towards renewal? The Polish Academy Of Sciences in postsocialist context. Science and Public Policy, 45(2):246–256, April 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/2/246/4557155.

Hellstrom:2018:CEC

[Hel18] Tomas Hellström. Centres of excellence and capacity building: from strategy to impact. Science and Public Policy, 45(4):543–552, August 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/4/543/4652938.

Hemphill:2012:BSU

[Hem12] Thomas A. Hemphill. The biotechnology sector and US gene patents: Legal challenges to intellectual property rights and the impact on basic research and development. Science and Public Policy, 39 (6):815–826, December 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hemmert:2017:KAU

[Hem17] Martin Hemmert. Knowledge acquisition by university researchers through company collaborations: Evidence from South Korea. Science and Public Policy, 44(2):199– ??, April 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
2/199/2525556/Knowledge-
acquisition-by-university-
researchers.

Henderson:2019:PEC

[Hen19]  Dylan Henderson.  Policy

entrepreneurship in context: Understanding the emergence of novel policy solutions for services innovation in Finland and Ireland. Science and Public Policy, 46(5):668–678, October 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/5/668/5485750.

Hayter:2017:DEP

[HF17] Christopher S. Hayter and Mary K. Feeney. Determinants of external patenting behavior among university scientists. Science and Public Policy, 44(1):111–??, February 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
1/111/2725287/Determinants-
of-external-patenting-behavior-
among.

Hardeman:2015:CCI

[HFNW15] Sjoerd Hardeman, Koen Frenken, Önder Nomaler, and Anne L. J. Ter Wal. Characterizing and comparing innovation systems by different ‘modes’ of knowledge production: a proximity approach. Science and Public Policy, 42(4):530–548, August 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hickey:2013:MES

[HFS^{+}13] Gordon M. Hickey, Patrick Forest, Jean L. Sandall, Briony M. Lalor, and Rodney J. Keenan. Managing the environmental science–policy nexus in government: Perspectives from public servants in Canada and Australia. Science and Public Policy, 40(4): 529–543, August 2013. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Huang:2011:OID

[HFW11] Wan-Ling Huang, Mary K. Feeney, and Eric W. Welch. Organizational and individual determinants of patent production of academic scientists and engineers in the United States. Science and Public Policy, 38(6):463–479, July 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Haak:2012:ESP

[HFW^{+}12] Laurel L. Haak, Will Ferriss, Kevin Wright, Michael E. Pollard, Kirk Barden, Matt A. Probus, Michael Tartakovsky, and Charles J. Hackett. The electronic Scientific Portfolio Assistant: Integrating scientific knowledge databases to support program impact assessment. Science and Public Policy, 39(4):464–475, August 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hickman:2019:AMC

[HFWL19] Charles F. Hickman, Eric A.

Fong, Allen W. Wilhite, and Yeolan Lee. Academic misconduct and criminal liability: Manipulating academic journal impact factors. Science and Public Policy, 46(5):661–667, October 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/5/661/5488509.

Hessels:2011:ECI

[HGS11] Dr. Laurens K. Hessels, John Grin, and Ruud E. H. M. Smits. The effects of a changing institutional environment on academic research practices: three cases from agricultural science. Science and Public Policy, 38(7):555–568, August 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hallonsten:2012:IPT

[HH12] Olof Hallonsten and Thomas Heinze. Institutional persistence through gradual organizational adaptation: Analysis of national laboratories in the USA and Germany. Science and Public Policy, 39(4):450– 463, August 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hallonsten:2013:PPP

[HH13] Olof Hallonsten and Thomas Heinze. From particle physics to photon science: Multidimensional and multi-level renewal at DESY and SLAC. Science and Public Policy, 40(5):591–603, October 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hallonsten:2015:FEN

[HH15] Olof Hallonsten and Thomas Heinze. Formation and expansion of a new organizational field in experimental science. Science and Public Policy, 42 (6):841–854, December 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Haegeman:2010:ISS

[HHJ10] Karel Haegeman, Dr. Jennifer C. Harper, and Ron Johnston. Introduction to a special section: Impacts and implications of future-oriented technology analysis for policy and decision-making. Science and Public Policy, 37(1):3–6, February 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hicks:2012:HES

[Hic12] Diana Hicks. Book review:

How Economics Shapes Science by Paula Stephan. Science and Public Policy, 39 (6):840–841, December 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Higginbotham:2016:ASE

[Hig16] Brian D. Higginbotham. Assessing the state of entrepreneurship and innovation research. Science and Public Policy, 43(6):877–??, December 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/
6/877/2525530/Assessing-
the-state-of-entrepreneurship-
and.

Himmelsbach:2016:BRP

[Him16] Raffael Himmelsbach. Book review: The Planet Remade: How Geoengineering Could Change the World by Oliver Morton. Science and Public Policy, 43(6):872–??, December 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/

                                        6/872/2525528/The-Planet-
                                        Remade-How-Geoengineering-
                                        Could-Change.

Himmelsbach:2018:HSA

[Him18] Raffael Himmelsbach. How scientists advising the European Commission on research priorities view climate engineering proposals. Science and Public Policy, 45(1):124–133, February 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/1/124/4158232.

Hellstrom:2017:TOP

[HJS17] Tomas Hellström, Merle Jacob, and Karolin Sjöö. From thematic to organizational prioritization: the challenges of implementing RDI priorities. Science and Public Policy, 44(5):599–608, October 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

                                        //academic.oup.com/spp/
                                        article/44/5/599/2907804.

Hirsch-Kreinsen:2014:KIE

[HKS14] Hartmut Hirsch-Kreinsen and Isabel Schwinge, editors. Knowledge-intensive entrepreneurship in low-tech industries. Edward Elgar, Cheltenham, UK, 2014. ISBN 1-78347-203-0 (hardcover). xiii + 249 pp. LCCN HB615 .K56 2014.

Horta:2011:HDS

[HL11] Hugo Horta and T. Austin Lacy. How does size matter for science? Exploring the effects of research unit size on academics’ scientific productivity and information exchange behaviors. Science and Public Policy, 38(6):449–460, July 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Harris:2019:ACN

[HL19] Richard Harris and Trinh Le.

Absorptive capacity in New Zealand firms: Measurement and importance. Science and Public Policy, 46(2):290–309, April 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/2/290/5133283.

Hlavka:2017:PSC

[Hlá17] Jakub P. Hlávka. Portfolio society: On the capitalist mode of prediction. Science and Public Policy, 44(6):879–880, December 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

//academic.oup.com/spp/
article/44/6/879/3897315.

Harmon:2013:GRE

[HLH13] Shawn H. E. Harmon, Graeme Laurie, and Gill Haddow. Governing risk, engaging publics and engendering trust:

New horizons for law and social science? Science and Public Policy, 40(1):25–33, February 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hansson:2011:FPP

[HLI^{+}11] Kristofer Hansson, Susanne Lundin, Markus Idvall, Jekaterina Kaleja, and Aivita Putnina. Framing the public: the policy process around xenotransplantation in Latvia and Sweden 1970–2004. Science and Public Policy, 38(8):629– 637, October 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hall:2014:RRD

[HLL14] Michael J. Hall, Stephen K. Layson, and Albert N. Link. The returns to R&D: Division of Policy Research and Analysis at the National Science Foundation. Science and Public Policy, 41(4):458–463, July 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Howlett:2010:ELV

[HM10] M. Howlett and A. Migone. Explaining local variation in agri-food biotechnology policies: ‘green’ genomics regulation in comparative perspective. Science and Public Policy, 37(10):781–795, December 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hodge:2014:NRR

[HMB14a] Graeme A. Hodge, Andrew D.

Maynard, and Diana M. Bowman. Nanotechnology: Rhetoric, risk and regulation. Science and Public Policy, 41 (1):1–14, February 2014. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hommels:2014:VTC

[HMB14b] Anique Hommels, Jessica Mesman, and Wiebe E. Bijker, editors. Vulnerability in technological cultures: new directions in research and governance. Inside technology. MIT Press, Cambridge, MA, USA, 2014. ISBN 0-262-02710-0 (hardcover), 0-262-52580-1 (paperback). xi + 382 pp. LCCN T14.5 .V85 2014.

Heink:2015:CCR

[HMH^{+}15] Ulrich Heink, Elisabeth Marquard, Katja Heubach, Kurt Jax, Carolin Kugel, Carsten Neßhöver, Rosmarie K. Neumann, Axel Paulsch, Sebastian Tilch, Johannes Timaeus, and Marie Vandewalle. Conceptualizing credibility, relevance and legitimacy for evaluating the effectiveness of science–policy interfaces: Challenges and opportunities. Science and Public Policy, 42(5):676–689, October 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /spp.oxfordjournals.org/
                                        content/42/5/676.

Hennen:2015:NWT

[HN15] Leonhard Hennen and Linda Nierling. A next wave of technology assessment? Barriers and opportunities for establishing TA in seven European countries. Science and Public Policy, 42(1):44–58, February 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hillman:2011:FST

[HNRM11] Karl Hillman, Måns Nilsson, Annika Rickne, and Thomas Magnusson. Fostering sustainable technologies: a framework for analysing the governance of innovation systems. Science and Public Policy, 38 (5):403–415, June 2011. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ho:2019:AES

[HO19] Jae-Yun Ho and Eoin O’Sullivan. Addressing the evolving standardisation challenges of ‘smart systems’ innovation: Emerging roles for government? Science and Public Policy, 46(4):552–569, August 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/46/4/552/5418570.

Holbrook:2010:IPS

[Hol10] J. Adam Holbrook. Innovating in the public sector services? Science and Public Policy, 37(2):161–162, March 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Holbrook:2015:IHH

[Hol15] Adam Holbrook. Book review: The Innovator’s Hypothesis: How Cheap Experiments Are Worth More Than Good Ideas by Michael Schrage. Science and Public Policy, 42(2):290–291, April 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/42/2/290.

Hempel:2013:TSI

[HOSV13] Leon Hempel, Lars Ostermeier, Tobias Schaaf, and Dagny Vedder. Towards a social impact assessment of security technologies: a bottom-up approach. Science and Public Policy, 40(6):740–754, December 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hoy:2013:BRB

[Hoy13] Frank Hoy. Book review:

Handbook of Research on Innovation and Entrepreneurship, edited by David B. Audretsch, Oliver Falck, Stephan Heblich and Adam Lederer.

Science and Public Policy, 40 (2):274–275, April 2013. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Huzair:2012:UBC

[HP12] Farah Huzair and Theo Papaioannou. UK Biobank: Consequences for commons and innovation. Science and Public Policy, 39(4):500–512, August 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Holl:2012:TSB

[HR12] Adelheid Holl and Ruth Rama. Technology sourcing: Are biotechnology firms different? An exploratory study of the Spanish case. Science and Public Policy, 39(3):304– 317, June 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Holl:2019:LCI

[HR19] Adelheid Holl and Ruth Rama. Local cooperation for innovation in ICT — domestic groups with collaborations for innovation abroad and foreign subsidiaries. Science and Public Policy, 46(4):599–610, August 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/46/4/599/5423864.

Hrotic:2014:BRB

[Hro14]  Steven Hrotic.  Book re-

view: Creativity and Entrepreneurship: Changing Currents in Education and Public Life edited by Lynn Book and David P. Phillips. Science and Public Policy, 41(2):261–263, April 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/41/2/261.full.pdf+
html.

Huang:2016:GTL

[HS16] Can Huang and Naubahar Sharif. Global technology leadership: The case of China. Science and Public Policy, 43(1):62–73, February 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/43/1/62.

Hassan:2016:TII

[HSM16] Saeed-Ul Hassan, Raheem Sarwar, and Amina Muazzam. Tapping into intraand international collaborations of the Organization of Islamic Cooperation states across science and technology disciplines. Science and Public Policy, 43(5):690–??, October 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/

                                        5/690/2725285/Tapping-into-
                                        intra-and-international.

Huang:2017:ICR

[Hua17]  Hsini Huang.  Invisible

constraints: The relationship among non-competition agreements, inventor mobility, and patent commercialization. Science and Public Policy, 44(3):341–??, June 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/44/
                                        3/341/2525574/Invisible-
                                        Constraints-The-Relationship-
                                        among-Non.

Hubbard:2014:FBN

[Hub14] Christopher Hubbard. Fukushima and beyond: nuclear power in a low-carbon world. Ashgate, Farnham, Surrey, England, 2014. ISBN 1-4094-5491-6 (hardcover), 1-4094-5492-4 (e-book), 1-4724-0730-X (epub). xii + 202 pp. LCCN TK9153 .H83 2014.

Huggins:2013:EEM

[Hug13]  Laura E. Huggins.  En-

vironmental entrepreneurship: markets meet the environment in unexpected places. Edward Elgar Pub. Ltd., Cheltenham, UK, 2013. ISBN 1-78195-396-1, 1-78195-397-X (e-book). v + 215 pp. LCCN HD60 .H84 2013.

Harsh:2018:RETa

[HWC^{+}18a] Matthew Harsh, Thomas S.

Woodson, Susan Cozzens, Jameson M. Wetmore, Ogundiran Soumonni, and Rodrigo Cortes. The role of emerging technologies in inclusive innovation: the case of nanotechnology in South Africa. Science and Public Policy, 45(5):597–607, October 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/5/597/4711134.

Harsh:2018:RETb

[HWC^{+}18b] Matthew Harsh, Thomas S.

Woodson, Susan Cozzens, Jameson M. Wetmore, Ogundiran Soumonni, and Rodrigo Cortes. The role of emerging technologies in inclusive innovation: the case of nanotechnology in South Africa. Science and Public Policy, 45(5):746, October 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/5/746/5060715.

Ivancheva:2011:CCM

[IG11] Ludmila Ivancheva and Elissaveta Gourova. Challenges for career and mobility of researchers in Europe. Science and Public Policy, 38(3):185– 198, April 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Isabelle:2011:MSI

[IH11] Diane A. Isabelle and Louise A. Heslop. Managing for success in international scientific collaborations: views from Canadian government senior science managers. Science and Public Policy, 38(5):349–364, June 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Izsak:2017:ERI

[IR17] Kincsö Izsak and Slavo Radošević. EU research and innovation policies as factors of convergence or divergence after the crisis. Science and Public Policy, 44(2): 274–??, April 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/44/
                                        2/274/2725289/EU-Research-
                                        and-Innovation-Policies-
                                        as-Factors-of.

Iskandarova:2017:ISI

[Isk17] Marfuga Iskandarova. From the idea of scale to the idea of agency: an actornetwork theory perspective on policy development for renewable energy. Science and Public Policy, 44(4):476– ??, August 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/44/
                                        4/476/2567815/From-the-

idea-of-scale-to-the-idea-
of-agency-An.

Jackson:2012:BRB

[Jac12] Paul Jackson. Book review:

Red, Black, and Objective: Science, Sociology, and Anarchism by Sal Restivo. Science and Public Policy, 39(6):839– 840, December 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jeong:2015:CRA

[JC15] Seongkyoon Jeong and Jae Young Choi. Collaborative research for academic knowledge creation: How team characteristics, motivation, and processes influence research impact. Science and Public Policy, 42(4):460–473, August 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Janssen:2018:SIC

[JC18] Matthijs J. Janssen and Carolina Castaldi. Services, innovation, capabilities, and policy: Toward a synthesis and beyond. Science and Public Policy, 45(6):863–874, December 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/6/863/4904189.

Jonkers:2010:IPS

[JCC10] Koen Jonkers and Laura Cruz-Castro. The internationalisation of public sector research through international joint laboratories. Science and Public Policy, 37(8):559–570, October 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jeong:2014:DIC

[JCK14] Seongkyoon Jeong, Jae Young Choi, and Jang-Yun Kim. On the drivers of international collaboration: The impact of informal communication, motivation, and research resources. Science and Public Policy, 41(4):520–531, July 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jensen:2014:TCD

[Jen14] Eric Allen Jensen. The therapeutic cloning debate: global science and journalism in the public sphere. Ashgate, Farnham, Surrey, UK, 2014. ISBN 1-4094-2982-2 (hardcover), 1-4094-2983-0 (e-book), 1-4724-0380-0 (e-pub). vi + 240 pp. LCCN P96.H84 J37 2014.

Jacob:2012:PBS

[JH12] Merle Jacob and Olof Hallonsten. The persistence of big science and megascience in research and innovation policy. Science and Public Policy, 39(4):411–415, August 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jungblut:2017:ADA

[JJ17] Jens Jungblut and Marc Jungblut. All different? All equal? Differentiation of universities’ mission statements and excellence initiatives in Germany. Science and Public Policy, 44(4):535– ??, August 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
4/535/2738672/All-different-
All-equal-differentiation-
of.

Jung:2019:DEB

[JL19] Euy-Young Jung and Xielin Liu. The different effects of basic research in enterprises on economic growth: Income-level quantile analysis. Science and Public Policy, 46(4):570–588, August 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/4/570/5369805.

Jackson:2018:UICa

[JMSS18a] Paul Jackson, Reza Kiani Mavi, Yuliani Suseno, and Craig Standing. University– industry collaboration within the triple helix of innovation: The importance of mutuality. Science and Public Policy, 45(1):142, February 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/1/142/4812026.

Jackson:2018:UICb

[JMSS18b] Paul Jackson, Reza Kiani Mavi, Yuliani Suseno, and Craig Standing. University– industry collaboration within the triple helix of innovation: The importance of mutuality. Science and Public Policy, 45(4):553–564, August 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/4/553/4712015.

Johnsen:2016:EEA

[JNP16] Hans Christian Garmann Johnsen, Roger Normann, and Rómulo Pinheiro. External engagement and the academic heartland: The case of a regionally-embedded university. Science and Public Policy, 43(6):787–??, December 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/43/
                                        6/787/2525522/External-
                                        engagement-and-the-academic-
                                        heartland-The.

Johnson:2011:UPR

[Joh11] Björn Johnson. From userproducer relations to the learning economy. Science and Public Policy, 38(9):703–711, November 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Johnson:2017:BPW

[Joh17] Ericka Johnson. Big pharma, women and the labour of love. Science and Public Policy, 44(3):431–??, June 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
3/431/2525563/Big-Pharma-
Women-and-the-Labour-of-
Love.

Jensen-Ryan:2019:ESP

[JRG19] Danielle K. Jensen-Ryan and Laura A. German. Environmental science and policy: a meta-synthesis of case studies on boundary organizations and spanning processes. Science and Public Policy, 46(1):13–27, February 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/1/13/4960916.

Jonkers:2018:DIE

[JS18] Koen Jonkers and Frédérique Sachwald. The dual impact of ‘excellent’ research on science and innovation: the case of Europe. Science and Public Policy, 45(2):159–174, April 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/2/159/4992965.

Juliano:2012:FNP

[Jul12] R. L. Juliano. The future of nanomedicine: Promises and limitations. Science and Public Policy, 39(1):99–104, February 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Juliano:2013:PIP

[Jul13] R. L. Juliano. Pharmaceutical innovation and public policy: The case for a new strategy for drug discovery and development. Science and Public Policy, 40(3):393–405, June 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jacobsson:2014:MWA

[JVH14] Staffan Jacobsson, Eugenia Perez Vico, and Hans Hellsmark. The many ways of academic researchers: How is science made useful? Science and Public Policy, 41(5):641– 657, October 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kalar:2016:SCA

[KA16] Barbara Kalar and Bostjan Antoncic. Social capital of academics and their engagement in technology and knowledge transfer. Science and Public Policy, 43(5):646– ??, October 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/

spp/article-abstract/43/
5/646/2725283/Social-capital-
of-academics-and-their-
engagement.

Kolltveit:2017:DSI

[KA17] Kristoffer Kolltveit and Jostein Askim. Decentralisation as substantial and institutional policy change: scrutinising the regionalisation of science policy in Norway. Science and Public Policy, 44(4):546– ??, August 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
4/546/2870675/Decentralisation-
as-substantial-and-institutional.

Kahn:2018:CAP

[Kah18] Michael Kahn. Co-authorship as a proxy for collaboration: a cautionary tale. Science and Public Policy, 45(1):117–123, February 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/1/117/4159476.

Kahn:2019:CBS

[Kah19] Michael Kahn. The contract between science and society: a South African case study. Science and Public Policy, 46(1):116–125, February 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/1/116/5004402.

Kamath:2011:DTE

[Kam11] Anant Kamath. Does technical education in India contribute to its Core-HRST? A case study of IIT Madras. Science and Public Policy, 38(4): 293–305, May 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kamenetzky:2013:OIS

[Kam13] Julia R. Kamenetzky. Opportunities for impact: Statistical analysis of the National Science Foundation’s broader impacts criterion. Science and Public Policy, 40(1):72– 84, February 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kastrinos:2010:PCO

[Kas10] Nikos Kastrinos. Policies for co-ordination in the European research area: a view from the social sciences and humanities. Science and Public Policy, 37(4):297–310, May 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kastrinos:2013:FCG

[Kas13] Nikos Kastrinos. The financial crisis and Greek R&D policy from a Schumpeterian perspective. Science and Public Policy, 40(6):779–791, December 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kastenhofer:2017:BGS

[Kas17] Karen Kastenhofer. Bridging gaps and squaring circles: attempts at a cross-European technology assessment. Science and Public Policy, 44(5):728–729, October 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

//academic.oup.com/spp/
article/44/5/728/3858430.

Kassab:2019:DPO

[Kas19] Omar Kassab. Does public outreach impede research performance? Exploring the ‘researcher’s dilemma’ in a sustainability research center. Science and Public Policy, 46(5):710–720, October 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/5/710/5491606.

Kropp:2011:MSS

[KB11] Kristoffer Kropp and Anders Blok. Mode-2 social science knowledge production? The case of Danish sociology between institutional crisis and new welfare stabilizations. Science and Public Policy, 38(3):213–224, April 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kim:2017:IRC

[KB17] Do Han Kim and Hee-Je Bak.

Incentivizing research collaboration using performancebased reward systems. Science and Public Policy, 44(2): 186–??, April 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
2/186/2525553/Incentivizing-
research-collaboration-
using.

Kowalczewska:2019:HPD

[KB19] Katarzyna Kowalczewska and Jelle Behagel. How policymakers’ demands for usable knowledge shape sciencepolicy relations in environmental policy in Poland. Science and Public Policy, 46(3):381–390, June 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/3/381/5211077.

Kivimaa:2017:CUI

[KBA17] Paula Kivimaa, Wouter Boon, and Riina Antikainen. Commercialising university inventions for sustainability — a case study of (non- )intermediating ‘cleantech’ at aalto University. Science and Public Policy, 44(5): 631–644, October 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

//academic.oup.com/spp/
article/44/5/631/2838388.

Krick:2019:BSP

[KCH19] Eva Krick, Johan Christensen, and Cathrine Holst. Between ‘scientization’ and a ‘participatory turn’. Tracing shifts in the governance of policy advice. Science and Public Policy, 46(6):927–939, December 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/6/927/5559560.

Keenan:2012:OIS

[KCM^{+}12] Michael Keenan, Paul Cutler, John Marks, Richard Meylan, Carthage Smith, and Emilia Koivisto. Orienting international science cooperation to meet global ‘grand challenges’. Science and Public Policy, 39 (2):166–177, March 2012. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kim:2012:CUN

[KCS12] Youngjae Kim, Elizabeth A.

Corley, and Dietram A. Scheufele. Classifying US nano-scientists: Of cautious innovators, regulators, and technology optimists. Science and Public Policy, 39(1):30– 38, February 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kim:2017:NPI

[KCS17] Youngjae Kim, Elizabeth A.

Corley, and Dietram A.

Scheufele. Nanoscientists and political involvement: Which characteristics make scientists more likely to support engagement in political debates? Science and Public Policy, 44(3):317– ??, June 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
3/317/2525570/Nanoscientists-
and-political-involvement-
Which.

Kim:2014:IPV

[KCW14] Seoyong Kim, Sang-Ok Choi, and Jaesun Wang. Individual perception vs. structural context: Searching for multilevel determinants of social acceptance of new science and technology across 34 countries. Science and Public Policy, 41(1):44–57, February 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kurt-Dickson:2010:AML

[KD10] Aygen Kurt-Dickson. Analysing multi-level governance and learning dynamics in European clusters: a complex picture? Science and Public Policy, 37(8):643–644, October 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Knockaert:2019:RBO

[KDL19] Mirjam Knockaert, Matthias Deschryvere, and Laura Lecluyse.

The relationship between organizational interdependence and additionality obtained from innovation ecosystem participation. Science and Public Policy, 46(4):490–503, August 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/4/490/5309056.

Kannebley:2018:SPC

[KdLABdP18] Júnior Sérgio Kannebley, Renata de Lacerda Antunes Borges, and Diogo de Prince. Scientific production and its collective determinants: an econometric analysis for the Brazilian research labs. Science and Public Policy, 45(5):661–672, October 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/5/661/4791402.

Kaltenbrunner:2017:QOE

[KdR17] Wolfgang Kaltenbrunner and Sarah de Rijcke. Quantifying ‘output’ for evaluation: Administrative knowledge politics and changing epistemic cultures in Dutch law faculties. Science and Public Policy, 44(2):284–??, April 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
2/284/2525564/Quantifying-
Output-for-Evaluation-Administrative.

Kearnes:2013:PSW

[Kea13] Matthew Kearnes. Performing synthetic worlds: Situating the bioeconomy. Science and Public Policy, 40(4):453– 465, August 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Keelor:2013:BRB

[Kee13] Bradley Keelor. Book review: Greenhouse Governance: Addressing Climate Change in America by Barry G. Rabe. Science and Public Policy, 40 (3):420–421, June 2013. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Keelor:2016:BRI

[Kee16] Bradley Keelor. Book review: Innovating in Urban Economies: Economic Transformation in Canadian City-Regions edited by David A. Wolfe. Science and Public Policy, 43(3):429–430, June 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Keelor:2018:PSG

[Kee18] Bradley E. Keelor. Pursuing sustainability: a guide to the science and practice. Science and Public Policy, 45(3):430, June 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/3/430/4794723.

Kennedy:2016:BRC

[Ken16] Eric B. Kennedy. Book review: The Climate Resilient Organization: Adaptation and Resilience to Climate Change and Weather Extremes by Martina K. Linnenluecke and Andrew Griffiths. Science and Public Policy, 43(5):724– ??, October 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/
5/724/2525477/The-Climate-
Resilient-Organization-
Adaptation-and.

Kent:2017:BRS

[Ken17] Alex Kent. Book review:

Sustaining Biodiversity and Ecosystem Functions: Economic Issues by Clement A. Tisdell. Science and Public Policy, 44(1):146–??, February 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
1/146/2725288/Sustaining-
Biodiversity-and-Ecosystem-
Functions.

Klerkx:2017:AFR

[KG17a] Laurens Klerkx and José Guimón. Attracting foreign R&D through international centres of excellence: early experiences from Chile. Science and Public Policy, 44(6):763–774, December 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

                                        //academic.oup.com/spp/
                                        article/44/6/763/3192393.

Kruss:2017:UII

[KG17b] Glenda Kruss and Michael Gastrow. Universities and innovation in informal settings: Evidence from case studies in South Africa. Science and Public Policy, 44(1):26– ??, February 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/44/
                                        1/26/2525494/Universities-
                                        and-innovation-in-informal-
                                        settings.

Klenk:2012:ISR

[KH12a] Nicole Lisa Klenk and Gordon M. Hickey. Improving the social robustness of research networks for sustainable natural resource management: Results of a Delphi study in Canada. Science and Public Policy, 39(3):357– 372, June 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Konnola:2012:EFT

[KH12b] Totti Könnölä and Karel Haegeman. Embedding foresight in transnational research programming. Science and Public Policy, 39(2):191–207, March 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kloet:2013:UCD

[KHZ^{+}13] Roy R. Kloet, Laurens K.

Hessels, Marjolein B. M. Zweekhorst, Jacqueline E. W. Broerse, and Tjard de Cock Buning. Understanding constraints in the dynamics of a research programme intended as a niche innovation. Science and Public Policy, 40(2):206– 218, April 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kim:2014:GRF

[Kim14] So Young Kim. Government R&D funding in economic downturns: Testing the varieties of capitalism conjecture. Science and Public Policy, 41(1):107–118, February 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kingiri:2011:CAC

[Kin11] Ann Njoki Kingiri. Conflicting advocacy coalitions in an evolving modern biotechnology regulatory subsystem: policy learning and influencing Kenya’s regulatory policy process. Science and Public Policy, 38(3):199–211, April 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kim:2018:EVI

[KJ18] Leo Dhohoon Kim and Deok-Ho Jang. Expert views on innovation and bureaucratization of science: Semantic network analysis of discourses on scientific governance. Science and Public Policy, 45(1):36–44, February 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/1/36/4080734.

Kirs:2017:SBR

[KKL17] Margit Kirs, Erkki Karo, and Priit Lumi. Strategic behaviour of research groups within the entrepreneurial university policy rhetoric: the Estonian biotechnology sector. Science and Public Policy, 44(6):802–820, December 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

//academic.oup.com/spp/
article/44/6/802/3769304.

Klemens:2012:CLP

[Kle12]  Ben Klemens.  Book re-

view: Copyright Law and the Progress of Science and the Useful Arts by Alina Ng. Science and Public Policy, 39(5): 692–693, October 2012. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Klochikhin:2013:IST

[Klo13] Evgeny A. Klochikhin. Innovation system in transition: Opportunities for policy learning between China and Russia. Science and Public Policy, 40(5):657–673, October 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kashani:2013:CFP

[KM13] Ebrahim Souzanchi Kashani and Erik Millstone. Contested framings and policy controversies: Analysing biosafety policy-making in Iran. Science and Public Policy, 40(5):616– 627, October 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kundu:2019:RCF

[KM19] Oishee Kundu and Nicholas E.

Matthews. The role of charitable funding in university research. Science and Public Policy, 46(4):611–619, August 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/4/611/5418575.

Ko:2015:PIP

[Ko15] Yongsu Ko. Policy ideas and policy learning about ‘basic research’ in South Korea. Science and Public Policy, 42(4): 448–459, August 2015. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kortright:2014:BRB

[Kor14] Chris Kortright. Book review: Cosmopolitan Commons: Sharing Resources and Risks across Borders, edited by Nil Disco and Eda Kranakis. Science and Public Policy, 41(1):135–137, February 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kaiser:2010:PSE

[KPG10] Robert Kaiser and Heiko Prange-Gstöhl. A paradigm shift in European R&D policy? The EU Budget Review and the economic crisis. Science and Public Policy, 37(4): 253–265, May 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kuhlmann:2018:NGI

[KR18] Stefan Kuhlmann and Arie Rip. Next-generation innovation policy and grand challenges. Science and Public Policy, 45(4):448–454, August 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/4/448/4915392.

Krick:2015:NEP

[Kri15] Eva Krick. Negotiated expertise in policy-making: How governments use hybrid advisory committees. Science and Public Policy, 42(4):487–500, August 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kroll:2019:HEI

[Kro19a]  Henning Kroll.  How to

evaluate innovation strategies with a transformative ambition? A proposal for a structured, process-based approach. Science and Public Policy, 46(5):635–647, October 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/5/635/5490657.

Kropp:2019:BRE

[Kro19b] Kristoffer Kropp. Book review: The European Research Council by Thomas König. Science and Public Policy, 46(5):784–785, October 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/5/784/5476598.

Kaldewey:2017:PPS

[KS17] David Kaldewey and Désirée Schauz. “The politics of pure science” revisited. Science and Public Policy, 44 (6):883–886, December 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

//academic.oup.com/spp/
article/44/6/883/4587618.

Konnola:2012:FFS

[KSC^{+}12] Totti Könnölä, Ahti Salo, Cristiano Cagnin, Vicente Carabias, and Eeva Vilkkumaa. Facing the future: Scanning, synthesizing and sensemaking in horizon scanning. Science and Public Policy, 39(2):222–231, March 2012.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Knockaert:2013:UNI

[KVH13] Mirjam Knockaert, Elien Vandenbroucke, and Annelore Huyghe. Unraveling the need for innovation support services in new technology-based firms: The impact of commercialization strategy. Science and Public Policy, 40(1):85– 96, February 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kwiek:2012:CHE

[Kwi12]  Marek Kwiek.  Changing

higher education policies: From the deinstitutionalization to the reinstitutionalization of the research mission in Polish universities. Science and Public Policy, 39(5):641– 654, October 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kwiek:2018:ATE

[Kwi18] Marek Kwiek. Academic top earners. Research productivity, prestige generation, and salary patterns in European universities. Science and Public Policy, 45(1):1–13, February 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/1/1/3836941.

Kim:2017:HSC

[KWJ17] KyungWoo Kim, Eric W.

Welch, and Timothy Johnson. Human and social capital determinants of translational activity in medical sciences. Science and Public Policy, 44(5):609–619, October 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

//academic.oup.com/spp/
article/44/5/609/2949256.

Kwon:2011:CEU

[Kwo11]  Ki-Seok Kwon.  The co-

evolution of universities’ academic research and knowledgetransfer activities: the case of South Korea. Science and Public Policy, 38(6):493–503, July 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Laberge:2014:BRB

[Lab14]  Yves Laberge.  Book re-

view: Sustainable Development, Evaluation and Policy-Making: Theory, Practise and Quality Assurance edited by Anneke von Raggamby and Frieder Rubik. Science and Public Policy, 41(3):403–404, June 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Laidlaw:2017:HSP

[Lai17]  Tess Laidlaw.  Handbook

for science public information officers. Science and Public Policy, 44(3):428–??, June 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/44/
                                        3/428/2525561/Handbook-
                                        for-Science-Public-Information-
                                        Officers.

Langford:2010:ICN

[Lan10a] Cooper H. Langford. International cluster networks. Science and Public Policy, 37(8): 637–638, October 2010. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Langford:2010:RET

[Lan10b] Cooper H. Langford. A revolution in energy technologies. Science and Public Policy, 37 (3):235–236, April 2010. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Laranja:2012:NGI

[Lar12] Manuel Laranja. Network governance of innovation policies: The Technological Plan in Portugal. Science and Public Policy, 39(5):655–668, October 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Laursen:2011:UPI

[Lau11] Keld Laursen. User-producer interaction as a driver of innovation: costs and advantages in an open innovation model.

Science and Public Policy, 38 (9):713–723, November 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lyall:2013:RFA

[LBMM13] Catherine Lyall, Ann Bruce, Wendy Marsden, and Laura Meagher. The role of funding agencies in creating interdisciplinary knowledge. Science and Public Policy, 40(1):62– 71, February 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Langfeldt:2015:EGD

[LBS^{+}15] Liv Langfeldt, Mats Benner, Gunnar Sivertsen, Ernst H. Kristiansen, Dag W. Aksnes, Siri Brorstad Borlaug, Hanne Foss Hansen, Egil Kallerud, and Antti Pelkonen. Excellence and growth dynamics: a comparative study of the Matthew effect. Science and Public Policy, 42(5):661–675, October 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/42/5/661.

Liu:2018:THM

[LC18] Cui Liu and Yuzhuo Cai.

Triple helix model and institutional logics in Shenzhen Special Economic Zone. Science and Public Policy, 45 (2):221–231, April 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/2/221/4364868.

Loi:2015:TMU

[LD15] Michela Loi and Maria Chiara Di Guardo. The third mission of universities: an investigation of the espoused values. Science and Public Policy, 42(6):855–870, December 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ladwig:2012:PFF

[LDB^{+}12] Pete Ladwig, Kajsa E. Dalrymple, Dominique Brossard, Dietram A. Scheufele, and Elizabeth A. Corley. Perceived familiarity or factual knowledge? Comparing operationalizations of scientific understanding. Science and Public Policy, 39(6):761–774, December 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lehoux:2010:ECN

[LDL^{+}10] Pascale Lehoux, Geneviève Daudelin, John Lavis, Julia Abelson, Jean-Louis Denis, and F. A. Miller. Exploring the conundrum of the new knowledge production regime: an ethnographic case study on the governance and outcomes of a science/policy network in genetics. Science and Public Policy, 37(10):737–750, December 2010. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Li:2019:BTF

[LDS19] Feng Li, Jing Ding, and Wangbing Shen. Back on track: Factors influencing Chinese returnee scholar performance in the reintegration process. Science and Public Policy, 46(2):184–197, April 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/2/184/5052295.

Leisyte:2011:UCP

[Lei11] Dr. Liudvika Leisyte. University commercialization policies and their implementation in The Netherlands and the United States. Science and Public Policy, 38(6):437–448, July 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lempert:2016:BRE

[Lem16] Robert J. Lempert. Book review: Experiment Earth: Responsible Innovation in Geoengineering by Jack Stilgoe A Case for Climate Engineering by David Keith. Science and Public Policy, 43(6):873– ??, December 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/
6/873/2525529/Experiment-

                                        Earth-Responsible-Innovation-
                                        in.

Lemay:2018:CST

[Lem18] Margaret A. Lemay. Canadian science, technology and innovation policy: The innovation economy and society nexus. Science and Public Policy, 45(3):428–429, June 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/3/428/4555478.

Leslie:2013:BRB

[Les13] Christopher Leslie. Book review: Media Clusters: Spatial Agglomeration and Content Capabilities edited by Charlie Karlsson and Robert G. Picard. Science and Public Policy, 40(4):559–560, August 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Levine:2012:SSC

[Lev12]  Aaron D. Levine.  State

stem cell policy and the geographic preferences of scientists in a contentious emerging field. Science and Public Policy, 39(4):530–541, August 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Levenda:2015:BRC

[Lev15] Anthony M. Levenda. Book review: Cheap and Clean: How Americans Think About Energy in the Age of Global Warming by Stephen Ansolabehere and David M. Konisky. Science and Public Policy, 42(4):595–597, August 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lyall:2013:EIC

[LF13] Catherine Lyall and Isabel Fletcher. Experiments in interdisciplinary capacitybuilding: The successes and challenges of large-scale interdisciplinary investments. Science and Public Policy, 40(1): 1–7, February 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Li:2016:SAN

[LG16] Yanchao Li and Luke Georghiou.

Signaling and accrediting new technology: Use of procurement for innovation in China. Science and Public Policy, 43 (3):338–351, June 2016. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Langfeldt:2012:IME

[LGGK12] Liv Langfeldt, Helge Godø, Åse Gornitzka, and Aris Kaloudis. Integration modes in EU research: Centrifugality versus coordination of national research policies. Science and Public Policy, 39(1): 88–98, February 2012. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Leisyte:2011:ISI

[LH11] Dr. Liudvika Leisyte and Dr. Hugo Horta. Introduction to a special issue: Academic knowledge production, diffusion and commercialization: policies, practices and perspectives. Science and Public Policy, 38(6):422–424, July 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lalor:2013:ESP

[LH13] Briony M. Lalor and Gordon M. Hickey. Environmental science and public policy in Executive government: Insights from Australia and Canada. Science and Public Policy, 40(6):767–778, December 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lewis:2014:WEC

[LHGW14] Kate V. Lewis, Colette Henry, Elizabeth J. Gatewood, and John Watson, editors. Women’s entrepreneurship in the 21st century: an international multi-level research analysis. Edward Elgar Publishing, Cheltenham, UK, 2014. ISBN 1-78254-460-7. LCCN HD6072.5 .W665 2014.

Li:2011:MMP

[Li11] Ben Li. Microchips are made of people. Science and Public Policy, 38(3):251–252, April 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ligtvoet:2014:BRB

[Lig14] Andreas Ligtvoet. Book review: Handbook on the Economic Complexity of Technological Change edited by Cristiano Antonelli. Science and Public Policy, 41(2): 260–261, April 2014. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http://

spp.oxfordjournals.org/
content/41/2/260.full.pdf+
html.

Link:2014:ISS

[Lin14] Albert N. Link. Introduction to special section: a retrospective look at US science and technology policy. Science and Public Policy, 41(4):409– 410, July 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Link:2019:TTU

[Lin19] Albert N. Link. Technology transfer at the US National Institute of Standards and Technology. Science and Public Policy, 46(6):906–912, December 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/6/906/5543532.

Livesey:2012:BRB

[Liv12] Finbarr Livesey. Book review: Unlocking Energy Innovation: How America can build a Lowcost, Low-carbon Energy System by Richard K. Lester and David M. Hart. Science and Public Policy, 39(4):543–544, August 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lindstrom:2017:UIE

[LK17] Maria Duclos Lindstrøm and Kristoffer Kropp. Understanding the infrastructure of European Research Infrastructures — the case of the European Social Survey (ESS-ERIC). Science and Public Policy, 44(6): 855–864, December 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

//academic.oup.com/spp/
article/44/6/855/3751184.

Liefner:2016:RDP

[LKP16] Ingo Liefner, Henning Kroll, and Arman Peighambari. Research-driven or partypromoted? Factors affecting patent applications of private small and medium-sized enterprises in China’s Pearl River Delta. Science and Public Policy, 43(6):849–??, December 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/

6/849/2525469/Research-
driven-or-party-promoted-
Factors.

Lis:2019:CPE

[LKR19] Aleksandra Lis, Kärg Kama, and Leonie Reins. Coproducing European knowledge and publics amidst controversy: The EU expert network on unconventional hydrocarbons. Science and Public Policy, 46(5):721–731, October 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/5/721/5513307.

Levine:2013:OHE

[LLH13] Aaron D. Levine, T. Austin Lacy, and James C. Hearn. The origins of human embryonic stem cell research policies in the US states. Science and Public Policy, 40(4):544–558, August 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lehoux:2016:HVC

[LMDU16] P. Lehoux, F. A. Miller, G. Daudelin, and D. R. Urbach. How venture capitalists decide which new medical technologies come to exist. Science and Public Policy, 43(3):375–385, June 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Leitch:2014:FRR

[LMMD14] Shirley Leitch, Judy Motion, Elizabeth Merlot, and Sally Davenport. The fall of research and rise of innovation: Changes in New Zealand science policy discourse. Science and Public Policy, 41(1):119– 130, February 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lorenz:2011:DLM

[Lor11] Edward Lorenz. Do labour markets and educational and training systems matter for innovation outcomes? A multilevel analysis for the EU-27. Science and Public Policy, 38 (9):691–702, November 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lovbrand:2011:CPE

[Löv11] Eva Lövbrand. Co-producing European climate science and policy: a cautionary note on the making of useful knowledge. Science and Public Policy, 38(3):225–236, April 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lovering:2016:BRF

[Lov16] Jessica R. Lovering. Book review: Fukushima and Beyond: Nuclear Power in a Low-Carbon World by Christopher Hubbard, Ashgate, Farnham, UK, 2014, 214 pages, £60.00 (US$109.95) (hardcover), ISBN 978-1-4094-5491-5. Science and Public Policy, 43(1):145–146, February 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/43/1/145.

Levidow:2016:PDN

[LP16] Les Levidow and Theo Papaioannou. Policy-driven, narrative-based evidence gathering: UK priorities for decarbonisation through biomass. Science and Public Policy, 43(1):46–61, February 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/43/1/46.

Link:2012:SVQ

[LS12a] Albert N. Link and John T.

Scott. On the social value of quality: an economic evaluation of the Baldrige Performance Excellence Program. Science and Public Policy, 39(5):680–689, October 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Linkova:2012:PAP

[LS12b] Marcela Linková and Tereza Stöckelová. Public accountability and the politicization of science: The peculiar journey of Czech research assessment. Science and Public Policy, 39(5):618–629, October 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lawson:2014:REC

[LS14] Cornelia Lawson and Valerio Sterzi. The role of early-career factors in the formation of serial academic inventors. Science and Public Policy, 41(4): 464–479, July 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lawson:2015:IRV

[LS15] Cornelia Lawson and Sotaro Shibayama. International research visits and careers: an analysis of bioscience academics in Japan. Science and Public Policy, 42 (5):690–710, October 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /spp.oxfordjournals.org/
                                        content/42/5/690.

Link:2018:TAU

[LS18] Albert N. Link and John T. Scott. Toward an assessment of the US Small Business Innovation Research Program at the National Institutes of Health. Science and Public Policy, 45(1):83–91, February 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/1/83/4111198.

Llopis:2018:SEK

[LSBOPCM18] Oscar Llopis, Mabel Sánchez-Barrioluengo, Julia Olmos-Peñuela, and Elena Castro-Martínez. Scientists’ engagement in knowledge transfer and exchange: Individual factors, variety of mechanisms and users. Science and Public Policy, 45(6):790–803, December 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/6/790/4925383.

Lamers:2017:CDM

[LSKvA17] Dieuwke Lamers, Marc Schut, Laurens Klerkx, and Piet van Asten. Compositional dynamics of multilevel innovation platforms in agricultural research for development. Science and Public Policy, 44(6):739–752, December 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

//academic.oup.com/spp/
article/44/6/739/3103023.

Lengyel:2015:CRI

[LSL15] Balázs Lengyel, Tamás Sebestyén, and Loet Leydesdorff. Challenges for regional innovation policies in Central and Eastern Europe: Spatial concentration and foreign control of US patenting. Science and Public Policy, 42(1):1–14, February 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lee:2016:RSS

[LSPC16] Sungjoo Lee, Jong Ku Son, Sun Young Park, and Chanwoo Cho. R&D support services for small and mediumsized enterprises: The different perspectives of clients and service providers, and the roles of intermediaries. Science and Public Policy, 43(6):859–??, December 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/43/
                                        6/859/2525480/R-amp-D-support-
                                        services-for-small-and-
                                        medium.

Liu:2017:BCC

[LSTC17] Xielin Liu, Sylvia Schwaag Serger, Ulrike Tagscherer, and Amber Y. Chang. Beyond catch-up — can a new innovation policy help China overcome the middle income trap? Science and Public Policy, 44(5):656–669, October 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

                                        //academic.oup.com/spp/
                                        article/44/5/656/3058734.

Lin:2019:DET

[LSW19] Shoufu Lin, Ji Sun, and Shanyong Wang. Dynamic evaluation of the technological innovation efficiency of China’s industrial enterprises. Science and Public Policy, 46(2):232–243, April 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/2/232/5070427.

Lema:2016:CDW

[LSZ16a] Rasmus Lema, Ambuj Sagar, and Yuan Zhou. Convergence or divergence? Wind power innovation paths in Europe and Asia. Science and Public Policy, 43(3):400–413, June 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). See editorial remark [LSZ16b] about publication in incorrect journal issue.

Lema:2016:E

[LSZ16b] Rasmus Lema, Ambuj Sagar, and Yuan Zhou. Erratum. Science and Public Policy, 43(4): 580, August 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). See [LSZ16a].

Lander:2011:DBI

[LT11] Bryn Lander and Halla Thorsteinsdóttir. Developing biomedical innovation capacity in India. Science and Public Policy, 38(10):767–781, December 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Li:2019:WIM

[LT19] Feng Li and Li Tang. When international mobility meets local connections: Evidence from China. Science and Public Policy, 46(4):518–529, August 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/4/518/5364332.

Lindmark:2010:BRI

[LTU10] Sven Lindmark, Geomina Turlea, and Martin Ulbrich. Business R&D in the ICT sector: examining the European ICT R&D deficit. Science and Public Policy, 37(6):413– 428, July 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Levidow:2017:STC

[LU17] Les Levidow and Paul Upham. Socio-technical change linking expectations and representations: Innovating thermal treatment of municipal solid waste. Science and Public Policy, 44(2):211– ??, April 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
2/211/2525557/Socio-technical-
change-linking-expectations-
and.

Lundvall:2011:P

[Lun11] Bengt-Åke Lundvall. Publications. Science and Public Policy, 38(9):735–738, November 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lusk:2016:BRE

[Lus16] Greg Lusk. Book review:

The Ecological Footprint: New Developments in Policy and Practice, by Andrea Collins and Andrew Flynn. Science and Public Policy, 43(5):725– ??, October 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/
5/725/2525478/The-Ecological-
Footprint-New-Developments-
in.

Luukkonen:2014:ERC

[Luu14] Terttu Luukkonen. The European Research Council and the European research funding landscape. Science and Public Policy, 41(1):29–43, February 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lv:2014:HDO

[Lv14] Ping Lv. How does openness affect innovation? Evidence from national key laboratories in China. Science and Public Policy, 41(2):180–193, April 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/41/2/180.full.pdf+
html.

Loeber:2011:SLL

[LVG11] Anne Loeber, Wytske Versteeg, and Erich Griessler. Stop looking up the ladder: analyzing the impact of participatory technology assessment from a process perspective. Science and Public Policy, 38(8):599–608, October 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Long:2019:CPP

[LW19] Cheryl Xiaoning Long and Jun Wang. China’s patent promotion policies and its quality implications. Science and Public Policy, 46 (1):91–104, February 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/1/91/5004405.

Liu:2010:CCS

[LZ10] Xielin Liu and Tingting Zhi. China is catching up in science and innovation: the experience of the Chinese Academy of Sciences. Science and Public Policy, 37(5):331–342, June 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mane:2015:BRB

[Man15] Rahul S. Mane. Book review: The Internationalization of Business R&D, edited by Bernhard Dachs, Robert Stehrer and Georg Zahradnik. Science and Public Policy, 42(3):427–429, June 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/42/3/427.

Marks:2011:SCR

[Mar11] Dr. Nicola J. Marks. Stem cell researchers’ trust, ambivalence and reflexivity: opportunities for improved sciencepublic relations? Science and Public Policy, 38(7):541–554, August 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Martin:2013:BRB

[Mar13] Aaron K. Martin. Book review: Innovation, Dual Use, and Security: Managing the Risks of Emerging Biological and Chemical Weapons edited by Jonathan B. Tucker. Science and Public Policy, 40 (6):824–826, December 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Martin:2015:CFS

[Mar15] Brian Martin. Censorship and free speech in scientific controversies. Science and Public Policy, 42(3):377–386, June 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/42/3/377.

Martin:2016:TCI

[Mar16] Ben R. Martin. Twenty challenges for innovation studies. Science and Public Policy, 43 (3):432–450, June 2016. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mahroum:2013:DLR

[MAS13] Sami Mahroum and Yasser Al-Saleh. Demand-led related diversification: an innovation policy approach to economic diversification and development. Science and Public Policy, 40(3):406–418, June 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mahroum:2016:SMC

[MAS16] Sami Mahroum and Yasser Al-Saleh. The surrogate model of cluster creation: The case of Mubadala in Abu Dhabi. Science and Public Policy, 43(1):1–12, February 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /spp.oxfordjournals.org/
                                        content/43/1/1.

Matos:2010:MSC

[Mat10] Stelvia Matos. Mobilizing science: a close look at the social side of science. Science and Public Policy, 37(6):467– 468, July 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Matos:2017:NIS

[Mat17]  Stelvia Matos.  National

innovation systems, social inclusion and development:

The Latin American experience. Science and Public Policy, 44(3):432–??, June 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
3/432/2549061/National-
Innovation-Systems-Social-
Inclusion-And.

Mazzucato:2014:ESD

[Maz14] Mariana Mazzucato. The entrepreneurial state: debunking public vs. private sector myths. Anthem frontiers of global political economy; Anthem other canon economics. Anthem Press, London, UK, revised edition, 2014. ISBN 0-85728-252-2 (paperback). xxv + 237 pp. LCCN HB615 .M372797 2014.

Mejlgaard:2012:SSE

[MB12] Niels Mejlgaard and Carter Bloch. Science in society in Europe. Science and Public Policy, 39(6):695–700, December 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Martinez:2018:LBT

[MB18] Catalina Martinez and Lydia Bares. The link between technology transfer and international extension of university patents: Evidence from Spain. Science and Public Policy, 45(6):827–842, December 1, 2018. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/6/827/4903023.

Mejlgaard:2012:LSS

[MBD^{+}12] Niels Mejlgaard, Carter Bloch, Lise Degn, Mathias W. Nielsen, and Tine Ravn. Locating science in society across Europe: Clusters and consequences. Science and Public Policy, 39(6):741–750, December 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mohr:2012:MRO

[MBHD12] Alison Mohr, Helen Busby, Tamara Hervey, and Robert Dingwall. Mapping the role of official bioethics advice in the governance of biotechnologies in the EU: The European Group on Ethics’ Opinion on commercial cord blood banking. Science and Public Policy, 39(1):105–117, February 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mirnezami:2016:WDR

[MBL16] Seyed Reza Mirnezami, Catherine Beaudry, and Vincent Larivière. What determines researchers’ scientific impact? A case study of Quebec researchers. Science and Public Policy, 43(2):262–274, April 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mejlgaard:2019:RRI

[MBM19] Niels Mejlgaard, Carter Bloch, and Emil Bargmann Madsen. Responsible research and innovation in Europe: a cross-country comparative analysis. Science and Public Policy, 46(2):198–209, April 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/2/198/5055383.

McCormick:2013:SDC

[McC13] Patricia K. McCormick. Space debris: Conjunction opportunities and opportunities for international cooperation. Science and Public Policy, 40 (6):801–813, December 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mendritzki:2010:CAG

[Men10] Stefan Mendritzki. The case against growth. Science and Public Policy, 37(8):638–640, October 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Menzies:2012:RSE

[Men12] Malcolm B. Menzies. Researching scientific entrepreneurship in New Zealand. Science and Public Policy, 39(1):39– 59, February 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mercado:2012:SIS

[Mer12] Alexis Mercado. Social inclusion or social illusion: The challenges of social inclusion, social participation and social cohesion in Venezuelan S&T policy. Science and Public Policy, 39(5):592–601, October 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Meszaros:2014:WST

[Mes14] Jacqueline Meszaros. Book review: What Should Think Tanks Do?: a Strategic Guide to Policy Impact by Andrew Selee. Science and Public Policy, 41(3):402–403, June 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Milosevic:2016:VCF

[MF16] Miona Milosevic and Jacqueline Fendt. Venture capital and its French exception: Explaining performance through human capital, policy and institutional failures. Science and Public Policy, 43(5):660– ??, October 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/
5/660/2525448/Venture-capital-
and-its-French-exception.

Modic:2017:MHB

[MF17] Dolores Modic and Maryann P. Feldman. Mapping the human brain: comparing the US and EU Grand Challenges. Science and Public Policy, 44(3): 440–??, June 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
3/440/2883400/Mapping-the-
human-brain-comparing-the-
US-and-EU.

Marjanovic:2012:CBB

[MFC12] Sonja Marjanovic, Caroline Fry, and Joanna Chataway. Crowdsourcing based business models: In search of evidence for Innovation 2.0. Science and Public Policy, 39(3):318– 332, June 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mascarenhas:2018:UIC

[MFM18] Carla Mascarenhas, João J.

Ferreira, and Carla Marques. University–industry cooperation: a systematic literature review and research agenda. Science and Public Policy, 45(5):708–718, October 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/5/708/4829714.

Molyneux-Hodgson:2014:SBW

[MHB14] Susan Molyneux-Hodgson and Andrew S. Balmer. Synthetic biology, water industry and the performance of an innovation barrier. Science and Public Policy, 41(4):507–519, July 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Michelson:2011:IIN

[Mic11] Evan S. Michelson. The interdisciplinary impacts of nanotechnology: a look into the future. Science and Public Policy, 38(4):334–335, May 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Miller:2013:BRB

[Mil13] Jennifer M. Miller. Book review: The New Knowledge Workers by Dariusz Jemelniak. Science and Public Policy, 40(2):277–278, April 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). See corrigendum [Mil14].

Miller:2014:CNK

[Mil14] Jennifer M. Miller. Corrigendum to ‘The New Knowledge Workers by Dariusz Jemielniak’. Science and Public Policy, 41(2):267, April 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /spp.oxfordjournals.org/
                                        content/41/2/267.full.pdf+

html. See [Mil13].

Miller:2015:BRB

[Mil15] Jennifer M. Miller. Book review: Knowledge-Intensive Entrepreneurship in Low-Tech Industries edited by Hartmut Hirsch-Kreinsen and Isabel Schwinge. Science and Public Policy, 42(2):287–288, April 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/42/2/287.

Miscione:2011:IIU

[Mis11] Gianluca Miscione. Information infrastructures and understanding of global warming. Science and Public Policy, 38(4):332–333, May 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mote:2016:TBI

[MJH^{+}16] Jonathon Mote, Gretchen Jordan, Jerald Hage, Wilbur Hadden, and Aleia Clark. Too big to innovate? Exploring organizational size and innovation processes in scientific research. Science and Public Policy, 43(3):332–337, June 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Meyer:2013:ISS

[MK13] Morgan Meyer and Matthew Kearnes. Introduction to special section: Intermediaries between science, policy and the market. Science and Public Policy, 40(4):423–429, August 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Marcoux:2013:DRL

[ML13] Jean-Michel Marcoux and Lyne Létourneau. A distorted regulatory landscape: Genetically modified wheat and the influence of non-safety issues in Canada. Science and Public Policy, 40(4):514–528, August 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Meijer:2012:NRI

[MMGM12] Ingeborg Meijer, Jordi Molas-Gallart, and Pauline Mattsson. Networked research infrastructures and their governance: The case of biobanking. Science and Public Policy, 39(4):491–499, August 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Marques:2019:SNC

[MMHV19] Pedro Marques, Kevin Morgan, Adrian Healy, and Paul Vallance. Spaces of novelty: Can universities play a catalytic role in less developed regions? Science and Public Policy, 46(5):763–771, October 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/46/5/763/5498789.

Mote:2012:CEE

[Mot12] Jonathon Mote. Book review: Conceptualizing entrepreneurship as entrepreneuring: Entrepreneurial Imagination:

Time, Timing, Space and Place in Business Action by Björn Bjerke and Hans Rämö. Science and Public Policy, 39 (3):403–404, June 2012. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mote:2016:BRC

[Mot16] Jonathon Mote. Book review:

Corporate Venturing: Organizing for Innovation by Jessica van den Bosch and Geert Duysters. Science and Public Policy, 43(4):576–578, August 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

McLean:2012:RRM

[MP12] Craig McLean and Alan Patterson. The regulation of risk: Mobile phones and the siting of phone masts — the UK experience. Science and Public Policy, 39(6):827–836, December 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Moncada-Paterno-Castello:2010:ISI

[MPC10] Pietro Moncada-Paternò-Castello.

Introduction to a special issue: New insights on EU– US comparison of corporate R&D. Science and Public Policy, 37(6):391–400, July 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mali:2017:EFC

[MPP^{+}17] Franc Mali, Toni Pustovrh, Rok Platinovšek, Luka Kronegger, and Anuška Ferligoj. The effects of funding and coauthorship on research performance in a small scientific community. Science and Public Policy, 44(4):486– ??, August 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
4/486/2525579/The-effects-
of-funding-and-co-authorship-
on.

Muscio:2017:CRB

[MRR17] Alessandro Muscio, Laura Ramaciotti, and Ugo Rizzo. The complex relationship between academic engagement and research output: Evidence from Italy. Science and Public Policy, 44(2):235– ??, April 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
2/235/2525559/The-complex-
relationship-between-academic.

McKelvey:2018:RCR

[MSZ18] Maureen McKelvey, Rögnvaldur J. Saemundsson, and Olof Zaring. A recent crisis in regenerative medicine: Analyzing governance in order to identify public policy issues. Science and Public Policy, 45(5):608–620, October 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/5/608/4818297.

Mastroeni:2013:RIP

[MTR13] Michele Mastroeni, Joyce Tait, and Alessandro Rosiello. Regional innovation policies in a globally connected environment. Science and Public Policy, 40(1):8–16, February 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Moodysson:2017:PLS

[MTZ17] Jerker Moodysson, Michaela Trippl, and Elena Zukauskaite. Policy learning and smart specialization: balancing policy change and continuity for new regional industrial paths. Science and Public Policy, 44(3): 382–??, June 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
3/382/2525577/Policy-learning-
and-smart-specialization-
balancing.

Murmann:2013:CDI

[Mur13] Johann Peter Murmann. The co-development of industrial sectors and academic disciplines. Science and Public Policy, 40(2):229–246, April 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Musiani:2012:BRB

[Mus12] Francesca Musiani. Book review: Internet politics and STS: a much-needed encounter, The Power of Networks. Organizing the Global Politics of the Internet by Mikkel Flyverbom. Science and Public Policy, 39(3):407– 409, June 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Musiani:2016:GTC

[Mus16] Francesca Musiani. Governance of technological change: How to define it without reducing it? The Governance of Socio-Technical Systems: Explaining Change edited by Susana Borrás and Jakob Edler. Science and Public Policy, 43 (2):288–289, April 2016. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Munoz:2015:CMP

[MVFG15] V. Muñoz, F. Visentin, D. Foray, and P. Gaulé. Can medical products be developed on a non-profit basis? Exploring product development partnerships for neglected diseases. Science and Public Policy, 42(3): 315–338, June 2015. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http://

                                        spp.oxfordjournals.org/
                                        content/42/3/315.

Nehme:2012:CCO

[NdMSFC12] Claudio Chauke Nehme, Marcio de Miranda Santos, Lelio Fellows Filho, and Gilda Massari Coelho. Challenges in communicating the outcomes of a foresight study to advise decision-makers on policy and strategy. Science and Public Policy, 39(2):245– 257, March 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Neff:2018:PIU

[Nef18] Mark W. Neff. Publication incentives undermine the utility of science: Ecological research in Mexico. Science and Public Policy, 45(2):191–201, April 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/2/191/4372434.

Neff:2019:WMM

[Nef19] Mark W. Neff. Williams and Morrone misunderstand and inadvertently support my argument: Mexico’s SNI systematically steers ecological research. Science and Public Policy, 46(1):154–158, February 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/1/154/4976486.

See [WM18].

Nelson:2011:MGR

[Nel11] Richard R. Nelson. The Moon and the ghetto revisited. Science and Public Policy, 38 (9):681–690, November 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Neck:2014:TEP

[NGB14] Heidi M. Neck, Patricia G. (Patricia Gene) Greene, and Candida G. Brush. Teaching entrepreneurship: a practicebased approach. Edward Elgar Publishing, Cheltenham, UK, 2014. ISBN 1-78254-055-5 (hardcover). vii + 341 pp. LCCN HB615 .N433 2014.

Nielsen:2016:LMG

[Nie16a] Mathias W. Nielsen. Limits to meritocracy? Gender in academic recruitment and promotion processes. Science and Public Policy, 43(3):386– 399, June 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nielsen:2016:CRG

[Nie16b] Morten Velsing Nielsen. The concept of responsiveness in the governance of research and innovation. Science and Public Policy, 43(6):831–??, December 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/
6/831/2525458/The-concept-
of-responsiveness-in-the-
governance-of.

Nikzad:2015:SMS

[Nik15] Rashid Nikzad. Small and medium-sized enterprises, intellectual property, and public policy. Science and Public Policy, 42(2):176–187, April 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/42/2/176.

Njos:2018:PER

[NJ18] Rune Njøs and Stig-Erik Jakobsen. Policy for evolution of regional innovation systems: The role of social capital and regional particularities. Science and Public Policy, 45(2):257–268, April 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/2/257/4555422.

Niinikoski:2015:DNK

[NK15] Marja-Liisa Niinikoski and Stefan Kuhlmann. In discursive negotiation: Knowledge and the formation of Finnish innovation policy. Science and Public Policy, 42(1):86–106, February 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nilsson:2015:RIP

[NM15] Magnus Nilsson and Jerker Moodysson. Regional innovation policy and coordination: Illustrations from Southern Sweden. Science and Public Policy, 42(2):147–161, April 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/42/2/147.

Nishimura:2016:KRS

[NO16] Junichi Nishimura and Hiroyuki Okamuro. Knowledge and rent spillovers through government-sponsored R&D consortia. Science and Public Policy, 43(2):207–225, April 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nemethova:2019:PSF

[NSS19] Valeria Nemethova, Maria Siranova, and Miroslav Sipikal. Public support for firms in lagging regions — evaluation of innovation subsidy in Slovakia. Science and Public Policy, 46(2):173–183, April 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/2/173/5061548.

Navarro:2014:HAR

[NVAM14] Mikel Navarro, Jesús M. Valdaliso, Mari Jose Aranguren, and Edurne Magro. A holistic approach to regional strategies: The case of the Basque Country. Science and Public Policy, 41(4):532–547, July 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nakwa:2016:TMT

[NZ16] Karantarat Nakwa and Girma Zawdie. The ‘third mission’ and ‘triple helix mission’ of universities as evolutionary processes in the development of the network of knowledge production: Reflections on SME experiences in Thailand. Science and Public Policy, 43(5):622– ??, October 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/
5/622/2525535/The-third-
mission-and-triple-helix-
mission-of.

Ortega-Argiles:2010:EUD

[OAB10] Raquel Ortega-Argilés and Andries Brandsma. EU-US differences in the size of R&D intensive firms: do they explain the overall R&D intensity gap? Science and Public Policy, 37(6):429–441, July 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Olofinyehun:2018:SHQ

[OAE18] Adedayo O. Olofinyehun, Caleb M. Adelowo, and Abiodun A. Egbetokun. The supply of high-quality entrepreneurs in developing countries: evidence from Nigeria. Science and Public Policy, 45(2):269–282, April 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/2/269/4582333.

OBrien:2013:ESS

[O’B13] Timothy L. O’Brien. Education and support for scientists and elected officials in public policy decisions. Science and Public Policy, 40(3):340– 353, June 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Odumosu:2013:MWO

[Odu13] Tolu Odumosu. Making the world, one recipe at a time. Science and Public Policy, 40(5):698–699, October 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ogra:2017:WLE

[Ogr17] Anshu Ogra. A world to live in: an ecologist’s vision for a plundered planet. Science and Public Policy, 44(5):725–726, October 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

                                        //academic.oup.com/spp/
                                        article/44/5/725/3829323.

Oh:2013:GSG

[OL13] Sea-Hong Oh and Kyoung-Joo Lee. Governance system of governmental R&D programs: Formation and transformation of the Framework Act on Science and Technology in Korea. Science and Public Policy, 40(4):492–503, August 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Owen:2012:RRI

[OMS12] Richard Owen, Phil Macnaghten, and Jack Stilgoe. Responsible research and innovation: From science in society to science for society, with society. Science and Public Policy, 39(6):751–760, December 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

OBrien:2013:PIT

[OMW13] Liz O’Brien, Mariella Marzano, and Rehema M. White. ‘Participatory interdisciplinarity’: Towards the integration of disciplinary diversity with stakeholder engagement for new models of knowledge production. Science and Public Policy, 40(1):51–61, February 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Olmos-Penuela:2014:SMS

[OPBCM14] Julia Olmos-Peñuela, Paul Benneworth, and Elena Castro-Martínez. Are ‘STEM from Mars and SSH from Venus’?: Challenging disciplinary stereotypes of research’s social value. Science and Public Policy, 41(3):384– 400, June 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Olmos-Penuela:2016:DIT

[OPBCM16] Julia Olmos-Peñuela, Paul Benneworth, and Elena Castro-Martínez. Does it take two to tango? Factors related to the ease of societal uptake of scientific knowledge. Science and Public Policy, 43(6):751– ??, December 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/43/
                                        6/751/2525514/Does-it-take-
                                        two-to-tango-Factors-related-
                                        to-the.

Olmos-Penuela:2014:ICB

[OPMGCM14] Julia Olmos-Peñuela, Jordi Molas-Gallart, and Elena Castro-Martínez. Informal collaborations between social sciences and humanities researchers and non-academic partners. Science and Public Policy, 41(4):493–506, July 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Orozco:2010:QIB

[OR10] Jeffrey Orozco and Keynor Ruiz. Quality of interactions between public research organisations and firms: lessons from Costa Rica. Science and Public Policy, 37(7):527–540, August 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Orlic:2019:RDI

[ORB19] Edvard Orlic, Dragana Radicic, and Merima Balavac. R&D and innovation policy in the Western Balkans: Are there additionality effects? Science and Public Policy, 46(6):876–894, December 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/6/876/5543070.

Orstavik:2014:BRB

[Ors14] Finn Orstavik. Book review:

Competing through innovation: Technology Strategy and Antitrust Policies by David J. Teece. Science and Public Policy, 41(2):257–258, April 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/41/2/257.full.pdf+
html.

Page:2011:FBT

[Pag11] William Page. Farewell Beech Tree, hello Oxford University Press. Science and Public Policy, 38(10):742, December 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Parthasarathy:2010:BEB

[Par10] PhD Shobita Parthasarathy.

Breaking the expertise barrier: understanding activist strategies in science and technology policy domains. Science and Public Policy, 37(5):355– 367, June 2010. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Protogerou:2010:IEP

[PCS10] Aimilia Protogerou, Yannis Caloghirou, and Evangelos Siokas. The impact of EU policy-driven research networks on the diffusion and deployment of innovation at the national level: the case of Greece. Science and Public Policy, 37(4):283–296, May 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Protogerou:2013:RNT

[PCS13] Aimilia Protogerou, Yannis Caloghirou, and Evangelos Siokas. Research networking and technology fusion through EU-funded collaborative projects. Science and Public Policy, 40(5):576–590, October 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Parotte:2018:CPL

[PD18] Céline Parotte and Pierre Delvenne. Co-produced legitimacies: Parliamentary technology assessment and nuclear waste management in France. Science and Public Policy, 45(6):853–862, December 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/6/853/4851128.

Pelaez:2013:RPC

[PdSA13] Victor Pelaez, Letícia Rodrigues da Silva, and Eduardo Borges Araújo. Regulation of pesticides: a comparative analysis*. Science and Public Policy, 40(5):644–656, October 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pieczka:2013:DSI

[PE13] Magda Pieczka and Oliver Escobar. Dialogue and science: Innovation in policy-making and the discourse of public engagement in the UK. Science and Public Policy, 40(1):113– 126, February 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pearce:2016:RIO

[Pea16] J. M. Pearce. Return on investment for open source scientific hardware development. Science and Public Policy, 43 (2):192–195, April 2016. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pittens:2014:RAI

[PEV^{+}14] Carina A. C. M. Pittens, Janneke E. Elberse, Merel Visse, Tineke A. Abma, and Jacqueline E. W. Broerse. Research agendas involving patients: Factors that facilitate or impede translation of patients’ perspectives in programming and implementation. Science and Public Policy, 41(6):809–820, December 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Petersohn:2018:PBRa

[PH18a] Sabrina Petersohn and Thomas Heinze. Professionalization of bibliometric research assessment. insights from the history of the Leiden Centre for Science and Technology Studies (CWTS). Science and Public Policy, 45(4):565–578, August 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/4/565/4745838.

Petersohn:2018:PBRb

[PH18b] Sabrina Petersohn and Thomas Heinze. Professionalization of bibliometric research assessment. Insights from the history of the Leiden Centre for Science and Technology Studies (CWTS). Science and Public Policy, 45(4):596, August 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/4/596/4831257.

Philipps:2013:MSS

[Phi13] Axel Philipps. Mission statements and self-descriptions of German extra-university research institutes: a qualitative content analysis. Science and Public Policy, 40(5):686–697, October 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Phillips:2017:PBD

[Phi17] Fred Phillips. A perspective on ‘Big Data’. Science and Public Policy, 44 (5):730–737, October 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

//academic.oup.com/spp/
article/44/5/730/3738035.

Park:2015:ERC

[PJYL15] Joohyoung Park, Seongkyoon Jeong, Yongik Yoon, and Heesang Lee. The evolving role of collaboration in developing scientific capability: Evidence from Korean government-supported research institutes. Science and Public Policy, 42(2): 255–272, April 2015. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http://

spp.oxfordjournals.org/
content/42/2/255.

Patterson:2018:RRC

[PM18] Alan Patterson and Craig McLean. The regulation of risk: the case of fracking in the UK and The Netherlands. Science and Public Policy, 45(1):45–52, February 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/1/45/3980181.

Patterson:2019:PPW

[PM19] Alan Patterson and Craig McLean. The precautionary principle at work: The case of neonicotinoids and the health of bees. Science and Public Policy, 46(3):441–449, June 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/46/3/441/5262292.

Palmer:2019:PEP

[POD19] James Palmer, Susan Owens, and Robert Doubleday. Perfecting the ‘elevator pitch’? Expert advice as locallysituated boundary work. Science and Public Policy, 46(2):244–253, April 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/46/2/244/5106397.

Porteous:2016:MEO

[Por16] Mike Porteous. The magician’s hat: Evidence and openness in policy making. Science and Public Policy, 43 (2):157–168, April 2016. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Povoa:2010:TTU

[PR10] Luciano Martins Costa Póvoa and Márcia Siqueira Rapini. Technology transfer from universities and public research institutes to firms in Brazil: what is transferred and how the transfer is carried out. Science and Public Policy, 37(2): 147–159, March 2010. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Prem:2014:IIR

[Pre14] Erich Prem. Identifying international research cooperation capabilities in information and communication technologies. Science and Public Policy, 41(5):658–672, October 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pribilsky:2019:BRD

[Pri19] Jason Pribilsky. Book review: Drone: Remote control warfare by Hugh Gusterson. Science and Public Policy, 46(5):785–786, October 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/5/785/5485121.

Pohl:2010:RRK

[PRZ^{+}10] Christian Pohl, Stephan Rist, Anne Zimmermann, Patricia Fry, Ghana S. Gurung, Flurina Schneider, Chinwe Ifejika Speranza, Boniface Kiteme, Sébastian Boillat, Elvira Serrano, Gertrude Hirsch Hadorn, and Urs Wiesmann. Researchers’ roles in knowledge co-production: experience from sustainability research in Kenya, Switzerland, Bolivia and Nepal. Science and Public Policy, 37(4):267– 281, May 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Radosevic:2011:SIL

[Rad11] Slavo Radosevic. Scienceindustry links in Central and Eastern Europe and the Commonwealth of Independent States: conventional policy wisdom facing reality. Science and Public Policy, 38(5):365– 378, June 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Radicic:2019:EPP

[Rad19] Dragana Radicic. Effectiveness of public procurement of innovation versus supply-side innovation measures in manufacturing and service sectors. Science and Public Policy, 46(5):732–746, October 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/46/5/732/5498790.

Rahder:2015:BRB

[Rah15] Micha Rahder. Book review: Can we bridge the ‘Anthropocene Gap’? Global Environmental Governance, Technology and Politics: The Anthropocene Gap by Victor Galaz. Science and Public Policy, 42(3):431–432, June 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/42/3/431.

Ramirez:2019:SPS

[RB19] Luisa F. Ramirez and Brian M.

Belcher. Stakeholder perceptions of scientific knowledge in policy processes: a Peruvian case-study of forestry policy development. Science and Public Policy, 46(4):504–517, August 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/4/504/5306126.

Roelofsen:2010:EFT

[RBdCBB10] Anneloes Roelofsen, Jacqueline Broerse, Tjard de Cock Buning, and Joske Bunders. Engaging with future technologies: how potential future users frame ecogenomics. Science and Public Policy, 37(3): 167–179, April 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Russo:2019:IIP

[RCRR19] Margherita Russo, Annalisa Caloffi, Federica Rossi, and Riccardo Righi. Innovation intermediaries and performance-based incentives: a case study of regional innovation poles. Science and Public Policy, 46(1):1–12, February 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/46/1/1/4939323.

Renault:2016:COB

[RdMdAFY16] Thiago Renault, José Manoel Carvalho de Mello, Marcus Vinicius de Araújo Fonseca, and Sérgio Yates. A chip off the old block: Case studies of university influence on academic spin-offs. Science and Public Policy, 43(5):594– ??, October 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/43/
                                        5/594/2525537/A-chip-off-
                                        the-old-block-Case-studies-
                                        of.

Rushforth:2017:QMT

[RdR17] Alexander D. Rushforth and Sarah de Rijcke. Quality monitoring in transition: The challenge of evaluating translational research programs in academic biomedicine. Science and Public Policy, 44(4): 513–??, August 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/44/
                                        4/513/2525581/Quality-monitoring-
                                        in-transition-The-challenge-
                                        of.

Reinsborough:2015:IED

[Rei15] Michael Reinsborough. Institutional economics for a discipline at sea. Capitalism and Democracy: a Fragile Alliance by Theo van de Klundert. Science and Public Policy, 42 (1):139–141, February 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Renninger:2013:BRB

[Ren13] Bryce J. Renninger. Book review: Grace Hopper and the Information Age by Kurt W. Beyer. Science and Public Policy, 40(3):419–420, June 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rosenbloom:2017:ESS

[RG17] Joshua L. Rosenbloom and Donna K. Ginther. The effectiveness of social science research in addressing societal problems: Broadening participation in computing. Science and Public Policy, 44(2): 259–??, April 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
2/259/2584493/The-effectiveness-
of-social-science-research-
in.

Rodriguez-Gulias:2018:RKS

[RGFLRP^{+}18] María Jesús Rodríguez-Gulías, Sara Fernández-López, David Rodeiro-Pazos, Christian Corsi, and Antonio Prencipe. The role of knowledge spillovers on the university spin-offs innovation. Science and Public Policy, 45(6):875–883, December 1, 2018. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/6/875/4855832.

Robinson-Garcia:2018:UAC

[RGvLR18] Nicolas Robinson-Garcia, Thed N. van Leeuwen, and Ismael Ràfols. Using altmetrics for contextualised mapping of societal impact: From hits to networks. Science and Public Policy, 45(6):815–826, December 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/6/815/4925531.

Rekers:2015:IRG

[RH15] Josephine V. Rekers and Teis Hansen. Interdisciplinary research and geography: Overcoming barriers through proximity. Science and Public Policy, 42(2):242–254, April 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /spp.oxfordjournals.org/
                                        content/42/2/242.

Regan:2019:MSA

[RH19] Áine Regan and Maeve Henchion. Making sense of altmetrics: The perceived threats and opportunities for academic identity. Science and Public Policy, 46(4):479–489, August 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/46/4/479/5304700.

Rosenlund:2015:CNE

[RHJS15] Joacim Rosenlund, William Hogland, Anders W. Johansson, and Jackie Seddon. A cross-national environmental cluster collaboration: Shifting between an analytical and management level of the triple helix. Science and Public Policy, 42(4):583–593, August 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Reddy:2012:LRA

[RHS12] T. S. Vamsidhar Reddy, Andy Hall, and Rasheed Sulaiman. Locating research in agricultural innovation trajectories: Evidence and implications from empirical cases from South Asia. Science and Public Policy, 39(4):476–490, August 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ricard:2016:AIG

[Ric16] Lykke Margot Ricard. Aligning innovation with grand societal challenges: Inside the European Technology Platforms in wind, and carbon capture and storage. Science and Public Policy, 43(2):169– 183, April 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Revill:2014:TKB

[RJ14] James Revill and Catherine Jefferson. Tacit knowledge and the biological weapons regime. Science and Public Policy, 41(5):597–610, October 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lee:2011:PCC

[rL11] Kong rae Lee. Path-creating capability building and innovation in Korea’s leading firms. Science and Public Policy, 38(10):743–754, December 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Reiss:2014:ISS

[RM14] Thomas Reiss and Kate Millar. Introduction to special section: Assessment of emerging science and technology: Integration opportunities and challenges. Science and Public Policy, 41(3):269–271, June 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rask:2012:IPE

[RMZP12] Mikko Rask, Saule Maciukaite-Zviniene, and Jurgita Petrauskiene. Innovations in public engagement and participatory performance of the nations. Science and Public Policy, 39(6):710–721, December 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rodriguez-Navarro:2018:EPD

[RNN18] Alonso Rodríguez-Navarro and Francis Narin. European paradox or delusion — are European science and economy outdated? Science and Public Policy, 45(1):14–23, February 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/1/14/3788009.

Robbins:2014:BRW

[Rob14] Carol A. Robbins. Book review: Why innovation measurement is as hard as it looks: Handbook of Innovation Indicators and Measurement, edited by Fred Gault. Science and Public Policy, 41 (3):406–407, June 2014. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rodriguez:2010:GEC

[Rod10] Fernando Santiago Rodríguez.

Governing ethical clinical research in developing countries: exploring the case of Mexico. Science and Public Policy, 37(8):583–596, October 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rodder:2015:SMC

[Röd15] Simone Rödder. Science Media Centres and public policy. Science and Public Policy, 42(3):387–400, June 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/42/3/387.

Rolfstam:2013:PPI

[Rol13] Max Rolfstam. Public procurement and innovation: the role of institutions. Edward Elgar Publishing, Northampton, MA, USA, 2013. ISBN 1-84980-287-4 (hardcover). ???? pp. LCCN ???? URL

http://www.e-elgar.com/
shop/public-procurement-
and-innovation.

Rosales:2010:NGG

[Ros10] Dr. Janna Rosales. Nano goes global. Science and Public Policy, 37(8):641–642, October 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Radicic:2017:RPP

[RP17] Dragana Radicic and Geoffrey Pugh. R&D programmes, policy mix, and the ‘European Paradox’: Evidence from European SMEs. Science and Public Policy, 44(4):497– ??, August 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
4/497/2525580/R-amp-D-Programmes-
Policy-Mix-and-the-European.

Ranga:2016:NFU

[RPK16] Marina Ranga, Juha Perälampi, and Juha Kansikas. The new face of university–business cooperation in Finland. Science and Public Policy, 43(5):601– ??, October 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/
5/601/2770712/The-new-face-
of-university-business-
cooperation-in.

Radatz:2019:AES

[RRF^{+}19] Alecia Radatz, Michael Reinsborough, Erik Fisher, Elizabeth Corley, and David Guston. An assessment of engaged social science research in nanoscale science and engineering communities. Science and Public Policy, 46(6):853–865, December 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/6/853/5552834.

Rosenlund:2017:OTH

[RRH17] Joacim Rosenlund, Erik Rosell, and William Hogland. Overcoming the triple helix boundaries in an environmental research collaboration. Science and Public Policy, 44(2): 153–??, April 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
2/153/2525546/Overcoming-
the-triple-helix-boundaries-
in-an.

Rey-Rocha:2012:GFC

[RRMS12] Jesús Rey-Rocha and María José Martín-Sempere. Generating favourable contexts for translational research through the incorporation of basic researchers into hospitals: The FIS/Miguel Servet Research Contract Programme. Science and Public Policy, 39(6):787– 801, December 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ramos-Vielba:2018:ISS

[RVDWA18] Irene Ramos-Vielba, Pablo D’Este, Richard Woolley, and Nabil Amara. Introduction to a special section: Balancing scientific and societal impact — a challenging agenda for academic research. Science and Public Policy, 45(6):749–751, December 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/6/749/4994268.

Ryan:2010:MTI

[Rya10] Dr. Camille D. Ryan. Myths of technology: innovation and inequality. Science and Public Policy, 37(9):726–727, November 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ryan:2011:IEC

[Rya11] Dr. Jim Ryan. Irish experience of cross-sector research collaboration initiatives. Science and Public Policy, 38(2): 147–155, March 2011. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ryan:2015:GER

[Rya15] Lorna Ryan. Governance of EU research policy: Charting forms of scientific democracy in the European research area. Science and Public Policy, 42(3):300–314, June 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/42/3/300.

Schmitz:2016:IPE

[SA16] Hubert Schmitz and Tilman Altenburg. Innovation paths in Europe and Asia: Divergence or convergence? Science and Public Policy, 43(4):454– 463, August 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Saarela:2019:PSP

[Saa19] Sanna-Riikka Saarela. From pure science to participatory knowledge production? Researchers’ perceptions on science–policy interface in bioenergy policy. Science and Public Policy, 46(1):81–90, February 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/1/81/4976487.

Shibayama:2011:SRT

[SB11] Sotaro Shibayama and Yasunori Baba. Sharing research tools in academia: the case of Japan. Science and Public Policy, 38(8):649–659, October 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sanchez-Barrioluengo:2016:RHC

[SBC16] Mabel Sánchez-Barrioluengo and Davide Consoli. Regional human capital and university orientation: a case study on Spain. Science and Public Policy, 43(6):798–??, December 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/43/
                                        6/798/2525538/Regional-
                                        human-capital-and-university-
                                        orientation.

Sarpong:2018:MMR

[SBD18] David Sarpong, David Botchie, and Bidit Dey. From marginal to mainstream: The revival, transformation, and boom of plant medicine. Science and Public Policy, 45 (6):843–852, December 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/6/843/4911087.

Schneider:2019:RFP

[SBK^{+}19] Flurina Schneider, Tobias Buser, Rea Keller, Theresa Tribaldos, and Stephan Rist. Research funding programmes aiming for societal transformations: Ten key stages. Science and Public Policy, 46(3):463–478, June 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/3/463/5305067.

Sinha:2017:ERP

[SBKP17] Pawan Sinha, Peter Bex, Margaret Kjelgaard, and Flip Phillips. Enhancing research with Plenary Labs. Science and Public Policy, 44(3): 434–??, June 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
3/434/2628053/Enhancing-
research-with-Plenary-Labs.

Serger:2015:CUG

[SBL15] Sylvia Schwaag Serger, Mats Benner, and Li Liu. Chinese university governance: Tensions and reforms. Science and Public Policy, 42(6):871–886, December 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Suarez:2018:SDP

[SBY18] Diana Suárez, Florencia Barletta, and Gabriel Yoguel. Sectoral dialogue and promotion of innovation in Argentina. Science and Public Policy, 45(3):309–317, June 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/3/309/4094886.

Schmidt:2019:SDFa

[SC19a] Evanthia Kalpazidou Schmidt and Marina Cacace. Setting up a dynamic framework to activate gender equality structural transformation in research organizations. Science and Public Policy, 46(1):159, February 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/1/159/5232681.

Schmidt:2019:SDFb

[SC19b] Evanthia Kalpazidou Schmidt and Marina Cacace. Setting up a dynamic framework to activate gender equality structural transformation in research organizations. Science and Public Policy, 46(3):321– 338, June 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/3/321/5138293.

Su:2016:AGH

[SCB^{+}16] Leona Yi-Fan Su, Michael A. Cacciatore, Dominique Brossard, Elizabeth A. Corley, Dietram A. Scheufele, and Michael A. Xenos. Attitudinal gaps: How experts and lay audiences form policy attitudes toward controversial science. Science and Public Policy, 43(2):196–206, April 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Savaget:2019:EPP

[SCE19] Paulo Savaget, Tulio Chiarini, and Steve Evans. Empowering political participation through artificial intelligence. Science and Public Policy, 46(3):369– 380, June 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/3/369/5161215.

Schiller:2011:IPC

[Sch11] Daniel Schiller. Institutions and practice in cross-sector research collaboration: conceptual considerations with empirical illustrations from the German science sector. Science and Public Policy, 38(2): 109–121, March 2011. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Satterfield:2013:USP

[SCH^{+}13] Terre Satterfield, Joe Conti, Barbara Herr Harthorn, Nick Pidgeon, and Anton Pitts. Understanding shifting perceptions of nanotechnologies and their implications for policy dialogues about emerging technologies. Science and Public Policy, 40(2):247–260, April 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Schrage:2014:IHH

[Sch14] Michael Schrage. The innovator’s hypothesis: how cheap experiments are worth more than good ideas. MIT Press, Cambridge, MA, USA, 2014. ISBN 0-262-02836-0 (hardcover). xvi + 237 pp. LCCN HD45 .S367 2014.

Scharnhorst:2015:BRB

[Sch15] Andrea Scharnhorst. Book review: Beyond Bibliometrics: Harnessing Multidimensional Indicators of Scholarly Impact edited by Blaise Cronin and Cassidy R. Sugimoto. Science and Public Policy, 42(3):429–431, June 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /spp.oxfordjournals.org/
                                        content/42/3/429.

Scott:2014:UFT

[Sco14] John T. Scott. The US Federal Trade Commission’s Line of Business Program and innovation research. Science and Public Policy, 41(4):438– 448, July 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Suarez:2015:RPI

[SD15] Marcela Suárez and Gabriela Dutrénit. The role of policy incentives in the reproduction of asymmetries within nanotechnology knowledge networks. Science and Public Policy, 42(1):59–71, February 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Saad:2011:IGP

[SDP11] Professor Mohammed Saad, Dr. Surja Datta, and Dimitrios Papadakis. Innovation in global public goods: issues and challenges in provision of environment-monitoring services in Europe. Science and Public Policy, 38(4):307–318, May 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Seale:2013:BRB

[Sea13] Catherine Seale. Book review:

The New Harvest: Agricultural Innovation in Africa by Calestous Juma. Science and Public Policy, 40(6):817–818, December 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sen:2014:TRT

[Sen14] Avery Sen. Totally radical:

From transformative research to transformative innovation. Science and Public Policy, 41 (3):344–358, June 2014. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sen:2017:IBH

[Sen17] Avery Sen. Island + Bridge:

how transformative innovation is organized in the federal government. Science and Public Policy, 44(5):

707–721, October 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

//academic.oup.com/spp/
article/44/5/707/3744806.

Sevignani:2013:CPI

[Sev13] Sebastian Sevignani. The commodification of privacy on the Internet. Science and Public Policy, 40(6):733–739, December 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Schmid:2019:VPI

[SF19] Jon Schmid and Ayodeji Fajebe. Variation in patent impact by organization type: an investigation of government, university, and corporate patents. Science and Public Policy, 46(4):589–598, August 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/4/589/5382094.

Salles-Filho:2010:TOP

[SFBB10] Sergio Salles-Filho, Maria Beatriz, and Machado Bonacelli. Trends in the organization of public research organizations: lessons from the Brazilian case. Science and Public Policy, 37(3):193–204, April 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sleeboom-Faulkner:2018:RCBa

[SFCR18a] Margaret Sleeboom-Faulkner, Haidan Chen, and Achim Rosemann. Regulatory capacity building and the governance of clinical stem cell research in China. Science and Public Policy, 45(3):416–427, June 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/3/416/4675127.

Sleeboom-Faulkner:2018:RCBb

[SFCR18b] Margaret Sleeboom-Faulkner, Haidan Chen, and Achim Rosemann. Regulatory capacity building and the governance of clinical stem cell research in China. Science and Public Policy, 45(3):431, June 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/3/431/4833895.

Shen:2016:DTS

[SFZ16] Qunhong Shen, Kaidong Feng, and Xiaobin Zhang. Divergent technological strategies among leading electric vehicle firms in China: Multiplicity of institutional logics and responses of firms. Science and Public Policy, 43(4):492–504, August 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Schneider:2016:BVA

[SG16] Michael Schneider and Bernhard Gill. Biotechnology versus agroecology: Entrenchments and surprise at a 2030 forecast scenario workshop. Science and Public Policy, 43(1):74–84, February 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/43/1/74.

Salter:2014:CPG

[SH14a] Brian Salter and Alison Harvey. Creating problems in the governance of science: Bioethics and human/animal chimeras. Science and Public Policy, 41(5):685–696, October 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sorin:2014:EWE

[SH14b] Martin D. Sorin and Randall J. Hannum. Erratum to ‘Which extramural scientists were funded by the US National Institutes of Health from its ARRA funds?’. Science and Public Policy, 41(2):268, April 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/41/2/268.full.pdf+

html. See [SH14c].

Sorin:2014:WES

[SH14c] Martin D. Sorin and Randall J. Hannum. Which extramural scientists were funded by the US National Institutes of Health from its ARRA funds? Science and Public Policy, 41(1):58–75, February 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). See erratum [SH14b].

Sharma:2017:BRA

[Sha17] Aviram Sharma. Book review: American Environmental Policy: The Failures of Compliance, Abatement and Mitigation. Science and Public Policy, 44(1):150– ??, February 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
1/150/2525551/American-
Environmental-Policy-The-
Failures-of.

Sharma:2018:ADW

[Sha18] Aviram Sharma. Academic diary: Or why higher education still matters. Science and Public Policy, 45(1): 139–140, February 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/1/139/4037459.

Sheoran:2015:BRT

[She15] Nayantara Sheoran. Book review: The Therapeutic Cloning Debate: Global Science and Journalism in the Public Sphere by Eric A. Jensen. Science and Public Policy, 42(5):741–742, October 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /spp.oxfordjournals.org/
                                        content/42/5/741.

Shipp:2014:UCI

[Shi14] Stephanie S. Shipp. Book review: Urban Competitiveness and Innovation edited by Pengfei Ni and Zheng Qiongjie. Science and Public Policy, 41(5):698–699, October 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Shipp:2015:GUS

[Shi15] Stephanie S. Shipp. Governance for urban sustainability and resilience. Science and Public Policy, 42(6):904–905, December 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Suurna:2010:EIP

[SK10] Margit Suurna and Rainer Kattel. Europeanization of innovation policy in Central and Eastern Europe. Science and Public Policy, 37(9):646–664, November 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Spithoven:2011:RBC

[SK11] André Spithoven and Mirjam Knockaert. The role of business centres in firms’ networking capabilities and performance. Science and Public Policy, 38(7):569–580, August 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Schulze-Krogh:2018:FAC

[SK18] Ann Camilla Schulze-Krogh.

Firms’ absorptive capacity for research-based collaboration — an analysis of a Norwegian R&D brokering policy program. Science and Public Policy, 45(4):533–542, August 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/4/533/4587619.

Schmid:2018:PVEa

[SKY18a] Jon Schmid, Sergey A.

Kolesnikov, and Jan Youtie. Plans versus experiences in transitioning transnational education into research and economic development: a case study. Science and Public Policy, 45(1):103–116, February 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/1/103/4139726.

Schmid:2018:PVEb

[SKY18b] Jon Schmid, Sergey A.

Kolesnikov, and Jan Youtie. Plans versus experiences in transitioning transnational education into research and economic development: a case study. Science and Public Policy, 45(3):432, June 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/3/432/4930730.

Sa:2011:UIR

[SL11] Dr. Creso M. Sá and Dr. Jeffrey Litwin. Universityindustry research collaborations in Canada: the role of federal policy instruments. Science and Public Policy, 38 (6):425–435, July 2011. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Schuelke-Leech:2014:BRB

[SL14] Beth-Anne Schuelke-Leech. Book review: Cycles, Crises and Innovation: Path to Sustainable Development — A Kaleckian–Schumpeterian Synthesis by Jerry Courvisanos. Science and Public Policy, 41 (1):131–132, February 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Schuelke-Leech:2017:STE

[SL17] Beth-Anne Schuelke-Leech. Smart technologies and the end(s) of law. Science and Public Policy, 44(4):579– ??, August 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/44/
                                        4/579/3806863/Smart-Technologies-
                                        and-the-End-s-of-Law.

Schut:2010:RRR

[SLvP10] Marc Schut, Cees Leeuwis, and Annemarie van Paassen. Room for the river: Room for research? The case of depoldering De Noordwaard, The Netherlands. Science and Public Policy, 37(8):611–627, October 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sanders:2010:RNB

[SM10] Carrie B. Sanders and Fiona Alice Miller. Reframing norms: boundary maintenance and partial accommodations in the work of academic technology transfer. Science and Public Policy, 37(9):689–701, November 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Schlierf:2013:SKI

[SM13] Katharina Schlierf and Morgan Meyer. Situating knowledge intermediation: Insights from science shops and knowledge brokers. Science and Public Policy, 40(4):430–441, August 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Smith:2014:PIC

[Smi14] Katherine E. Smith. The politics of ideas: The complex interplay of health inequalities research and policy. Science and Public Policy, 41(5):561– 574, October 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Smirnova:2016:UIK

[Smi16] Yelena V. Smirnova. University– industry knowledge transfer in an emerging economy: Evidence from Kazakhstan. Science and Public Policy, 43(5):702–??, October 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/
5/702/2725286/University-
industry-knowledge-transfer-
in-an.

Sanchez:2016:BFS

[SMR16] Antonio García Sánchez, José Molero, and Ruth Rama. Are ‘the best’ foreign subsidiaries cooperating for innovation with local partners? The case of an intermediate country. Science and Public Policy, 43(4):532–545, August 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sanz-Menendez:2014:CSG

[SMVdP14] Luis Sanz-Menéndez, Gregg G. Van Ryzin, and Eloisa del Pino. Citizens’ support for government spending on science and technology. Science and Public Policy, 41(5):611– 624, October 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Strauss:2013:SNS

[SN13] Stefan Strauß and Michael Nentwich. Social network sites, privacy and the blurring boundary between public and private spaces. Science and Public Policy, 40(6):724–732, December 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sugimoto:2015:RBG

[SNL15] Cassidy R. Sugimoto, Chaoqun Ni, and Vincent Larivière. On the relationship between gender disparities in scholarly communication and countrylevel development indicators. Science and Public Policy, 42 (6):789–810, December 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sarkki:2014:BCR

[SNT^{+}14] Simo Sarkki, Jari Niemeldá, Rob Tinch, Sybille van den Hove, Allan Watt, and Juliette Young. Balancing credibility, relevance and legitimacy: a critical assessment of trade-offs in science–policy interfaces. Science and Public Policy, 41(2):194–206, April 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/41/2/194.full.pdf+
html.

Soete:2010:CFP

[Soe10] Luc Soete. Chris Freeman: the person. Science and Public Policy, 37(8):629–630, October 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sonmez:2014:ACE

[Sön14a]  Zafer Sönmez.  Book re-

view: Agglomeration, Clusters and Entrepreneurship: Studies in Regional Economic Development edited by Charlie Karlsson, Börje Johansson and Roger R. Stough. Science and Public Policy, 41(6):836– 838, December 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sonmez:2014:EIP

[Son14b] Zafer Sonmez. EU innovation policy: One size doesn’t fit all. Challenges for European Innovation Policy: Cohesion and Excellence from a Schumpeterian Perspective edited by Slavo Radosevic and Anna Kaderabkova. Science and Public Policy, 41(6):835– 836, December 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sonmez:2017:IMG

[Son17]  Zafer Sonmez.  Inventor

mobility and the geography of knowledge flows: evidence from the US biopharmaceutical industry. Science and Public Policy, 44 (5):670–682, October 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

//academic.oup.com/spp/
article/44/5/670/2964537.

Sonmez:2018:BRL

[Son18] Zafer Sonmez. Book review:

The Life Cycle of Clusters: a Policy Perspective, Edited by Dirk Fornahl. Science and Public Policy, 45(4):592–594, August 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/4/592/4911558.

Stampfer:2010:MRF

[SPH10] Michael Stampfer, Rupert Pichler, and Reinhold Hofer. The making of research funding in Austria: transition politics and institutional development, 1945–2005. Science and Public Policy, 37(10):765–780, December 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Salvador:2011:ISP

[SR11] Elisa Salvador and Secondo Rolfo. Are incubators and science parks effective for research spin-offs? Evidence from Italy. Science and Public Policy, 38(3):170–184, April 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sa:2017:ICK

[SRC17] Nelson Sá, Ana Paula Ribeiro, and Vítor Carvalho. International collaboration and knowledge creation: Evidence from economics in Portuguese academia. Science and Public Policy, 44(1):50– ??, February 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/44/
                                        1/50/2525509/International-
                                        collaboration-and-knowledge-
                                        creation.

Srinivas:2014:BRB

[Sri14] Krishna Ravi Srinivas. Book review: Handbook of Innovation Indicators and Measurement, edited by Fred Gault. Science and Public Policy, 41 (3):405, June 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Salter:2010:GIB

[SS10] Brian Salter and Charlotte Salter. Governing innovation in the biomedicine knowledge economy: stem cell science in the USA. Science and Public Policy, 37(2):87–100, March 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sorensen:2017:SNR

[SS17] Mads P. Sørensen and Jesper Wiborg Schneider. Studies of national research performance: a case of ‘methodological nationalism’ and ‘zombie science’? Science and Public Policy, 44(1):132–??, February 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
1/132/2979767/Studies-of-
national-research-performance-
A-case-of.

Scholvinck:2019:EMP

[SSB19] Anne-Floor M. Schölvinck, Tjerk Jan Schuitmaker, and Jacqueline E. W. Broerse. Embedding meaningful patient involvement in the process of proposal appraisal at the Dutch Cancer Society. Science and Public Policy, 46(2):254–263, April 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/2/254/5265274.

Silva:2017:DGD

[SSC17] Ana M. Silva, Sandra T. Silva, and Anabela Carneiro. Determinants of grant decisions in R&D subsidy programmes: Evidence from firms and S&T organisations in Portugal. Science and Public Policy, 44(5):683–697, October 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

//academic.oup.com/spp/
article/44/5/683/2996780.

Sigurdson:2015:LUS

[SSK15] Kristjan Sigurdson, Creso M. Sá, and Andrew Kretz. Looking under the street light: Limitations of mainstream technology transfer indicators. Science and Public Policy, 42(5):632–645, October 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /spp.oxfordjournals.org/
                                        content/42/5/632.

Simoes:2012:DFC

[SSTM12] Jorge Simões, Maria Jose Silva, Virginia Trigo, and Jacinta Moreira. The dynamics of firm creation fuelled by higher education institutions within innovation networks. Science and Public Policy, 39(5):630– 640, October 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stahl:2013:RRI

[Sta13] Bernd Carsten Stahl. Responsible research and innovation: The role of privacy in an emerging framework. Science and Public Policy, 40(6):708– 716, December 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stegenga:2010:FOU

[Ste10]  Jacob Stegenga.  A fun

overview for undergraduates. Science and Public Policy, 37(1):79–80, February 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stahl:2017:EEI

[STF17] Bernd Carsten Stahl, Job Timmermans, and Catherine Flick. Ethics of emerging information and communication technologies: On the implementation of responsible research and innovation. Science and Public Policy, 44(3): 369–??, June 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
3/369/2525576/Ethics-of-
Emerging-Information-and-
Communication.

Stine:2017:MET

[Sti17] Deborah D. Stine. Book review: Managing Emerging Technologies for Socio-Economic Impact, edited by Dimitris G. Assimakopoulos, Ilan Oshri and Krsto Pandza: Edward Elgar, Cheltenham, UK and Northampton, MA, 2015, 416 pages, £85.50, US$150.00 (hardback), ISBN 978-1-78254-787-7. Science and Public Policy, 44(2):298– ??, April 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
2/298/2525552/Managing-
Emerging-Technologies-for-
Socio-Economic.

Su:2014:BRB

[Su14] Xuhong Su. Book review: Is American Science in Decline? by Yu Xie and Alexandra A. Killewald. Science and Public Policy, 41(2):265–266, April 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /spp.oxfordjournals.org/
                                        content/41/2/265.full.pdf+
                                        html.

Su:2016:BRE

[Su16]  Xuhong Su.  Book re-

view: Environmental Taxation and Green Fiscal Reforms: Theory and Impact, edited by Larry Kreiser, Soocheol Lee, Kazuhiro Ueta, Janet E. Milne and Hope Ashiabor. Science and Public Policy, 43(5):727–??, October 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/43/
                                        5/727/2525479/Environmental-
                                        Taxation-and-Green-Fiscal-
                                        Reforms.

Sutherlin:2017:EEJ

[Sut17] John W. Sutherlin. Ethics, environmental justice and climate change. Science and Public Policy, 44(4):578–??, August 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/44/

4/578/2624552/Ethics-Environmental-
Justice-and-Climate-Change.

Scholten:2015:PEC

[SV15] Peter Scholten and Stijn Verbeek. Politicization and expertise: Changing researchpolicy dialogues on migrant integration in Europe. Science and Public Policy, 42(2): 188–200, April 2015. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http://

spp.oxfordjournals.org/
content/42/2/188.

Schut:2014:TDR

[SvPLK14] Marc Schut, Annemarie van Paassen, Cees Leeuwis, and Laurens Klerkx. Towards dynamic research configurations: a framework for reflection on the contribution of research to policy and innovation processes. Science and Public Policy, 41(2):207–218, April 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/41/2/207.full.pdf+
html.

Stange:2015:SLK

[SvTvL15] Kari Stange, Jan van Tatenhove, and Judith van Leeuwen. Stakeholder-led knowledge production: Development of a long-term management plan for North Sea Nephrops fisheries. Science and Public Policy, 42(4):501–513, August 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Saxon:2019:DSW

[SW19] Tyler Saxon and Stephan Weiler. Defence spending and women in research: a crosscountry comparison. Science and Public Policy, 46 (5):679–688, October 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/46/5/679/5485747.

Swayamprakash:2017:RCS

[Swa17] Ramya Swayamprakash. The rise of the city: Spatial dynamics in the urban century (new horizons in regional science series). Science and Public Policy, 44(3): 429–??, June 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/44/
                                        3/429/2525562/The-Rise-
                                        of-the-City-Spatial-Dynamics-
                                        in-the-Urban.

Sharif:2019:RGC

[SX19] Naubahar Sharif and Jack Linzhou Xing. Restricted generalizability of city innovation policies: The case of e-hailing in China. Science and Public Policy, 46(6):805–819, December 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/

article/46/6/805/5564003.

See erratum [SX20].

Sharif:2020:ERG

[SX20] Naubahar Sharif and Jack Linzhou Xing. Erratum: Restricted generalizability of city innovation policies: The case of e-hailing in China. Science and Public Policy, 47(1):148, February 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/47/1/148/5592229.

See [SX19].

Saad:2011:ISI

[SZ11] Professor Mohammed Saad and Dr. Girma Zawdie. Introduction to special issue: The emerging role of universities in socio-economic development through knowledge networking. Science and Public Policy, 38(1):3–6, February 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tamura:2013:GDS

[Tam13] Suguru Tamura. Generic definition of standardization and the correlation between innovation and standardization in corporate intellectual property activities. Science and Public Policy, 40(2):143–156, April 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tamtik:2017:PCC

[Tam17] Merli Tamtik. Policy coordination challenges in governments’ innovation policy — the case of Ontario, Canada. Science and Public Policy, 44(3):417–??, June 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/44/
                                        3/417/2525578/Policy-coordination-
                                        challenges-in-governments.

Tang:2010:DIE

[Tan10] Li Tang. Deciphering innovation and entrepreneurship in functional regions. Science and Public Policy, 37(8):640– 641, October 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tang:2013:BRB

[Tan13] Li Tang. Book review: Beyond the hype: Exploring interfaces of the knowledge economy. Foundations of the Knowledge Economy: Innovation, Learning and Clusters edited by Knut Ingar Westeren. Science and Public Policy, 40(6):814– 815, December 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tangney:2017:UCC

[Tan17] Peter Tangney. The UK’s 2012 Climate Change Risk Assessment: How the rational assessment of science develops policy-based evidence. Science and Public Policy, 44(2): 225–??, April 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
2/225/2525558/The-UK-s-
2012-Climate-Change-Risk-
Assessment-How.

Tang:2018:REB

[Tan18]  Li Tang.  Return en-

trepreneurs: bind and impact. Science and Public Policy, 45(2):287–289, April 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/2/287/4265543.

Tassey:2013:BBC

[Tas13] Gregory Tassey. Beyond the business cycle: The need for a technology-based growth strategy. Science and Public Policy, 40(3):293–315, June 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tassey:2014:IIP

[Tas14]  Gregory Tassey.  Innova-

tion in innovation policy management: The Experimental Technology Incentives Program and the policy experiment. Science and Public Policy, 41(4):419–424, July 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tigabu:2015:FEA

[TBvB15] Aschalew D. Tigabu, Frans Berkhout, and Pieter van Beukering. Functional evolution and accumulation of technological innovation systems: The case of renewable energy in East Africa. Science and Public Policy, 42 (5):614–631, October 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /spp.oxfordjournals.org/
                                        content/42/5/614.

Talamini:2012:MEL

[TD12] E. Talamini and H. Dewes. The macro-environment for liquid biofuels in Brazilian science and public policies. Science and Public Policy, 39(1): 13–29, February 2012. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Torres:2011:WFD

[TDBS11] Arturo Torres, Gabriela Dutrénit, Noé Becerra, and José Luis Sampedro. What are the factors driving universityindustry linkages in latecomer firms: evidence from Mexico. Science and Public Policy, 38(1):31–42, February 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Teirlinck:2013:CPC

[TDPV13] Peter Teirlinck, Henri Delanghe, Pierre Padilla, and Arnold Verbeek. Closing the policy cycle: Increasing the utilization of evaluation findings in research, technological development and innovation policy design. Science and Public Policy, 40(3):366– 377, June 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Teil:2011:SSW

[Tei11] Geneviève Teil. The sociology of scientific work. Science and Public Policy, 38(5):419– 420, June 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Teil:2012:BRB

[Tei12] G. Teil. Book review: Art Entrepreneurship edited by M. Scherdin and I. Zander. Science and Public Policy, 39(3): 409–410, June 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Teil:2014:BRB

[Tei14] Genevieve Teil. Book reviews: Societal Entrepreneurship: Positioning, Penetrating, Promoting edited by Karin Berglund, Bengt Johannisson and Birgitta Schwartz. Social Entrepreneurship: To Act as if and Make a Difference by Björn Bjerke and Mathias Karlsson. Science and Public Policy, 41(1):139– 140, February 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Thomas:2012:STP

[TFB12] Hernán Thomas, Mariano Fressoli, and Lucas Becerra. Science and technology policy and social ex/inclusion: Analyzing opportunities and constraints in Brazil and Argentina. Science and Public Policy, 39(5):579–591, October 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Turpin:2011:ISI

[TFE11] Professor Tim Turpin and Dr. Manuel Fernández-Esquinas. Introduction to special issue: The policy rationale for cross-sector research collaboration and contemporary consequences. Science and Public Policy, 38(2):82–86, March 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Thune:2011:IUI

[TG11] Taran Thune and Magnus Gulbrandsen. Institutionalization of university-industry interaction: an empirical study of the impact of formal structures on collaboration patterns. Science and Public Policy, 38(2):99–107, March 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Thompson:2013:GSEb

[Tho13] Charis Thompson. Good science: the ethical choreography of stem cells research. Inside technology. MIT Press, Cambridge, MA, USA, 2013. ISBN 0-262-02699-6 (paperback), 0-262-31903-9 (e-book), 0-262-31904-7 (e-book). x + 343 pp. LCCN QH588.S83.

Tholen:2016:VIC

[Tho16] Berry Tholen. The value of the issue context approach for scientific policy advice. Science and Public Policy, 43(2): 184–191, April 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tiffin:2011:MRU

[TK11] Dr. Scott Tiffin and Martin Kunc. Measuring the roles universities play in regional innovation systems: a comparative study between Chilean and Canadian natural resource-based regions. Science and Public Policy, 38(1): 55–66, February 2011. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tonurist:2017:CRD

[TK17] Piret Tõnurist and Rainer Kattel. Can research, development, and innovation policies cross borders? The case of Nordic–Baltic region. Science and Public Policy, 44(3): 328–??, June 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/

                                        3/328/2525571/Can-Research-
                                        Development-and-Innovation-
                                        Policies.

teKulve:2018:IDI

[tKBKS18] Haico te Kulve, Wouter Boon, Kornelia Konrad, and Tjerk Jan Schuitmaker. Influencing the direction of innovation processes: the shadow of authorities in demand articulation. Science and Public Policy, 45(4):455–467, August 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/4/455/4915397.

Tijssen:2018:REA

[TKM18] Robert Tijssen and Erika Kraemer-Mbula. Research excellence in Africa: Policies, perceptions, and performance. Science and Public Policy, 45(3):392–403, June 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/3/392/4600842.

Trench:2012:PPS

[TM12] Brian Trench and Steven Miller. Policies and practices in supporting scientists’ public communication through training. Science and Public Policy, 39(6):722–731, December 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tonismann:2019:PBS

[Tõn19] Teele Tõnismann. Paths of Baltic States public research funding 1989–2010: Between institutional heritage and internationalisation. Science and Public Policy, 46(3):391– 403, June 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/3/391/5203405.

Trapani:2014:BRB

[Tra14] Josh Trapani. Book review:

Transnational Culture in the Internet Age edited by Sean A. Pager and Adam Candeub. Science and Public Policy, 41 (1):134–135, February 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Trapani:2018:KSN

[Tra18] Josh Trapani. Knowledge for sale: The neoliberal takeover of higher education. Science and Public Policy, 45(2):284–285, April 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/2/284/4085884.

Thune:2016:UEE

[TRGA16] Taran Thune, Ingvild Reymert, Magnus Gulbrandsen, and Per Olaf Aamodt. Universities and external engagement activities: Particular profiles for particular universities? Science and Public Policy, 43(6):774–??, December 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/43/
6/774/2525521/Universities-
and-external-engagement-
activities.

Toivanen:2015:EIE

[TS15] Hannes Toivanen and Arho Suominen. Epistemic integration of the European Research Area: The shifting geography of the knowledge base of Finnish research, 1995– 2010. Science and Public Policy, 42(4):549–566, August 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tsipouri:2012:CIP

[Tsi12] Lena Tsipouri. Comparing innovation performance and science in society in the European member states. Science and Public Policy, 39(6):732– 740, December 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Turnhout:2013:NRS

[TSK^{+}13] Esther Turnhout, Marian Stuiver, Judith Klostermann, Bette Harms, and Cees Leeuwis. New roles of science in society: Different repertoires of knowledge brokering. Science and Public Policy, 40(3):354–365, June 2013.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tupasela:2011:GWC

[Tup11] Aaro Tupasela. From gift to waste: changing policies in biobanking practices. Science and Public Policy, 38(7):510– 520, August 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tuunainen:2011:HTH

[Tuu11] Juha Tuunainen. High-tech hopes: policy objectives and business reality in the biopharmaceutical industry. Science and Public Policy, 38(5): 338–348, June 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Turpin:2011:CSR

[TWGJ11] Tim Turpin, Richard Woolley, and Sam Garrett-Jones. Cross-sector research collaboration in Australia: the Cooperative Research Centres Program at the crossroads. Science and Public Policy, 38(2): 87–97, March 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Talwar:2011:UES

[TWR11] Sonia Talwar, Arnim Wiek, and John Robinson. User engagement in sustainability research. Science and Public Policy, 38(5):379–390, June 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Trencher:2014:BTM

[TYM^{+}14] Gregory Trencher, Masaru Yarime, Kes B. McCormick, Christopher N. H. Doll, and Steven B. Kraines. Beyond the third mission: Exploring the emerging university function of co-creation for sustainability. Science and Public Policy, 41(2):151–179, April 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/41/2/151.full.pdf+
html.

Thurner:2018:TTR

[TZ18] Thomas Wolfgang Thurner and Stanislav Zaichenko. Technology transfer into Russia’s agricultural sector — can public funding replace ailing business engagement? Science and Public Policy, 45(5):683–691, October 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/5/683/4836308.

Timmermans:2013:CUW

[TZI13] Bram Timmermans and Jon Mikel Zabala-Iturriagagoitia. Coordinated unbundling: a way to stimulate entrepreneurship through public procurement for innovation. Science and Public Policy, 40(5):674–685, October 2013. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Uriona:2019:ISP

[UG19] Mauricio Uriona and Sara S. (Saartjie) Grobbelaar. Innovation system policy analysis through system dynamics modelling: a systematic review. Science and Public Policy, 46(1):28–44, February 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/1/28/4962614.

Uctu:2014:BEB

[UJ14] Ramazan Uctu and Rachel C. C. Jafta. Bio-entrepreneurship as a bridge between science and business in a regional cluster: South Africa’s first attempts. Science and Public Policy, 41(2):219–233, April 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/41/2/219.full.pdf+
html.

Ulnicane:2015:WDI

[Uln15] Inga Ulnicane. Why do international research collaborations last? Virtuous circle of feedback loops, continuity and renewal. Science and Public Policy, 42(4):433–447, August 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ulnicane:2017:TSE

[Uln17]  Inga Ulnicane.  Towards

strengthening European scientific elite? Science and Public Policy, 44(4):580–??, August 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/44/
                                        4/580/2972858/Towards-strengthening-
                                        European-scientific-elite.

Valkenburg:2014:BRB

[Val14a] Govert Valkenburg. Book review: Crisis, Innovation and Sustainable Development: The Ecological Opportunity edited by Blandine Laperche, Nadine Levratto and Dimitri Uzunidis. Science and Public Policy, 41(4):548–549, July 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Vallverdu:2014:BRB

[Val14b] Jordi Vallverdú. Book review: Governance, Regulation and Innovation: Theory and Evidence from Firms and Nations edited by Mehmet Ugur. Science and Public Policy, 41(5): 697–698, October 2014. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Vanclay:2010:HEN

[Van10] Jerome K. Vanclay. Hallmarks of an effective nongovernmental organisation: the formation and management of Australia’s Wentworth Group. Science and Public Policy, 37(9):719–722, November 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Victorio:2016:MSC

[VCdS16] Carlos José Miranda Victório, Helder Gomes Costa, and Cristina Gomes de Souza. Modeling selection criteria of R&D projects for awarding direct subsidies to the private sector. Science and Public Policy, 43(2):275–287, April 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Valente:2015:MVS

[VCLA15] Adriana Valente, Tommaso Castellani, Maja Larsen, and Arja R. Aro. Models and visions of science-policy interaction: Remarks from a Delphi study in Italy. Science and Public Policy, 42(2):228–241, April 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/42/2/228.

vandeBurgwal:2019:AID

[vdBHC19] Linda H. M. van de Burgwal, Rana Hendrikse, and Eric Claassen. Aiming for impact: Differential effect of motivational drivers on effort and performance in knowledge valorisation. Science and Public Policy, 46(5):747–762, October 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/46/5/747/5506085.

vanDoren:2014:ARD

[vDFL14] Davy van Doren, Ellen-Marie Forsberg, and Ralf Lindner. Are assessments responding to a dynamic environment? Evidence from four emerging techno-scientific domains. Science and Public Policy, 41(3): 317–331, June 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

vanDoren:2014:SBT

[vDH14] Davy van Doren and Nils B. Heyen. Synthetic biology: Too early for assessments? A review of synthetic biology assessments in Germany. Science and Public Policy, 41(3): 272–282, June 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

vandeKlundert:2013:CDF

[vdK13] Theo van de Klundert. Capitalism and democracy: a fragile alliance. New horizons in institutional and evolutionary economics. Edward Elgar, Cheltenham, UK, 2013. ISBN 1-78195-615-4 (hardcover), 1-78195-616-2 (e-book). xxv + 209 pp. LCCN HB501 .K544 2013.

vanDongen:2019:PPS

[vDTC19] Peter van Dongen, Hester Tak, and Eric Claassen. Policies and patenting to stimulate the biotechnology sector: Evidence from The Netherlands. Science and Public Policy, 46(1):136–147, February 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/1/136/5049085.

vanderWeijden:2012:BBS

[vdWVvdB12] Inge van der Weijden, Maaike Verbree, and Peter van den Besselaar. From bench to bedside: The societal orientation of research leaders: The case of biomedical and health research in The Netherlands. Science and Public Policy, 39 (3):285–303, June 2012. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Vessuri:2012:ISS

[Ves12] Hebe Vessuri. Introduction to special section: The use of knowledge for social cohesion and social inclusion. Science and Public Policy, 39(5):545– 547, October 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Vignola-Gagne:2014:APS

[VG14] Etienne Vignola-Gagné. Argumentative practices in science, technology and innovation policy: The case of clinician-scientists and translational research. Science and Public Policy, 41(1):94–106, February 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Viales-Hurtado:2012:PAP

[VHAHGC12] Ronny J. Viales-Hurtado, Antonio Arellano-Hernández, and Rafael E. Granados-Carvajal. Perceptions about the political-scientific community and its role in formulating the problems of public policy for science, technology and innovation in Costa Rica. Science and Public Policy, 39(5):613–617, October 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

vonHippel:2017:MFD

[vHDdJ17] Eric von Hippel, Harold De-Monaco, and Jeroen P. J. de Jong. Market failure in the diffusion of cliniciandeveloped innovations: The case of off-label drug discoveries. Science and Public Policy, 44(1):121–??, February 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
1/121/2525545/Market-failure-
in-the-diffusion-of-clinician.

vanHolm:2019:CCN

[vHWW19] Eric Joseph van Holm, Yonghong Wu, and Eric W. Welch. Comparing the collaboration networks and productivity of China-born and US-born academic scientists. Science and Public Policy, 46(2):310–320, April 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/2/310/5137124.

Vico:2014:DSD

[Vic14] Eugenia Perez Vico. An indepth study of direct and indirect impacts from the research of a physics professor. Science and Public Policy, 41(6):701– 719, December 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Villasana:2011:FUI

[Vil11] Marcia Villasana. Fostering university-industry interactions under a Triple Helix model: the case of Nuevo Leon, Mexico. Science and Public Policy, 38(1):43–53, February 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Vitola:2015:IPM

[Vīt15]  Anete Vītola.  Innovation

policy mix in a multi-level context: The case of the Baltic Sea Region countries. Science and Public Policy, 42(3):401–414, June 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /spp.oxfordjournals.org/
                                        content/42/3/401.

Vico:2012:IEI

[VJ12] Eugenia Perez Vico and Staffan Jacobsson. Identifying, explaining and improving the effects of academic R&D: The case of nanotechnology in Sweden. Science and Public Policy, 39(4):513–529, August 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Valve:2010:DRC

[VMP10] Helena Valve, Ruth McNally, and Ari Pappinen. Doing research, creating impact: using ‘PROTEE’ to learn from a genetically modified tree field trial. Science and Public Policy, 37(5):369–379, June 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

vanRij:2010:JHS

[vR10] Victor van Rij. Joint horizon scanning: identifying common strategic choices and questions for knowledge. Science and Public Policy, 37(1):7–18, February 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Vick:2018:SLR

[VR18] Thais Elaine Vick and Maxine Robertson. A systematic literature review of UK university–industry collaboration for knowledge transfer: a future research agenda. Science and Public Policy, 45(4):579–590, August 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/4/579/4781592.

Venditti:2013:DUR

[VRL13] Michelina Venditti, Emanuela Reale, and Loet Leydesdorff. Disclosure of university research to third parties: a nonmarket perspective on an Italian university. Science and Public Policy, 40(6):792–800, December 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Waaijer:2017:PCP

[Waa17] Cathelijn J. F. Waaijer. Perceived career prospects and their influence on the sector of employment of recent PhD graduates. Science and Public Policy, 44(1):1–??, February 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article-abstract/44/1/1/
2525487/Perceived-career-
prospects-and-their-influence-
on.

Wade:2017:CGE

[Wad17] Imogen Wade. Concise guide to entrepreneurship, technology and innovation. Science and Public Policy, 44 (6):876–877, December 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

                                        //academic.oup.com/spp/
                                        article/44/6/876/3861500.

Wagner:2015:SGE

[Wag15] Aleksandra Wagner. Shale gas: Energy innovation in a (non-)knowledge society: a press discourse analysis. Science and Public Policy, 42(2):273–286, April 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /spp.oxfordjournals.org/
                                        content/42/2/273.

Walsh:2015:IFB

[Wal15] James P. Walsh. The impact of foreign-born scientists and engineers on American nanoscience research. Science and Public Policy, 42(1):107– 120, February 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wehrens:2011:CRP

[WBB11] Rik Wehrens, Marleen Bekker, and Roland Bal. Coordination of research, policy and practice: a case study of collaboration in the field of public health. Science and Public Policy, 38(10):755–766, December 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wadhwa:2015:SAS

[WBWW15] Kush Wadhwa, David Barnard-Wills, and David Wright. The state of the art in societal impact assessment for security research. Science and Public Policy, 42(3):339–354, June 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/42/3/339.

Warren:2014:PAT

[WCK^{+}14] David C. Warren, Sanya R.

Carley, Rachel M. Krause, John A. Rupp, and John D. Graham. Predictors of attitudes toward carbon capture and storage using data on world views and CCS-specific attitudes. Science and Public Policy, 41(6):821–834, December 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wardenaar:2014:VRC

[WdJH14] Tjerk Wardenaar, Stefan P. L.

de Jong, and Laurens K. Hessels. Varieties of research coordination: a comparative analysis of two strategic research consortia. Science and Public Policy, 41(6):780–792, December 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wesseling:2018:PPI

[WE18] Joeri H. Wesseling and Charles Edquist. Public procurement for innovation to help meet societal challenges: a review and case study. Science and Public Policy, 45(4):493–502, August 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/4/493/4924894.

Webb:2011:MCC

[Web11] Janette Webb. Making climate change governable: the case of the UK climate change risk assessment and adaptation planning. Science and Public Policy, 38(4):279–292, May 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Westfall:2012:IPM

[Wes12] Catherine Westfall. Institutional persistence and the material transformation of the US national labs: The curious story of the advent of the Advanced Photon Source. Science and Public Policy, 39(4): 439–449, August 2012. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wright:2013:IPE

[WF13] David Wright and Michael Friedewald. Integrating privacy and ethical impact assessments. Science and Public Policy, 40(6):755–766, December 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wieczorek:2012:CSI

[WH12a] Anna J. Wieczorek and Marko P. Hekkert. Corrigendum to ‘systemic instruments for systemic innovation problems: a framework for policy makers and innovation scholars’. Science and Public Policy, 39(6):842, December 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). See [WH12b].

Wieczorek:2012:SIS

[WH12b] Anna J. Wieczorek and Marko P. Hekkert. Systemic instruments for systemic innovation problems: a framework for policy makers and innovation scholars. Science and Public Policy, 39(1):74– 87, February 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). See corrigendum [WH12a].

Whalen:2019:RFE

[Wha19] Ryan Whalen. Research funding’s ‘endorsement effect’ on scientific boundary work and research production: Government legitimization of alternative medicine. Science and Public Policy, 46(1):71–80, February 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/1/71/4984520.

Weber:2012:CFC

[WHKB12] K. Matthias Weber, Jennifer Cassingena Harper, Totti Könnölä, and Vicente Carabias Barceló. Coping with a fast-changing world: Towards new systems of future-oriented technology analysis. Science and Public Policy, 39(2):153– 165, March 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wang:2019:CPM

[WHLhC19] Jue Wang, Rosalie Hooi, Andrew X. Li, and Meng hsuan Chou. Collaboration patterns of mobile academics: The impact of international mobility. Science and Public Policy, 46(3):450–462, June 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/46/3/450/5281269.

Wield:2013:TFC

[WHMS13] David Wield, Rebecca Hanlin, James Mittra, and James Smith. Twenty-first century bioeconomy: Global challenges of biological knowledge for health and agriculture. Science and Public Policy, 40(1):17–24, February 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wickson:2013:BRB

[Wic13a] Fern Wickson. Book review: New modes of thought required for the Age of Ecology Plato’s Revenge: Politics in the Age of Ecology by William Ophuls. Science and Public Policy, 40 (6):826–828, December 2013.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wickson:2013:RSN

[Wic13b]  Fern Wickson.  Research-

ing standards: No recipe, no risk. Science and Public Policy, 40(5):700–701, October 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Williams:2010:MMM

[Wil10] Dr. Amanda Williams. Metaphors as mirrors, magicians and mutineers. Science and Public Policy, 37(9):723–724, November 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Williams:2016:MRI

[Wil16] Amanda Williams. MOOCs:

Raising important questions about the future of higher education and the impact of new models for learning [review of MOOCs, Jonathan Haber, MIT Press, Cambridge, MA, 2014, 216 pages, US$15.95 (paperback), ISBN: 978-0-262-52691-3]. Science and Public Policy, 43 (1):143–144, February 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/43/1/143.

Wixted:2011:CTF

[Wix11] Dr. Brian Wixted. Cognition and the theory of the firm. Science and Public Policy, 38(8):661–662, October 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wixted:2015:LCD

[Wix15] Brian Wixted. Logistics clusters: Delivering value and driving growth. Science and Public Policy, 42(6):901–902, December 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Williams-Jones:2014:GDU

[WJOS14] Bryn Williams-Jones, Catherine Olivier, and Elise Smith. Governing ‘dual-use’ research in Canada: a policy review. Science and Public Policy, 41(1):76–93, February 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wigren-Kristoferson:2011:MGB

[WKGK11] Caroline Wigren-Kristoferson, Jonas Gabrielsson, and Fumi Kitagawa. Mind the gap and bridge the gap: research excellence and diffusion of academic knowledge in Sweden. Science and Public Policy, 38 (6):481–492, July 2011. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Winkler:2019:IUB

[WLA19] Anne E. Winkler, Sharon G.

Levin, and Michael T. Allison. Investigating the US biomedical workforce: Gender, field of training, and retention. Science and Public Policy, 46(6):913–926, December 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/6/913/5549136.

Wei:2012:RWT

[WLS12] Thomas E. Wei, Victoria Levin, and Lindsay M. Sabik. A referral is worth a thousand ads: Job search methods and scientist outcomes in the market for postdoctoral scholars. Science and Public Policy, 39(1):60–73, February 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wang:2015:TLC

[WLYCL15] Yuandi Wang, Jason Li-Ying, Jin Chen, and Zheng Lu. Technology licensing in China. Science and Public Policy, 42(3):293–299, June 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/42/3/293.

Williams:2018:SSM

[WM18] Trevor Williams and Juan J.

Morrone. Science is strengthened by Mexico’s researcher evaluation system: Factual errors and misleading claims by Neff. Science and Public Policy, 45(5):742–745, October 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/5/742/4816159.

See response [Nef19].

Witjes:2017:FTS

[WO17] Nina Witjes and Philipp Olbrich. A fragile transparency: satellite imagery analysis, non-state actors, and visual representations of security. Science and Public Policy, 44(4):524–??, August 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
4/524/3778300/A-fragile-
transparency-satellite-
imagery-analysis.

Wolfe:2014:IUE

[Wol14] David A. Wolfe, editor. Innovating in urban economies: economic transformation in Canadian city-regions. Innovation, creativity, and governance in Canadian cityregions. University of Toronto Press, Toronto, ON, Canada, 2014. ISBN 1-4426-4698-5 (hardcover), 1-4426-1476-5 (paper). xx + 372 pp. LCCN HC113 .I55 2014.

Woodson:2016:BRA

[Woo16] Thomas S. Woodson. Book review: Africa: Why Economists Get it Wrong by Morten Jerven. Science and Public Policy, 43(4):578–579, August 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wong:2016:APS

[WS16] Chan-Yuan Wong and Marsiti Md Salmin. Attaining a productive structure for technology: The Bayh– Dole effect on university– industry–government relations in developing economy. Science and Public Policy, 43(1):29–45, February 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/spp.oxfordjournals.org/
content/43/1/29.

Woolley:2015:RCS

[WSBTM15] Richard Woolley, Mabel Sánchez-Barrioluengo, Tim Turpin, and Jane Marceau. Research collaboration in the social sciences: What factors are associated with disciplinary and interdisciplinary collaboration? Science and Public Policy, 42(4):567–582, August 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wu:2013:CSD

[Wu13] Yonghong Wu. The crossstate distribution of federal funding in the USA: The case of financing academic research and development. Science and Public Policy, 40(3):316–326, June 2013. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wang:2017:UPI

[WW17] Chadwick (Chengwei) Wang and Luhao Wang. Unfolding policies for innovation intermediaries in China: a discourse network analysis. Science and Public Policy, 44(3): 354–??, June 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        http://academic.oup.com/
                                        spp/article-abstract/44/
                                        3/354/2525575/Unfolding-
                                        policies-for-innovation-
                                        intermediaries.

White:2010:CSL

[WWL^{+}10] Dave D. White, Amber Wutich, Kelli L. Larson, Patricia Gober, Timothy Lant, and Clea Senneville. Credibility, salience, and legitimacy of boundary objects: water managers’ assessment of a simulation model in an immersive decision theater. Science and Public Policy, 37(3):219– 232, April 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Xu:2015:ATF

[Xu15] Lei Xu. Advancing theories of firm growth by scrutinizing assumptions. Science and Public Policy, 42(6):902–904, December 2015. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Yang:2018:CKD

[Yan18]  Lihua Yang.  Collabora-

tive knowledge-driven governance: Types and mechanisms of collaboration between science, social science, and local knowledge. Science and Public Policy, 45(1):53–73, February 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/45/1/53/4159474.

Yaqub:2012:INM

[Yaq12] Ohid Yaqub. Innovation networks: More than just a metaphor? Science and Public Policy, 39(2):282–283, March 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Yang:2014:NAE

[YH14] Chang Hoon Yang and Jungeun Heo. Network analysis to evaluate crossdisciplinary research collaborations: The Human Sensing Research Center, Korea. Science and Public Policy, 41 (6):734–749, December 2014. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Yang:2012:WCT

[YHR12] Yan Yang, Jette Egelund Holgaard, and Arne Remmen.

What can Triple Helix frameworks offer to the analysis of eco-innovation dynamics? Theoretical and methodological considerations. Science and Public Policy, 39(3):373– 385, June 2012. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Yang:2017:TPT

[YHSH17] Yang Yang, Jin Hong, Ge Song, and Song Hong. Technology policy, technology strategy and innovation performance: evidence from Chinese aircraft and spacecraft manufacturing. Science and Public Policy, 44(5):620–630, October 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

//academic.oup.com/spp/
article/44/5/620/2877116.

Youtie:2017:ESS

[YPH17] Jan Youtie, Alan L. Porter, and Ying Huang. Early social science research about Big Data. Science and Public Policy, 44(1):65–??, February 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
1/65/2525527/Early-social-
science-research-about-
Big-Data.

Youtie:2019:RNE

[YSRF19] Jan Youtie, Philip Shapira, Michael Reinsborough, and Erik Fisher. Research network emergence: Societal issues in nanotechnology and the center for nanotechnology in society. Science and Public Policy, 46(1):126–135, February 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/1/126/5039639.

Yun:2015:ASE

[YWH^{+}15] JinHyo Joseph Yun, DongKyu Won, ByungYong Hwang, Jin-Won Kang, and DongHwan Kim. Analysing and simulating the effects of open innovation policies: Application of the results to Cambodia. Science and Public Policy, 42 (6):743–760, December 2015. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Yokakul:2011:KSS

[YZ11] Nattaka Yokakul and Dr. Girma Zawdie. The knowledge sphere, social capital and growth of indigenous knowledgebased SMEs in the Thai dessert industry. Science and Public Policy, 38(1):19–29, February 2011. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Zavale:2018:EVC

[Zav18] Nelson Casimiro Zavale. Expansion versus contribution of higher education in Africa: University–industry linkages in Mozambique from companies’ perspective. Science and Public Policy, 45 (5):645–660, October 1, 2018. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/45/5/645/4780191.

Zhang:2019:IEW

[Zha19]  Qiantao Zhang.  Innova-

tion and its enemies: Why people resist new technologies. Science and Public Policy, 46(4):632–633, August 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/46/4/632/5396727.

Zabala-Iturriagagoitia:2010:IPM

[ZI10] Jon Mikel Zabala-Iturriagagoitia. Innovation in policy and not more policies for innovation? Science and Public Policy, 37(1):80–82, February 2010. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Zou:2019:AOE

[ZLGG19] Bo Zou, Yanxia Li, Jinyu Guo, and Feng Guo. Antecedents and outcome of entrepreneurial identification: The moderating effect of role orientation. Science and Public Policy, 46(4):541–551, August 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

//academic.oup.com/spp/
article/46/4/541/5396728.

Zhang:2017:UKE

[ZLL17] Qiantao Zhang, Charles Larkin, and Brian M. Lucey. Universities, knowledge exchange and policy: a comparative study of Ireland and the UK. Science and Public Policy, 44(2): 174–??, April 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

http://academic.oup.com/
spp/article-abstract/44/
2/174/2525550/Universities-
knowledge-exchange-and-
policy-A.

Zhou:2016:CKB

[ZLLU16] Yuan Zhou, Xin Li, Rasmus Lema, and Frauke Urban. Comparing the knowledge bases of wind turbine firms in Asia and Europe: Patent trajectories, networks, and globalisation. Science and Public Policy, 43(4):476–491, August 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Zacharewicz:2019:PBR

[ZLRJ19] Thomas Zacharewicz, Benedetto Lepori, Emanuela Reale, and Koen Jonkers. Performancebased research funding in EU member states — a comparative assessment. Science and Public Policy, 46(1):105–115, February 2019. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:

                                        //academic.oup.com/spp/
                                        article/46/1/105/5037253.

Zapp:2017:MTM

[ZP17] Mike Zapp and Justin J. W. Powell. Moving towards Mode 2? Evidence-based policy-making and the changing conditions for educational research in Germany. Science and Public Policy, 44(5):645–655, October 1, 2017. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL https:

                                        //academic.oup.com/spp/
                                        article/44/5/645/2900558.

Zhang:2016:HSN

[ZWL16] Jing Zhang, Guoyu Wang, and Deming Lin. High support for nanotechnology in China: a case study in Dalian. Science and Public Policy, 43(1):115–127, February 2016. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /spp.oxfordjournals.org/
                                        content/43/1/115.